### v1

# Beyond Statistical Estimation: Differentially Private Individual Computation in the Shuffle Model

Abstract
The shuffle model of differential privacy (DP) has recently emerged as a powerful one for decentralized computation without fully trustable parties. Since it anonymizes and permutes messages from clients through a shuffler, the privacy can be amplified and utility can be improved. However, the shuffling procedure in turn restricts its applications only to statistical tasks that are permutation-invariant.
This work explores the feasibility of shuffle privacy amplification for prevalent non-statistical computations: spatial crowdsourcing, combinatorial optimization, location-based social systems, and federated learning with incentives, which suffer either computationally intractability or intolerable utility loss in existing approaches (e.g., secure MPC and local DP). We proposes a new paradigm of shuffle model that can provide critical security functionalities like message authorization and result access control pertained to permutation-equivariant tasks, meanwhile maintaining the most of privacy amplification effects. It incurs almost the same computation/communication costs on the client side as the non-private setting, and permits the server to run arbitrary algorithms on (noisy) client information in plaintext. Our novel technique is introducing statistically random identity into DP and force identical random distribution on all clients, so as to support secure functionalities even after message shuffling and to maintain privacy amplification simultaneously. Given that existing DP randomizers fails in the new shuffle model, we also propose a new mechanism and prove its optimality therein. Experimental results on spatial crowdsourcing, location-based social system, and federated learning with incentives, show that our paradigm and mechanism is fast as non-private settings, while reducing up to $90\%$ error and increasing utility performance indicates by $100\%$-$300\%$ relatively, and can be practical under reasonable privacy budget.
Index Terms:
differential privacy, shuffle model, digital identities, spatial crowdsourcing, federated learning
I Introduction
Personal information fuels a wide array of data-driven applications, e.g. statistical analytics, machine learning, recommendation systems, spatial crowdsourcing, mobile crowdsensing, combinatorial optimization, and location-based social systems. These applications deliver substantial value but depend on data collected from users, which is a prime target for attacks and carries a high risk of leakage or abuse. Data privacy concerns are escalating, especially after several high-profile data breach incidents. Despite the introduction of stricter privacy laws such as the EU’s General Data Protection Regulation, the California Consumer Privacy Act, and China’s Personal Information Protection Law, many users still distrust service providers and are hesitant to consent to the use of their data. To bridge this trust gap and encourage user participation, significant efforts have been made recently to develop privacy-enhancing technologies that enable private data processing without relying on a trusted third party.
Two prominent technologies addressing this problem are secure multiparty computation (MPC) [88] and differential privacy (DP) [28]. MPC uses interactive cryptographic protocols to allow mutually untrusted parties to jointly evaluate a function on their private data, with each party receiving an output but nothing else. However, the cryptographic nature of MPC imposes heavy overheads, making it challenging to scale for practical use. DP, on the other hand, adds random noise to data to ensure privacy, making it computationally more feasible since the computation is carried out on sanitized data without heavy cryptographic machinery. Traditional central DP still requires a trusted curator [28], but the local model (LDP [53]) allows each user to sanitize their data locally before sending it out. Due to its practicality and low trust assumption, LDP has been adopted by companies like Apple [74], Microsoft [24], and Google [30] in their real-world systems. However, LDP’s drawback is that each user must independently inject sufficient random noise to guarantee privacy, leading to significant utility loss.
Recently, within the realm of DP research, the shuffle model [17, 29] has emerged. This model introduces a shuffler that randomly permutes messages from users and then sends these anonymized messages to the computing server or analyzer. The server can then compute on these messages to derive the result. Trust in the shuffler is minimized because each user encrypts their locally sanitized data using the server’s public key. This way, the shuffler remains oblivious to the messages received from the users. A key advantage of the shuffle model over LDP is utility. It has been proven that anonymizing/shuffling messages amplifies the privacy guarantee provided by the local randomizer used by the users. For instance, shuffled messages from $n$ users each adopting local $\epsilon$-DP actually preserve differential privacy at the level $\epsilon_{c}=\tilde{O}(\sqrt{e^{\epsilon}/n})$ [31, 32]. Consequently, to achieve a predefined global privacy goal, each user consumes less privacy budget when sanitizing their data locally, meaning less noise needs to be added. This significantly improves the accuracy of the final result. Owing to these utility advantages, extensive studies have been conducted within the shuffle model, e.g., [10, 37, 38, 38, 39].
That said, the shuffle model has a noticeable limitation. The privacy amplification effect relies heavily on anonymizing/shuffling messages, which significantly restricts the types of computation that can be performed. So far, the sole form of computation achievable within the shuffle model is statistical estimation, i.e., the server takes the shuffled messages, aggregates them, and computes a single output from them, e.g., a count, sum, or histogram. However, many real-world applications are non-statistical in nature. When multiple users pool their data together for joint computation, they expect an individualized output that may differ for each user. We coined the term “individual computation” for such tasks. Examples of individual computation tasks include:
  • I.
    Combinatorial optimization: Spatial crowdsourcing [80], advertisement allocation [61], and general combinatorial optimization [55], where two or more parties are often matched together based on their private information. Each party should get their own list of “best matches” whatever that means.
  • II.
    Information retrieval: Mobile search [52], location-based systems [4, 22], where the query results (e.g., nearby restaurants or neighboring users) depend on the private information of the inquirer.
  • III.
    Incentive mechanisms: In federated learning [89] or data crowdsensing [72], incentives play a vital role to encourage well-behaved participation. The amount of rewarding incentive must be computed for individual users based on their contribution (e.g., via Shapley values [69]).
At first glance, private individual computation seems unattainable within the shuffle model because the need for individualized output clashes with the anonymization required for privacy amplification. However, this is not necessarily true. Our observation is that many individual computing functionalities are equivariant to shuffling. That is, the permutation applied to the inputs does not affect the computation. This means shuffling does not prevent the server from producing personalized answers for each client—the server does not know and does not need to know which answer is for whom. Nevertheless, there is a problem: how to pass the output back to the right user without breaking anonymity. A straightforward approach is to let the shuffler maintain a long-term duplex connection channel between each client and the server (e.g., as in an Onion routing network [41]), but this method is costly (due to storing communication states) and may be subject to de-anonymization attacks on anonymous channels [63, 65]. Additionally, the duplex connection is not supported in current shuffler implementations, such as the one in the seminal work [17]. This is the first technical challenge we need to address.
The second challenge we face is how to design optimal randomizers for individual computation in the shuffle model. Often the randomizer in a shuffle model protocol needs to be tailored to the tasks. For statistical tasks within the shuffle model, several studies have developed near-optimal randomizers, as seen in histogram estimation [36, 31] and one-dimensional summation estimation [9]. However, the new setting of individual computation is different: the focus is on the accuracy of the output for each user (instead of the statistical accuracy of the population). This gap renders existing randomization strategies (e.g., randomizers utilizing dimension sampling, budget splitting, or data sketching, as reviewed in [86]), along with prevalent randomizers (such as adding Laplace [28] or Staircase [35] noises), less effective for the new setting. Therefore we need to take a fresh look at the fundamental privacy-utility trade-offs and re-design the underlying randomizers.
In this work, we introduce a new paradigm of the shuffle model that provides security functionalities (e.g., message authorization, integrity verification, and access control) for permutation-equivariant computations, and term it the private individual computation model. The properties of the model are as follows.
  • I.
    Security: The server and other clients can verify the owner of a message using a fingerprint submitted alongside the message. The server can also securely deliver personalized backward information to each owner (i.e., access control). These functionalities enable the secure completion of equivariant tasks.
  • II.
    Privacy: There is no degradation of local privacy while providing these additional security functionalities. Furthermore, when task results are sparse (e.g., each user is matched to one taxi driver in spatial crowdsourcing), the degradation of shuffle privacy amplification is negligible. Additionally, our local and shuffle privacy guarantees are information-theoretic (as long as the shuffling is statistically random). This distinguishes our new model from computational DP [62], which also combines cryptographic functions and privacy but guarantees differential privacy only against polynomial-power adversaries.
  • III.
    Utility: Since clients still benefit from privacy amplification via shuffling, the utility is significantly improved. With the increasing of number of clients, the utility increases and approaches to non-private settings. In addition, the server in our model can run arbitrary algorithms on the (noisy) messages in plaintext. Our model is compatible with existing achievements and future advancements in server-side algorithms (e.g., noisy-aware matching algorithms for spatial crowdsourcing).
  • IV.
    Scalability: As the server runs algorithms on plain-text messages and there is no heavy computation burden on the client side, our model is highly efficient and scales to systems with massive clients. This eliminates the need for secure multi-party computation which necessitates impractical costs. In our model, the computation/communication costs are essentially the same to non-private settings (see Appendix -I).
Main techniques. Our approach to realizing additional security functionalities is based on two critical observations: (1) If the message payload satisfies $\epsilon$-LDP, appending arbitrary random variables (independent to the private value) to the payload incurs no additional privacy loss. Specifically, when random variables are sampled from an identical distribution for all clients, there is no excess loss in privacy amplification via shuffling. (2) Post-processing the message payload and appended random variables does not incur excess privacy loss. The first observation motivates us to append a random fingerprint (e.g., a public key) to the payload, allowing the client to use the corresponding secret key to negotiate shared secret and establish secure communication with other parties in later phases. A side consequence is that the server can use the public key pertained to each message to securely deliver information to the message owner. The second observation motivates us to further append a signature to protect the integrity of the fingerprint and payload.
Challenge (2): Optimal randomizers for non-statistical computations in the shuffle model. In the local model of DP, existing research [83, 66, 77, 81, 82] primarily focuses on designing randomizers in a high privacy regime (e.g., local budget $\epsilon=O(1)$). Conversely, in the shuffle model, the local budget can extend to $\tilde{O}(\log(n))$. For statistical tasks within this model, several studies have developed near-optimal randomizers, as seen in histogram estimation [36, 31] and one-dimensional summation estimation [9]. However, the new setting of non-statistical computation tasks in the shuffle model—often involving multi-dimensional user data and emphasizing an individual report’s accuracy (instead of statistical accuracy of the population)—has not been explored. This gap renders existing randomization strategies (e.g., randomizers utilizing dimension sampling, budget splitting, or data sketching, as reviewed in [86]), along with prevalent randomizers (such as adding Laplace [28] or Staircase [35] noises), less effective for non-statistical computation in the shuffle model. The fundamental privacy-utility trade-offs in non-statistical computation for this new model also present significant theoretical interest.
Our work establishes the error lower bounds for single reports in non-statistical computation within the shuffle model, demonstrating the sub-optimality of existing randomizers. To address this, we introduce a novel series of local randomizers—the Minkowski response—aligned with these error lower bounds. The core principle involves reporting values within a calibrated Minkowski distance from the true value with high probability.
The main contributions of this work are summarized as follows:
  • We propose equivariant shuffle model of DP, which supports various security functionalities and benefit large-scale equivariant computations from privacy amplification. The new model exhibits almost the same computation/communication costs on client/server side as non-private settings.
  • We establish tight minimax error lower bounds of $\tilde{O}((n\epsilon^{2}_{c})^{-\frac{2}{d+2}})$ for the new model of DP, where $d$ is the user data dimension, $\epsilon_{c}$ is global privacy budget, and $n$ is the number of users. Since existing data randomizers fail, we then design an attainable randomizer: Minkowski response, for the new model to match the lower bound.
  • Through experiments on three typical permutation-equivariant tasks: privacy-preserving spatial crowdsourcing, location-based nearest neighbor queries, and federated learning with incentives, we demonstrate that the new model, in conjunction with proposed randomizers, reduces error by $90\%$ meanwhile incurring negligible excess burdens in both the client and server side.
Organization. The remainder of this paper is organized as follows. Section II reviews related works. Section III provides preliminary knowledge about privacy definitions and security primitives. Section IV formalizes the problem setting. Section V presents the PIC model. Section VI provides optimal LDP randomizers. Section VII evaluates the utility and efficiency performances of our proposals. Finally, Section VIII concludes the paper.
II Related Works
This section reviews various approaches to private computation, primarily concentrating on permutation-equivariant tasks.
II-A Cryptography Methods
Numerous works aim to securely solve matching and allocation problems involving personal sensitive information from multiple parties, revealing only the output. They utilize homomorphic encryption [2, 44], secret sharing [3] or garbled circuits (e.g., in [40, 15]) to perform secure multi-party computations (MPC). Specifically, when there is a trusted message shuffler, every client can establish information-theoretically secure communication with other clients and accomplish arbitrary multi-party computation (including DP algorithms) within a few rounds [48, 13]. In general, MPC-based approaches impose high computation and interaction overheads, especially when the computation algorithm gets sophisticated. As comparison, the proposed paradigm in this work permits running arbitrary algorithms on plaintext, and requires only semi-trustness/honest-but-curious assumption on the shuffler (as the shuffler only sees ciphertexts, see Section V).
II-B Curator and Local DP Methods
Many works study matching, allocation, or general combinatorial optimization problems within the curator DP model [28] in the presence of a trusted party collecting raw data from clients (e.g., in [60, 23, 76, 75]). Since the assumption of a trustworthy party is often unrealistic in decentralized settings, many studies adopt the local model of DP (e.g., in [83, 66, 77, 81]), where each client sanitizes data locally and sends the noisy data to the server for executing corresponding matching/allocation algorithms. As each client must injects sufficient noises into data to satisfy local DP, the execution results often maintain low utility.
II-C Shuffle Model of DP
The recently proposed shuffle model [17, 29] combines the advantages of the curator model (e.g., high utility) and the local model (e.g., minimal trust). Depending on the number of messages each client can send to the intermediate shuffler, the shuffle model can be categorized as single-message [9, 29, 32] and multi-message [10, 36]. The single-message shuffle model leverages privacy amplification via shuffling to enhance data utility compared to the local model. A substantial body of work [29, 31, 32] demonstrates that $n$ shuffled messages from clients, each adopting a same $\epsilon$-LDP randomizer, can actually preserve $\tilde{O}(\sqrt{e^{\epsilon}/n})$-DP. By removing the constraint of sending one message, the multi-message shuffle model can achieve better utility than the single-message model and might be comparable to the curator model (e.g., in [10, 36]). However, each multi-message protocol is tailored to a specific statistical query (e.g., summation), rendering them unsuitable for permutation-equivariant tasks with non-linear computations. There is a line of works on the shuffle model for private information retrieval (e.g., in [48, 47, 34] with cryptography security and in [78, 1] with statistical DP), where the query is represented as multiple secret shares before sent to the shuffler, and the server holding the database entries returns linear-transformed entries for each query, using the duplex shuffled communication channel. This kind of duplex-communication shuffle model can be vulnerable to anonymity attacks [63, 65], and is pertained to the linear computation in private information retrieval. It can not be extended to other PIC tasks that involve with non-linear computations, and can not handle party-to-party communication in tasks like spatial crowdsourcing and social systems.
Overall, existing works in the shuffle model primarily focus on statistical queries that are naturally permutation-invariant. This work, for the first time, explore the shuffle model for non-statistical applications (i.e., combinatorial optimization, location-based social systems, and incentive mechanisms). Technically, our work extends the shuffle model to support identity-related security primitives, such as message authentication and access control.
II-D Combining Cryptography and DP
While cryptographic tools can protect data secrecy during multi-party computation, they do not necessarily preserve output’s privacy. DP can be employed to enhance the privacy of the outputting result of secure multiparty computation through decentralized noisy addition [43]. To account for privacy loss due to intermediate encrypted views in MPC, researchers have proposed the relaxed notion of computational DP [62] against polynomial-time adversaries. Computational DP protocols inherited the computation/communication complexity of MPC (refer to garbled-circuits-based approaches in Table VIII).
In contrast to enhancing relaxed privacy in the cryptographic cipher domain, our work initializes the study in the opposite direction, by integrating simple security functionalities into the privacy realm. Our model exposes intermediate results with information-theoretical DP guarantee. Therefore, it is significantly more efficient (i.e., without computation on ciphertext and requiring few rounds of interaction) and flexible (i.e., supporting arbitrary computations on noisy plaintext) than MPC with computational DP.
III Background and Problem Settings
TABLE I: List of notations.
Notation Description
$[i]$$\{1,2,...,i\}$
$[i:j]$$\{i,i+1,...,j\}$
$\mathcal{S}$ the shuffling procedure
$\mathcal{R}$ the randomization algorithm
$G_{i}$ the $i$-th group of users ($i\in[m]$)
$n_{i}$ the number of users in group $G_{i}$
$u_{i,j}$ the $j$-th user in group $G_{i}$ where $j\in[n_{i}]$
$d$ the dimension of user’s data
$\mathbb{X}$ the domain of input data
$\mathbb{Y}$ the domain of a sanitized message
$\epsilon$ the local privacy budget
$\epsilon_{c}$ the amplified privacy level
$sk$, $pk$ the secret key and the public key, respectively
$\lambda$ the security parameter of cryptography
III-A Privacy Definitions
Definition III.1 (Hockey-stick divergence [70])
For two probability distributions $P$ and $Q$, the Hockey-stick divergence between them with parameter $e^{\epsilon}$ is as follows:
$D_{\epsilon}(P||Q)=\int_{z\in\mathcal{Z}}\max\{0,P(z)-e^{\epsilon}Q(z)\}\mathrm{d}z.$
Differential privacy imposes divergence constraints on output probability distributions with respect to changes in the input. In the curator model of differential privacy, a trusted party collects raw data $x_{i}\in\mathbb{X}$ from all users to form a dataset $T=\{x_{1},\ldots,x_{n}\}$ and applies a randomization algorithm $\mathcal{R}$ to release query results $\mathcal{R}(T)$. For two datasets $T$ and $T^{\prime}$ of the same size and differing in only one element, they are referred to as neighboring datasets. Differential privacy ensures that the Hockey-stick divergence between $\mathcal{R}(T)$ and $\mathcal{R}(T^{\prime})$ is bounded by a sufficiently small value (i.e., $\delta=O(1/n)$). The formal definition of curator differential privacy is as follows:
Definition III.2 ($(\epsilon,\delta)$-DP [28])
A randomization mechanism $\mathcal{R}$ satisfies $(\epsilon,\delta)$-differential privacy iff $\mathcal{R}(T)$ and $\mathcal{R}(T^{\prime})$ are $(\epsilon,\delta)$-indistinguishable for any neighboring datasets $T,T^{\prime}\in\mathbb{X}^{n}$. That is, $\max(D_{\epsilon}(\mathcal{R}(T)\|\mathcal{R}(T^{\prime})),D_{\epsilon}(\mathcal{R}(T^{\prime})\|\mathcal{R}(T)))\leq\delta$.
In the local model of differential privacy, each user applies a randomization mechanism $\mathcal{R}$ to their own data $x_{i}$, with the objective of ensuring that the Hockey-stick divergence between $\mathcal{R}(x)$ and $\mathcal{R}(x^{\prime})$ is 00 for any $x,x^{\prime}\in\mathbb{X}$ (see Definition III.3).
Definition III.3 (local $\epsilon$-DP [53])
A randomization mechanism $\mathcal{R}$ satisfies local $\epsilon$-DP iff $D_{\epsilon}(\mathcal{R}(x)||\mathcal{R}(x^{\prime}))=0$ for any $x,x^{\prime}\in\mathbb{X}$.
The data processing inequality is considered a key feature of distance measures used for evaluating privacy. It asserts that the privacy guarantee cannot be weakened by further analysis of a private mechanism’s output.
Definition III.4 (Data processing inequality)
A distance measure $D:\Delta(\mathcal{S})\times\Delta(\mathcal{S})\to[0,\infty]$ on the space of probability distributions satisfies the data processing inequality if, for all distributions $P$ and $Q$ in $\Delta(\mathcal{S})$ and for all (possibly randomized) functions $g:\mathcal{S}\to\mathcal{S^{\prime}}$,
$D(g(P)\|g(Q))\leq D(P\|Q).$
III-B The Classical Shuffle Model
Following the conventions of the randomize-then-shuffle model [21, 9], we define a single-message shuffle protocol $\mathcal{P}$ as a list of algorithms $\mathcal{P}=(\mathcal{R},\mathcal{A})$, where $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ is local randomizer on client side, and $\mathcal{A}:\mathbb{Y}^{n}\to\mathbb{Z}$ is the analyzer on the server side. We refer to $\mathbb{Y}$ as the protocol’s message space and $\mathbb{Z}$ as the output space. The overall protocol implements a mechanism $\mathcal{P}:\mathbb{X}^{n}\to\mathbb{Z}$ as follows. User $i$ holds a data record $x_{i}$ and a local randomizer $\mathcal{R}$, then computes a message $y_{i}=\mathcal{R}(x_{i})$. The messages $y_{1},\ldots,y_{n}$ are shuffled and submitted to the analyzer. We denote the random shuffling step as $\mathcal{S}(y_{1},\ldots,y_{n})$, where $\mathcal{S}:\mathbb{Y}^{n}\to\mathbb{Y}^{n}$ is a shuffler that applies a uniform-random permutation to its inputs. In summary, the output of $\mathcal{P}(x_{1},\ldots,x_{n})$ is represented by $\mathcal{A}\circ\mathcal{S}\circ\mathcal{R}(X)=\mathcal{A}(\mathcal{S}(\mathcal{R}(x_{1}),\ldots,\mathcal{R}(x_{n})))$.
In particular, when all users adopt an identical $\epsilon$-LDP mechanism $\mathcal{R}$, recent works [31, 32] have derived that $n$ shuffled $\epsilon_{l}$-LDP messages satisfy $(O((1-e^{-\epsilon})\sqrt{e^{\epsilon}\log(1/\delta)/n}),\delta)$-DP. We denote the amplified privacy level as:
$\epsilon_{c}=\texttt{Amplify}(\epsilon,\delta,n),$
the tight value of which can also be numerically computed [56, 85].
III-C Security Primitives
We employ several well-established security primitives for secure point-to-point communication, including key aggreement, symmetry encryption and digital signature.
A Diffie-Hellman key agreement protocol with security parameter $\lambda$ comprises two primary steps:
  • $sk_{i},pk_{i}\leftarrow\textsf{DH-Gen}(\lambda)$: a party $i$ generates a public-private key pair $(sk_{i},pk_{i})$, where $pk_{i}$ is the public key and $sk_{i}$ is the secret/private key;
  • $K_{ij}\leftarrow\textsf{DH-Compute}(sk_{i},pk_{j})$: the two parties involved in the key exchange use their private keys $sk_{i}$ and $sk_{j}$ to compute a shared key $K_{ij}$ from the public keys $pk_{j}$ and $pk_{i}$ of each party (i.e., $K_{ij}=K_{ji}$.
The resulting shared secret key, $K_{ij}$, can be used for subsequent data symmetry encryption/decryption. Specifically, a symmetry encryption scheme with secret key $sk$ consists of two steps:
  • $c\leftarrow\textsf{Enc}(sk,x)$: generates a ciphertext $c$ by encrypting data $x$ with the encryption key $sk$;
  • $x\leftarrow\textsf{Dec}(sk,c)$: decrypts ciphertext $c$ with the key $sk$, outputting $x$.
A digital signature scheme permits the receipt to authorize ciphertexts and to ensure message integrity, through the following steps:
  • $t\leftarrow\textsf{Sign}(sk,c)$: returns a tag $t$ as the output of signing data $c$ with the private key $sk$;
  • $accept/reject\leftarrow\textsf{Verify}(pub,t,c)$: outputs $accept$ when $t=\textsf{Sign}(sk,c)$, and outputs $reject$ with a probability of $1-negl(k)$ when $t\neq\textsf{Sign}(sk,c)$.
IV Problem Settings
We aim to extend the capability of the shuffle model to support private individual computation. This would have multiple benefits: (1) we can get a formal privacy guarantee (in the DP sense) on a much wider range of data-driven computational tasks; (2) because the computation is done on already sanitized data, we can avoid heavy cryptographic protocols and scale to large data and/or user size; (3) the privacy amplification effect will allow much better utility, compared to sanitizing and computing directly using LDP. In this section, we will set the stage by first presenting a few motivating applications, and then formally defining private individual computation as an ideal functionality.
Figure 1: An illustration of conceptual overview of PIC model for spatial crowdsourcing.
IV-A Motivating Applications
We describe three prevalent exemplar computation tasks: spatial crowdsourcing, location-based services, and federated learning with incentives:
Spatial crowdsourcing. A spatial crowdsourcing system normally consists of three roles: users, workers, and the orchestrating server (the service platform), and proceed as following major steps:
  • I.
    Task submission: each participated user $i\in G_{a}$ submits a task (e.g., taxi calling requests, sensing/photo requests) $x_{i}=(i,l_{i},v_{i})$ containing the location $l_{i}$ and possibly other information $v_{i}$;
  • II.
    Worker reporting: every enrolled worker $j\in G_{b}$ reports $x_{j}$ containing the location $l_{j}$ and other information $v_{j}$;
  • III.
    Task assignment: the server receives $\{x_{i}\}_{i\in G_{a}}$ and $\{x_{j}\}_{k\in G_{b}}$, and outputs a matching $M:G_{a}\times G_{b}\mapsto\{0,1\}$ between $G_{a}$ and $G_{b}$ based on some criterion (e.g., minimizing total traveling costs, or maximizing matches).
  • IV.
    Task performing: users and workers retrieve matching results $M$ and collaboratively complete the task.
Location-based services. In a location-based service, a set of users and a server interact as the following:
  • I.
    Querying: Each participating user $i\in G_{a}$ submits a request $x_{i}=(i,l_{i},v_{i})$, which includes information like location $l_{i}$ and preferences $v_{i}$.
  • II.
    Generating recommendation: The server receives $\{x_{i}\}_{i\in G_{a}}$ and generates a list of recommendations for each user based on specific criteria, such as proximity and user preferences.
  • III.
    Retrieving: Finally, each user retrieves the recommendation results.
Federated learning with incentives. Federated learning involves a set of clients and a server:
  • I.
    Submitting gradient: Each client $i\in G_{a}$ in each epoch computes an intermediate gradient information $x_{i}\in[-1,1]^{d}$ with its local model and data, then submits to the server;
  • II.
    Computing incentive: The server computes the averaged gradient $\overline{x}=\frac{1}{n}\sum_{i\in[n]}x_{i}$. To incentive participation, the server may reward clients with monetary tokens (e.g., via cryptocurrency) according to a profit allocation algorithm $V:[-1,1]^{d\times n}\times[-1,1]^{d}\mapsto[0,1]$ (e.g., Shapley value [69]).
  • II.
    Receiving incentive: Each client retrieves the token and claims its monetary incentive.
A fundamental distinction between the above applications and those currently studied in the shuffle model is that now each participant expects an output that differs individually, instead of a single collectively aggregated output. Informally, in such applications, we would need to safeguard data privacy so that apart from the party who generates the data, no one should be certain about that party’s data (up to the leakage allowed by DP). To be able to amplify privacy, we also need to maintain anonymity so that for any message, an adversary (e.g. the server, an observer, an unmatched user) should only know this message comes from a user belonging to a particular group, but nothing more. Note that following the convention in the shuffle model, the shuffler is trusted for providing anonymity. So the shuffler will know the random permutation used in the shuffling process and will not leak it to other parties, although we prevent the shuffler from observing plaintext messages through encryption.
IV-B The Ideal Functionality
Following the ideal-real world paradigm, we capture private individual computation formally as an ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$, which is shown in Figure 2.
Functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ Parameters: $m\in\mathbb{N}$; $m$ groups of parties $G_{1},G_{2},\cdots,G_{m}$, where $n_{i}=|G_{i}|$ denotes the number of parties in group $G_{i}$; the data randomization mechanisms $\mathcal{R}_{i}$ for group $G_{i}$; a server ${S}$. Functionality: Upon receiving $n=\sum_{i\in[m]}n_{i}$ inputs $\{x_{i,j}\}_{i\in[m],j\in[|G_{i}|]}$ from all clients, and the description of a function $f$ to be computed over parties’ inputs from the server, do the following: Compute $x^{\prime}_{i,j}\leftarrow R_{i}(x_{i,j})$ for all $i\in[m]$ and $j\in[|G_{i}|]$. Sample $m$ random permutations $\pi_{1},\pi_{2},\cdots,\pi_{m}$, where $\pi_{i}:[n_{i}]\rightarrow[n_{i}]$, perform shuffle over inputs and obtain $L=\left(\{x_{1,\pi_{1}(j)}^{\prime}\}_{j\in[n_{1}]},\cdots,x_{m,\pi_{m}(j)}^{\prime}\}_{j\in[n_{m}]}\right)$, and compute ($\left(\{y_{1,\pi_{1}(j)}\}_{j\in[n_{1}]},\cdots,\{y_{m,\pi_{m}(j)}\}_{j\in[n_{m}]}\right)\leftarrow f(L)$. Send $y_{i,\pi_{i}(j)}$ to party $u_{i,j}$ for all $i\in[m]$ and $j\in n_{i}$. Additionally, send $(L,f(L))$ to $S$.
Figure 2: The functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$
Essentially, the ideal functionality represents a fully trusted party that interacts with $m$ groups of clients and one server $S$. It sanitizes the input from each client, then shuffles them randomly, and applies a function from the server to compute the output for each client. At the end, each client receives its function output, and the server gets the list of shuffled sanitized data from the clients and a list of the function output. This captures the functional requirement of individual computation in the real world: a computing task is performed by a server using the joint inputs from a set of clients (which are sanitized and shuffled), and each client gets an individualized output. It also captures the privacy requirements: each party should receive precisely the specified output, nothing more.
Remark A careful reader may notice that the ideal function does not capture differential privacy. This is intentional for the sake of security analysis. Later in the security analysis, we will decouple different privacy requirements into two sets of proofs: the first set shows that our concrete protocol executed by real-world parties realizes the ideal functionality, meaning that no additional information about parties’ inputs is revealed, except the output given to each party; then we prove that the output conforms with differential privacy.
V A Concrete Protocol
V-A The Protocol
We now present a concrete protocol for private individual computation. Because there are already secure protocols for shuffling, to simplify the description, we present the protocol in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model, in which parties can communicate as usual, and in addition have access to an ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ that does the shuffling. The ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ (Figure 3) is parameterized with a list of parties that are corrupted by the adversary and collude with the server. For those parties, the adversary should know the correspondence between their messages before and after shuffling, hence $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ leaks this part of the permutation to the adversary.
Functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ Parameters: $n\in\mathbb{N}$; $n$ parties $P_{1},P_{2},\cdots,P_{n}$; a server ${S}$; the corrupted party set $\mathcal{C}$; the leakage $\mathcal{L}(\pi)=\{(i,\pi(i))\}_{i\in[\mathcal{C}]}$ for the permutation $\pi$ being used. Functionality: Upon receiving $n$ inputs $\{x_{i}\}_{i\in[n]}$ from $P_{1},P_{2},\cdots,P_{n}$, respectively. Sample a random permutation $\pi\in\mathbf{S}_{n}$. Define $\{y_{i}\}_{i\in[n]}$ such that $y_{i}=x_{\pi(i)}$. Send $\{y_{i}\}_{i\in[n]}$ to the server $S$. Additionally if $S\in\mathcal{C}$, send $\mathcal{L}(\pi)$ to the adversary $S$.
Figure 3: The functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$
The Private Individual Computation protocol is outlined below:
  1. 1.
[…]
  2. 2.
    We denote the $j$-th client in group $G_{i}$ as $u_{i,j}$. Each client generates a key pair $(pk_{i,j},sk_{i,j})\leftarrow\textsf{Gen}(\lambda)$. Each client then randomizes its private data $x_{i,j}$ with mechanism $\mathcal{R}_{i}$ and obtains $x_{i,j}^{\prime}\leftarrow\mathcal{R}_{i}(x_{i,j})$. Then the sanitized input is concatenated with its own public key, and encrypted with the server’s public key $x_{i,j}^{\prime\prime}\leftarrow\textsf{Enc}_{pk_{c}}(pk_{i,j}||x^{\prime}_{i,j})$.
  3. 3.
    Each client in group $G_{i}$ invokes $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ with $x_{i,j}^{\prime\prime}$, and $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ outputs the shuffled messages $\{x_{i,\pi_{i}(j)}^{\prime\prime}\}_{j\in[n_{i}]}\leftarrow\mathcal{S}(\{x_{i,j}^{\prime\prime}\}_{j\in[n_{i}]})$ to the server, where $\pi^{-1}_{i}$ is the (secret) random permutation used during $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ for group $G_{i}$.
  4. 4.
    The server decrypts each set of shuffled messages and obtains a list $L$ for all $m$ groups as:
    $L=$ ($\left(\{pk_{1,\pi_{1}(j)}||x_{1,\pi_{1}(j)}^{\prime}\}_{j\in[n_{1}]},\cdots\right.$
    $\ \ \ \left.\{pk_{m,\pi_{m}(j)}||x_{m,\pi_{m}(j)}^{\prime}\}_{j\in[n_{m}]}\right)$
    It then computes the function $f$ over $L$ to produce output for each anonymous client:
    ($\left(\{y_{1,\pi_{1}(j)}\}_{j\in[n_{1}]},\cdots,\{y_{m,\pi_{m}(j)}\}_{j\in[n_{m}]}\right)\leftarrow f(L)$
  5. 5.
    The server publishes the computation results to a public bulletin board as a list of pairs: $\{(pk_{i,\pi_{i}(j)},\textsf{Enc}_{pk_{i,\pi_{i}(j)}}(y_{i,\pi_{i}(j)})\}_{j\in[n_{i}]}$, for each group $i\in[m]$.
  6. 6.
    Every client downloads the list, finds in the list the entry with its own public key, and decrypts the payload to get the computation result.
V-B Security Analysis
This section presents the formal security analysis of the protocol in the previous section. In the following, we will consider a semi-honest adversary who can statically corrupt parties in the protocol. That is, the adversary will faithfully follow the protocol specifications but try to learn more information than allowed through protocol interaction. Also, before running the protocol, the adversary specifies the corrupted parties. The adversary controls the corrupted parties and knows their internal states. We use $\mathcal{C}$ to denote the collection of corrupted parties and $\mathcal{C}\subset G_{1}\cup G_{2}\cdots G_{n}\cup\{S\}$.
We first show that our protocol securely realizes the ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model. This means the protocol leaks no more information than what is allowed by $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ and $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$. More precisely, each client gets their output from $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$, the server gets $(L,f(L))$, and in the case of colluding with some clients in $G_{i}$, the partial permutation $\mathcal{L}(\pi_{i})$. Formally we have the following theorem:
Theorem V.1 (Security)
The Private Individual Computation (PIC) protocol in §V securely computes $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model in the presence of any PPT adversary with static corruption.
The proof can be found in Appendix -J. The proof is simulation-based. It shows that for any PPT adversary $\mathcal{A}$ in the real world, there exists a PPT simulator $\mathcal{S}$ in the ideal world that can generate a simulated view given the corrupted parties’ inputs and outputs. Security means that the simulated view is indistinguishable from the view of real-world execution.
The above theorem states that the adversary learns strictly no more than the allowed output and leakage by engaging in the protocol execution. Next, we will show how much differential privacy we can get in the presence of such an adversary with such knowledge. We consider an honest user $u_{\scriptscriptstyle i^{*},j^{*}}$, where $i^{*}\in[m],j^{*}\in[n_{i^{*}}]$. At the same time, the set of corrupted users in $G_{i^{*}}$ is denoted as $C_{i^{*}}\subset G_{i^{*}}$. Differential privacy in our case means that on two neighboring inputs $X=\big{(}X_{1},\cdots,X_{i^{*}}\cdots,X_{m}\big{)}$ and $X^{\prime}=\big{(}X_{1},\cdots,X^{\prime}_{i^{*}}\cdots,X_{m}\big{)}$, the output and leakage obtained by the adversary, denoted as $\mathcal{A}(X)$ and $\mathcal{A}(X^{\prime})$, are close in distribution. Formally, we have the following theorem:
Theorem V.2 (Differential Privacy)
The Private Individual Computation protocol satisfies $(\epsilon_{c},\delta)$-DP, i.e.
$\max(D_{\epsilon_{c}}(\mathcal{A}(X)||\mathcal{A}(X^{\prime})),D_{\epsilon_{c}}(\mathcal{A}(X^{\prime})||\mathcal{A}(X)))\leq\delta.$
In particular, when all users in $G_{i^{*}}$ use an identical $\epsilon$-LDP mechanism as the data randomizer $\mathcal{R}_{i^{*}}$, and for $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|\geq\tilde{\Omega}(e^{\epsilon})$ we have:
$\small\epsilon_{c}=\log\Bigg{(}1+\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\Big{(}\sqrt{\frac{32(e^{\epsilon}+1)\log 4/\delta}{n^{\prime}_{i^{*}}}}+\frac{4(e^{\epsilon}+1)}{n^{\prime}_{i^{*}}}\Big{)}\Bigg{)}.\normalsize$ (1)
The analysis can be divided into two cases: in the first case, $S\in\mathcal{C}$, i.e. S is corrupted. In this case, since the honest client locally randomizes their input, the input enjoys at least $\epsilon$-DP. Then the shuffling will amplify the privacy guarantee. Since the corrupted server receives the partial permutation as the leakage, the amplification depends on the number of uncorrupted users in the same group ($n^{\prime}_{i^{*}}$) as the honest user. The amplified $\epsilon_{c}$ can then be derived following [31, 32]. In the second case where $S\not\in\mathcal{C}$, the knowledge of the adversary is $y_{i,j}$ for each corrupted user $u_{i,j}\in\mathcal{C}$. The tricky part is that how much information $y_{i,j}$ leaks depends on the function $f$ being evaluated by the server. Hence we consider the worst case where $f$ output $y_{i,j}=L$. Continuing along the same line of thought, we conclude that the level of differential privacy assurance is no less than in the first case. The full proof can be found in Appendix -A.
V-C Discussion on Post-computation Communication
In certain scenarios, such as spatial crowdsourcing and location-based social systems, there may be additional client-to-client communications following the execution of the PIC protocol. For instance, consider a ridesharing application where passengers are in a group $G_{1}$ and taxi drivers in a group $G_{2}$. After receiving the matching result at the end of the protocol, the passengers must send their locations to the matched drivers, who need to know where to pick them up. On the other hand, the drivers also need to share their identities and locations with the matched passengers. Inevitably, a party has to sacrifice their privacy to the matched parties.
The post-computation communication may also have privacy implications for other clients not in the matched pair. Privacy amplification via shuffling against an adversary relies on the number of parties that remain anonymous. Recall that in Equation 1, the amplified $\epsilon_{c}$ relies on the number $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|$ of uncorrupted users in a particular group $G_{i^{*}}$. From the perspective of the adversary, if the post-computation communication compromises the anonymity of an additional set $U$ of the clients in $G_{i^{*}}$, then $n^{\prime}_{i^{*}}$ becomes $|G_{i^{*}}-C_{i^{*}}-U|$. Accordingly, the privacy amplification effect for users in $G_{i^{*}}$ that are still anonymous is weakened.
We emphasize that preventing the loss or weakening of privacy via technical means is not feasible, because the information is necessary for the proper functioning of the application. However, managerial countermeasures, such as ensuring sufficiently large user groups and limiting post-computation exposure according to the need-to-know principle, can mitigate potential privacy risks arising from post-computation communication.
VI Optimal Randomizers
VI-A Inadequacy of Existing Randomizers
In the PIC model, each user first sanitizes their data using an $\epsilon$-LDP randomizer. The design of the randomizer significantly impacts the utility of the tasks. While the PIC model can be seen as an extension of the shuffle model, it has unique characteristics that render the existing LDP randomizers commonly used in the shuffle model inadequate.
The main discrepancy between the shuffle model and the PIC model is that the former emphasizes statistical utility, whereas the latter focuses on the utility of each report. The shuffle model aims to estimate certain statistics from the noisy data collected from users, so it cares about how close the estimation is to the true value of the desired statistic. In the literature, utility is often measured by the expected square error (i.e., the variance) of the estimation:
$\mathbb{E}_{\scriptscriptstyle T\in\mathbb{X}^{n}}[\|\tilde{f}(T)-f(T)\|_{2}^{2}]=\textsf{Var}_{\scriptscriptstyle T\in\mathbb{X}^{n}}[\tilde{f}(T)]$
where $f$ is a statistical function, and $\tilde{f}$ is its estimation output by the shuffle protocol. On the other hand, in the PIC model, the tasks are often non-statistical. For example, in location-based matching, the required computation is to take two users’ locations and compute the distance between them. Hence, the above utility measure is no longer suitable. It is more natural to measure utility by the single report error:
$\max_{x_{j}\in\mathbb{X}}\mathbb{E}[\|\mathcal{R}(x_{j})-x_{j}\|_{2}^{2}],$
where $\mathcal{R}$ is the LDP-randomizer employed by the users. Additionally, in the PIC model, user data is typically multi-dimensional (e.g., location, gradient), making sanitization significantly more difficult compared to scalar data. Another notable characteristic of the PIC model is that the local differential privacy budget $\epsilon$ is relatively large, often scaling linearly with $\log(n^{\prime}_{i^{*}})$, as implied by Theorem V.2. These factors together create issues when existing LDP randomizers are applied directly in the PIC model.
To see the problem, we first examine a class of LDP randomizers [64, 84, 58, 50]) that work by sampling a few dimensions from $[d]$. Each user submits an incomplete report that contains only the sampled dimensions (with added noise) in their local data. This is an effective strategy for statistical estimation, but will totally ruin the utility of a single report in PIC as unsampled dimensions are missing. Clearly, they are unsuitable for use in the PIC model.
There have been LDP randomizers that submit complete reports. One obvious strategy is to explicitly split the local budget into $d$ parts and then apply a one-dimensional LDP mechanism independently to each dimension, or implicitly distribute the budget among dimensions, as seen in the Laplace [27], PlanarLaplace [4], PrivUnit [18], and PrivUnitG mechanisms [8]. However, these approaches are sub-optimal in the high budget regime. Specifically, even if we use the optimal one-dimension randomizer [9], splitting the budget across each dimension and then applying any randomizer for each dimension will result in a mean squared error (MSE) of at least $\frac{d}{(e^{\epsilon/d}-1)^{2/3}}$. The Laplace/PlanarLaplace mechanisms introduce an MSE rate of $\frac{d}{\epsilon^{2}}$, while the PrivUnit/PrivUnitG mechanisms incur an MSE rate of $\frac{d}{\min\{\epsilon,\epsilon^{2}\}}$. In contrast, later we will show that the MSE rate can be improved to $(e^{\epsilon}-1)^{-2/(d+2)}$ (see Section VI-C and Appendix -B).
Another strategy for submitting complete reports involves using additional randomization techniques such as random projection, data sketches, public randomness, or quantization. Randomizers employing this strategy [73, 33, 71, 19, 49, 6] avoid the issue of incomplete reports but introduce additional noise because of the extra randomization. While the additional noise is not significant in the low budget regime (i.e., $\epsilon=O(1)$) that is typical in the LDP model, it becomes dominant in the PIC model where the local budget can be as large as $\tilde{O}(\log n^{\prime}_{i^{*}})$. The resulting additional error will never diminish even when $\epsilon\to+\infty$.
VI-B Randomizer Design
We now introduce an asymptotic optimal randomizer, tailored for the PIC model. The randomizer uses an LDP mechanism, which we termed as Minkowski Response. For ease of presentation, here we will focus primarily on the $\ell_{2}$ case (and the $\ell_{+\infty}$ case in Appendix -B), but the mechanism can be generalized to other Minkowski distances.
Without loss of generality, we assume the user data domain to be a $\ell_{2}$-bounded hyperball $\mathbb{X}=\{x\ |\ x\in\mathbb{R}^{d}\ and\ \|x\|_{2}\leq 1\}$. Most, if not all, real-world data domains can be normalized to $\mathbb{X}$ (e.g. gradient vector, set-valued data, and location data). We also denote an $\ell_{2}$-bounded hyperball with radius $r$ centered at any $x\in\mathbb{R}^{d}$ as follows:
$\mathbb{B}_{r}(x)=\{x^{\prime}\ |\ x^{\prime}\in\mathbb{R}^{d}\ and\ \|x^{\prime}-x\|_{2}\leq r\},$
and it is shorted as $\mathbb{B}_{r}$ when $x=\vec{0}$.
Minkowski Response works by first defining a distance $r$ based on the local privacy budget as the following:
$r=\big{(}(e^{\epsilon}-1)^{1/(d+2)}-1\big{)}^{-1}.$
Then given the input domain $\mathbb{X}$, the output domain $\mathbb{Y}_{r}$ is defined by expanding $\mathbb{X}$ by $r$:
$\mathbb{Y}_{r}=\{y\ |\ y\in\mathbb{R}^{d}\ and\ \exists x\in\mathbb{X}\ that\ y\in\mathbb{B}_{r}(x)\}.$
Figure 4: Illustrative probability design of Minkowski response mechanism over two dimensions with cap area radius $r$, given different inputs $x_{i}$ in (a) and (b) respectively.
For any input $x\in\mathbb{X}$, Minkowski Response outputs an output $y\in\mathbb{Y}_{r}$ with relatively high probability in the cap area $\mathbb{B}_{r}(x)$ and relatively low probability in remaining domain $\mathbb{Y}_{r}\backslash\mathbb{B}_{r}(x)$ (see Figure 4). Formally:
$y=\left\{\begin{array}[]{@{}ll@{}}\text{uniform}(\mathbb{B}_{r}(x)),&\text{with prob. $\frac{\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}$;}\\ \text{uniform}(\mathbb{Y}_{r}),&\text{with prob. $\frac{\mathrm{V}(\mathbb{Y}_{r})}{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}$,}\\ \end{array}\right.$ (2)
where $V(*)$ denote the volume of the corresponding domain.
Lastly, $y$ is debiased to $\tilde{x}$ so that $\mathbb{E}[\tilde{x}]={x}$ as follows:
$\tilde{x}=y\cdot\frac{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}{V(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}.$ (3)
It is obvious that the output of Minkowski Response is informative in every dimension. Therefore it avoids problems brought up by incomplete reports. Also, intuitively it has a better utility because the mechanism is more likely to output a value near the true value (i.e., in $\mathbb{B}_{r}(x)$), than from other parts of the output domain (i.e. $\mathbb{Y}_{r}$). When the budget $\epsilon$ gets large, the error rate of Minkowski Response decays faster than in previous LDP mechanisms. More specifically, the decay rate of Minkowski Response is $(e^{\epsilon}-1)^{-2/(d+2)}$ (Equation 5 in Appendix -B), while that of the previous mechanisms is $d/(e^{\epsilon/d}-1)^{2/3}$ or $d/\epsilon^{2}$. Therefore, the utility advantage of Minkowski Response becomes more significant when $\epsilon$ gets larger. When $n^{\prime}_{i^{*}}\to+\infty$ (and thus $\epsilon\to+\infty$), $r$ becomes 0, and the error goes to zero (i.e. no additional error).
VI-C Analysis of Minkowski Response
The local privacy guarantee of the randomizer is presented in Theorem VI.1.
Theorem VI.1 (Local Privacy Guarantee)
Given input domain $\mathbb{X}=\mathbb{B}_{1}$, the Minkowski response mechanism defined in Equation 2 satisfies $\epsilon$-LDP.
Proof:
It is observed that the output probability distribution in Equation 2 is valid for any input $x\in\mathbb{X}$, the probability density in the cap area $\mathbb{B}_{r}(x)$ is $\frac{e^{\epsilon}}{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}$, and the density in the non-cap area $\mathbb{Y}_{r}\backslash\mathbb{B}_{r}(x)$ is $\frac{1}{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}$. Therefore, for any $x,x^{\prime}\in\mathbb{X}$, we have $\frac{\mathbb{P}[\mathcal{R}(x)=y]}{\mathbb{P}[\mathcal{R}(x^{\prime})=y]}\leq e^{\epsilon}$ for all possible $y\in\mathbb{Y}_{r}$, establishing the local $\epsilon$-DP guarantee of the Minkowski response mechanism $\mathcal{R}$. ∎
Next, we analyze the utility of the Minkowski Response in the PIC model. As previously mentioned, in the PIC model, the single report error is a more appropriate measure of utility compared to the statistical errors used in the conventional shuffle model. In Theorem VI.2, we examine the single report error in the PIC model and establish its lower bound. The proof is provided in Appendix -C.
Theorem VI.2 (Error Lower Bounds)
Given $d\in\mathbb{N}$, $\epsilon_{c}>0$, $\delta\in(0,0.5]$, $\mathbb{X}=\mathbb{B}_{1}(\{0\}^{d})$, then for any randomizer $\mathcal{R}:\mathbb{X}\mapsto\mathbb{R}^{d^{\prime}}$ such that $\mathcal{S}\circ\mathcal{R}(X)$ and $\mathcal{S}\circ\mathcal{R}(X^{\prime})$ are $(\epsilon_{c},\delta)$-indistinguishable for all possible neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, and for any estimator $f:\mathbb{R}^{d^{\prime}}\mapsto\mathbb{X}$, we have $\max_{x\in\mathbb{X}}\mathbb{E}\big{[}\|f\circ\mathcal{R}(x)-x\|_{2}^{2}\big{]}\geq\tilde{\Omega}\big{(}{1}/{n^{\frac{2}{d+2}}}\big{)}.$
The above theorem suggests that in the PIC model, for any randomizer, the single report error is at least $\tilde{\Omega}\big{(}{1}/{n^{\frac{2}{d+2}}}\big{)}$. Clearly, a randomizer offers better utility if its error is closer to this bound. For the randomizer using Minkowski Response, as presented in section VI-B, we can demonstrate that its single report error has an upper bound. This is summarized in Theorem VI.3, with the proof provided in Appendix -B.
Theorem VI.3 (Error Upper Bounds)
For any $d\in\mathbb{N}$, $\epsilon_{c}>0$, $\delta\in(0,0.5]$, $\mathbb{X}=\mathbb{B}_{1}(\{0\}^{d})$, if $\epsilon_{c}\leq O(1)$ and $n>\max\{16\log(1/\delta),\frac{2^{d+7}\log(1/\delta)}{(e^{\epsilon_{c}}-1)^{2}}\}$, then there exist a randomizer $\mathcal{R}:\mathbb{X}\mapsto\mathbb{R}^{d}$ such that $\mathcal{S}\circ\mathcal{R}(X)$ and $\mathcal{S}\circ\mathcal{R}(X^{\prime})$ are $(\epsilon_{c},\delta)$-indistinguishable for all possible neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, and:
$\max_{x\in\mathbb{X}}\mathbb{E}\big{[}\|\mathcal{R}(x)-x\|_{2}^{2}\big{]}\leq O\Bigg{(}\Big{(}\frac{\log(1/\delta)}{n\epsilon_{c}^{2}}\Big{)}^{\frac{2}{d+2}}\Bigg{)}.$
Given Theorems VI.2 and VI.3, we observe that as $n\rightarrow\infty$, the error upper bound $\tilde{O}(\frac{1}{n^{2/(d+2)}})$ of the Minkowski Response randomizer is the same as the error lower bound of the PIC model (the $\tilde{O}$ notation suppresses factors about $\epsilon_{c},\delta$ for simplicity). This means that the utility of the Minkowski Response is asymptotically optimal. In contrast, existing LDP randomizers with privacy amplification in the PIC model will suffer the single report error of $\tilde{O}(\frac{d}{n^{2/(3d)}})$ or $\tilde{O}(\frac{d}{\log^{2}n})$, whose denominator grows much slower than $n^{2/(d+2)}$ when $d>1$ and $n$ is relatively large. In practice, $n$ does not need to be very large; in our experiments, Minkowski Response outperforms existing LDP randomizers in PIC model once $n$ reaches the order of $10^{2}$. When comparing the PIC model to local DP model, the gap is undoubtedly larger, as numerical methods of privacy amplification [56, 85] will always give a local budget $\epsilon$ that is larger than the global privacy $\epsilon_{c}$ as long as $n\geq 2$, creating an extra gap in between the local and PIC models of these LDP randomizers. Actually, any randomizers in the LDP model must endures an error of $\Omega({d}/{\epsilon_{c}^{2}})$ for single report when $\epsilon_{c}\leq O(1)$ [18, 25].
VII Experimental Evaluation
We evaluate the utility and efficiency performances of proposed PIC model in three representative individual computation tasks: spatial crowdsourcing, location-based nearest neighbors, and federated learning with incentives, aim to answer the following research questions:
  • RQ1.
    How much do our proposals outperform existing approaches in accuracy/utility?
  • RQ2.
    What the computational/communication costs of our model for supporting security functionalities?
TABLE II: Mean $\ell_{2}$-error (and running time) comparison of local $\epsilon$-LDP randomizers on location domain $[-1,1]\times[-1,1]$. All results are the expected value of $1000$ repeated experiments.
randomizer $\epsilon=0.5$ $\epsilon=1.0$ $\epsilon=2.0$ $\epsilon=3.0$ $\epsilon=5.0$ $\epsilon=8.0$ $\epsilon=10.0$
PrivUnit [16] 8.94 (49us) 4.68 (42us) 2.25 (61us) 1.44 (63us) 0.81 (113us) 0.32 (305us) 0.18 (868us)
PrivUnitG[8] 8.73 (76ms) 4.63 (75ms) 2.27 (77ms) 1.51 (76ms) 0.96 (74ms) 0.63 (75ms) 0.53 (81ms)
Laplace [28] 12.97 (0.1us) 6.56 (0.1us) 3.27 (0.1us) 2.13 (0.1us) 1.30 (0.1us) 0.81 (0.1us) 0.64 (0.1us)
PlanarLaplace [4] 11.17 (12us) 5.63 (11us) 2.84 (11us) 1.88 (11us) 1.14 (11us) 0.71 (12us) 0.56 (12us)
Staircase[35] 13.19 (49us) 6.40 (48us) 3.13 (46us) 2.01 (48us) 1.05 (44us) 0.46 (51us) 0.28 (50us)
SquareWave[57] 11.87 (1.9us) 5.72 (2.5us) 2.65 (1.9us) 1.68 (2.2us) 0.92 (1.7us) 0.53 (2.2us) 0.42 (2.1us)
MinkowskiResponse 10.42 (0.7us) 4.50 (0.6us) 1.78 (0.8us) 0.98 (0.8us) 0.39 (0.6us) 0.14 (0.7us) 0.074 (0.8us)
VII-A Spatial Crowdsourcing
Datasets and settings. In the experiments for spatial crowdsourcing, we use use two real datasets, GMission dataset [20] and EverySender dataset [79], for scientific simulation and campus-based micro-task completion respectively. The number of users/workers, location domain range, and serving radius of workers about these two datasets are summarized in Table III (more information are presented in Appendix -F). In our experiments, for compliance with data randomizers (see following parts), we rescale the location domain range to $[-1,1]\times[-1,1]$ and rescale the corresponding serving radius to $0.41.0\cdot\frac{1-(-1)}{5-0}=$ 0.4.
TABLE III: Summary statistics of spatial crowdsourcing datasets.
Dataset users workers location domain serving radius
GMission 713 532 $[0,5.0]\times[0,5.0]$ $1.0$
EverySender 4036 817 $[0,5.0]\times[0,5.0]$ $1.0$
Cryptography implementation. Follow the protocol in Section V, we implement the PIC model using Elliptical Curve Diffie-Hellman key agreement (ECDHE [11]) for establishing shared secret between client/server and matched user/worker, meanwhile using AES encryption [68] and HMAC signature [14] for secure point-to-point communication after key agreement.
Data randomizers and server-side algorithms. We use the Minkowski response for location randomization. As comparison, in the local model of DP (e.g., in [83, 77, 81]), we compare with existing mechanisms including Laplace [28], Staircase [35], PlanarLaplace with geo-indistinguishability [4], SquareWave [57], PrivUnit [16] and its Gaussian variant PrivUnitG in [8]. On the server side, we use either minimum weighted full matching [51] or maximum matching [45] over bipartite graph as the user/worker matching algorithm based on (noisy) locations, to obtain a match $M:G_{a}\times G_{b}\mapsto\mathbb{R}$. Specifically, the minimum weighted full matching aims to minimize the overall travelling costs between users and works over the (noisy) fully-connected bipartite graph; the maximum matching aims to maximize the number of successfully matched user/worker pairs over the (noisy) partially-connected bipartite graph, where user/worker pairs with distance surpassing $0.4$ (i.e. the serving radius) are deemed as not connected.
Performance metrics. We evaluate the performance of spatial crowdsourcing in terms of travel costs, successful matching ratio, and running time. The travel costs are the total actual Euclidean distance between all pairing matched user/workers:
$\text{travel costs}=\sum\nolimits_{(i,j)\in G_{a}\times G_{b}}\llbracket M(i,j)>0\rrbracket\cdot\|l_{i}-l_{j}\|_{2},$
[…]
$\text{success ratio}=\frac{\sum_{(i,j)\in G_{a}\times G_{b}}\llbracket M(i,j)>0\rrbracket\cdot\llbracket\|l_{i}-l_{j}\|_{2}\leq\tau\rrbracket}{\min\{|G_{a}|,|G_{b}|\}}.$
We evaluate both the client-side and server-side running time of our protocol on a laptop computer embedded with Intel i5555-8250U CPU @1.6GHz and $8$GB memory.
Figure 5: Expected $\ell_{2}$ error of reported locations.
Performance of location reports. Before exploring spatial crowdsourcing tasks, we first evaluate our proposed Minkowski response mechanism against existing mechanisms, as presented in Table II. This evaluation focuses on errors and running times associated with reporting locations under local $\epsilon$-DP constraints. We quantify the error using the expected $\ell_{2}$ distance between each true and noisy location pair. Notably, the Minkowski response showcases an $\ell_{2}$ error comparable to state-of-the-art (SOTA) mechanisms when $\epsilon\leq 1.0$. Its superiority is evident when $\epsilon\geq 2.0$, displaying a significantly reduced error. This supports the theoretical assertions made in Section VI, underscoring the importance of the Minkowski response for the shuffle model, over other existing randomizers. Additionally, the Minkowski response is highly efficient, demanding less than $1$us on the user end. This speed is nearly equivalent to merely adding Laplace noises and is roughly $50$x faster than the previously SOTA PrivUnit.
We further contrast the utility performances of the mechanisms in the local model with those in the shuffle model. Privacy amplification in the shuffle model is achieved using the latest analyses from [32]. As illustrated in Figure 5, the shuffle model presents a significantly reduced location reporting error compared to the local model. This disparity becomes more pronounced as the number of users increases (notably, the EverySender dataset comprises more users than GMission). The introduction of an apt randomizer—such as the Minkowski response—can further elevate utility. This is evident when comparing the Minkowski response and PrivUnit, both of which operate within the shuffle model.
Performance of minimum weight matching. Figure 6 presents the expected travel costs associated with minimum weighted full matching. For a comprehensive overview of the experimental results involving more competitors, readers are directed to -F. Notably, the Squarewave mechanism outperforms PrivUnit in matchings, prompting its primary representation in the shuffle model discussions. The equivalent shuffle model yields a cost reduction of approximately $50\%$, with the savings escalating as the number of users increases. The Minkowski response consistently exhibits the lowest costs across all scenarios. Together, they result in a substantial reduction of about $75\%$ when compared to existing approaches.
Figure 6: Travel costs of minimum weighted matching.
Performance of maximum matching. Figure 7 illustrates the success ratios for maximum matching. The equivalent shuffle model shows a relative improvement of approximately $200\%$ in the success ratio. The Minkowski response consistently yields higher ratios across all scenarios. When the user count becomes relatively substantial (for instance, a few thousands as in EverySender), our results indicate that spatial crowdsourcing maximum matching, with a reasonable level of privacy protection (specifically, $\epsilon\leq 3$), can be made practical, achieving a success ratio exceeding $90\%$.
Figure 7: Success ratios of maximum matching.
TABLE IV: Running time and communication overheads of spatial crowdsourcing in the PIC model.
User-side Procedures Time Communication Server-side Procedures Time Communication
location randomization 0.7us - decrypt one message 2.1ms -
encrypt information 2.9ms - min-weight match (GMission) 31ms 1.2MB
send message to shuffler - 1.3KB maximum match (GMission) 15ms 1.2MB
retrieve matching result - 1.8KB min-weight match (EverySender) 79ms 4.3MB
key agreement with worker 2.6ms - maximum match (EverySender) 63ms 4.3MB
Running time and communication overheads. We present the computation and communication overheads of the PIC model in Table IV. When considering multiple procedures on the user side, the total running time for each user is only a few milliseconds (with the data randomization time being negligible), and the total communication overhead is under $4$KB. On the server side, the running time of the matching algorithm on noisy plaintext is consistent with non-private settings, scaling with the number of users/workers and taking dozens of milliseconds. The additional decryption/encryption runtime is capped at several seconds, mirroring the costs of HTTPS decryption/encryption. Specifically, it utilizes ECDHE for key agreement and AES encryption as with HTTPS on TLS 1.3 [67] for client-server communication. In essence, the running times on both the client and server sides of the PIC model closely resemble those in non-private settings that use HTTPS communication (for a fair comparison).
VII-B Location-based Social Systems
In this part, we evaluate the performance of PIC model for privacy preserving nearest neighbor (NN) queries on users, a prevalent task in location-based social networks [22]. Like the exemplar spatial crowdsourcing tasks, the nearest location senders’ information (including their corresponding public keys) w.r.t. $\ell_{2}$-distance) of a user is securely return to the user. These matched neighboring users can then securely communicate with other to accomplish subsequent tasks.
TABLE V: Statistics of location social network datasets.
Dataset users check-ins location domain
Gowalla 138368 6442892 $[37.54,37.79]\times[-122.51,-122.38]$
Foursquare 1083 227428 $[40.55,40.99]\times[-74.27,-73.68]$
Datasets and settings. We use use two real check-in datasets, Gowalla dataset [22] and Foursquare dataset [87]. The statistics of two datasets are summarized in Table V. The location domains are rescaled to the standard domain $[-1,1]\times[-1,1]$. As for nearest neighbor queries, we set the nearest radius $\tau$ for Gowalla to $0.1$ and the nearest radius $\tau$ for Foursquare to $0.2$, so that users have a few tens or hundreds neighbors (vary due to check-ins densities). Accordingly, the population size for privacy amplification in PIC model is set to $n-1000$. The concrete cryptography implementations and data randomizers for nearest neighbor queries are the same as the one for spatial crowdsourcing, while the server simply runs a radius-based nearest neighbor algorithm and return the neighboring (public key, noisy location) lists to each user.
Performance metrics. We evaluate the performance of nearest neighboring queries in terms of precision and recall, and $F_{1}$ score of neighboring lists. Let $N_{i}$ denote the true neighbor lists that have $\ell_{2}$ distance to user $i$ less than the parameter $\tau$, and let $\widehat{N}_{i}$ denote the retrieved neighbor lists that have $\ell_{2}$ distance to user $i$ based on noisy (sanitized) locations less than the parameter $\tau$. The precision of nearest neighboring queries is defined as $\text{precision}={(\sum_{i\in[n]}|N_{i}\bigcap\widehat{N}_{i}|)}/{(\sum_{i\in[n]}|N_{i}|)}$, the recall of nearest neighboring queries is $\text{recall}={(\sum_{i\in[n]}|N_{i}\bigcap\widehat{N}_{i}|)}/{(\sum_{i\in[n]}|\widehat{N}_{i}|)}$. We use a balanced indicate
$F_{1}\text{ score}=2\cdot\text{precision}\cdot\text{recall}/(\text{precision}+\text{recall})$
as the major measurement. All presented results are the average value of $10$ experimental runs.
Experimental results. Figure 8 displays the $F_{1}$ scores for radius nearest neighbor queries. To evaluate the effects of varying user numbers, we adjusted the number of users $n$ in the Gowalla dataset from $10000$ to $40000$. Notably, as $n$ surpasses $50000$, the PIC model’s scores see an approximate relative increase of $100\%$ compared to the local model. The Minkowski response consistently records the highest scores. When used in conjunction, they can achieve up to $90\%$ of the $F_{1}$ scores within a reasonable privacy loss (e.g., $\epsilon\leq 3$), especially when $n$ is relatively large.
Figure 8: F1 scores of radius nearest neighbor queries in location-based social systems.
VII-C Federated Learning with Incentives
Recall that in federated learning with rigorous privacy protection, many clients collaboratively contribute gradient updates of model parameters so as to build a machine learning model. Each client $i$ holds a local dataset $D_{i}$. In each iteration $t$, a client $i$ downloads the latest model parameter $W_{t-1}$, and computes the local gradient $g_{i}$ based on local dataset $D_{i}$. Then, the client sanitizes $g_{i}$ and sends it to the shuffler/server. Without loss of generality, we assume the server uses Shapley value [69] to determine the gradient payoffs, where the utility function for computing Shapley value is based on the model parameters’ performance (e.g., accuracy or the negative of loss function) on a validation dataset $D_{val}$. The Shapley values themselves may severely leak sensitive information about user data [59]. Despite the vast non-linear computations involved in evaluating Shapley values with neural networks, no rigorous protection has been offered for federated learning with incentives in existing literature. In this study, we demonstrate the feasibility of computing Shapley value without incurring additional privacy loss.
TABLE VI: Model architecture with 4292 trainable parameters.
Layer Parameters
Convolution 8 filters of 4$\times$4, stride 2
Max-pooling 2$\times$2
Convolution 6 filters of 5$\times$5, stride 2
Max-pooling 2$\times$2
Softmax 10 units
Datasets and settings. Our experiments with federated learning use the MNIST handwritten digit dataset, comprising $60,000$ training, $5,000$ validation, and $5,000$ test images. We employ a neural network outlined in Table VI with $d=4292$ trainable parameters. It achieves $98.6\%$ accuracy using non-private, untruncated vanilla SGD and $92.7\%$ accuracy with non-private, truncated SGD. For our experiments, we simulate $50,000$ clients with each client possessing one sample (i.e., $n=50,000$). During each private SGD step, $s=10,000$ clients are randomly selected. Each client subsamples $0.5\%$ of the gradient vector dimensions, then clips the $\ell_{1}$-norm of the block gradient with a threshold of $c=0.001$. We compare our PIC model against a local model, both implementing Laplace mechanism for gradient randomization.
Implementation details. We use Elliptical curve encryption to perform secure point-to-point information delivery. Specifically, to delivering monetary incentives with anonymity protection, clients and server may employ untraceble crypto-currency (e.g., ZCash [46]). Each client creates one random temporary wallet address and append it to the message, the server computes incentive payments of the message and then send corresponding amount of coins to the wallet address.
Performance metrics. We assess performances using gradient/model accuracy and Shapley value accuracy. Gradient accuracy is gauged using the mean squared error (MSE) between the subsampled, clipped, and privatized gradients $\hat{g}_{i}$ and the corresponding non-private gradients ${g}_{i}$:
$\text{gradient}\ MSE=\|\sum\nolimits_{i\in[s]}\hat{g}_{i}-\sum\nolimits_{i\in[s]}{g}_{i}\|_{2}^{2}.$
Model accuracy refers to the image classification accuracy on the test dataset. Shapley value accuracy is the MSE between Shapley values of privatized gradients $\hat{g}_{i}$ and non-private gradients ${g}_{i}$:
$\text{shapley}\ MSE=\sum\nolimits_{i\in[s]}(\widehat{\texttt{Shapley}}_{i}-{\texttt{Shapley}}_{i})^{2},$
where the Shapley values for privatized gradients use the utility function (see Appendix -H for detail):
$\widehat{U}(S)=\frac{<\texttt{grad}_{val},\sum_{i\in S}\hat{g}_{i}>}{\|\texttt{grad}_{val}\|_{2}\cdot\|\sum_{i\in S}\hat{g}_{i}\|_{2}}.$
The presented gradient/shapley MSE averages over $50$ aggregation rounds, and the model accuracy is from the final round.
Experimental results. Since the curator model of DP and other cyrptographic MPC approaches (e.g., secure aggregation [18]) are not applicable to privacy preserving joint gradient aggregation and Shapley value computation (with complicate non-linear computation), we compare equivalent shuffle model with the local model of DP. Both models use Laplace noises [28] to privatizing local gradients, and use tight numerical composition for sequential gradient averaging [42]. We present the experimental results in Table VII, with varying overall privacy budget $\epsilon$. Every result is the average value of $5$ simulations.
TABLE VII: Utility performance of joint gradient aggregation and Shapley-value incentive for federated learning with differential privacy guarantee on MNIST dataset.
Setting
Privacy
Budget
Test
Accuracy
Gradient
MSE
Shapley
MSE
local model $\epsilon=1$ $19.1\%$ $1.7$e-4 $8.9$e-3
$\epsilon=3$ $28.9\%$ $1.3$e-4 $6.1$e-3
$\epsilon=5$ $35.6\%$ $1.1$e-4 $5.2$e-3
PIC model $\epsilon=1$ $77.6\%$ $2.5$e-5 $4.6$e-3
$\epsilon=3$ $81.2\%$ $1.4$e-5 $4.2$e-3
$\epsilon=5$ $85.1\%$ $1.1$e-5 $3.9$e-3
It is observed that PIC model enjoys significantly better gradient accuracy, increase model test accuracy by about $50\%$ in the absolute sense, reduces about $90\%$ gradient aggregation error, when compared to the local model. The PIC model also reduces about $40\%$ error on Shapley values, permit accurate payoff distribution to incentive honest/informative participants.
VIII Conclusion
In this paper, we proposed private individual computation model, for multi-party individual computation with DP preservation, having broad application to spatial crowdsourcing, resource allocation, location-based systems, and federated learning with incentives. The paradigm introduces statistically random identities into the traditional shuffle model of DP, so as to support secure functionalities required by individual computations, meanwhile maintaining the most benefits of privacy amplification via shuffling. It incurs only $O(d)$ computation and communication costs (matching the non-private setting) on the client side, meanwhile permitting plaintext computation on the server side. Additionally, with the increasing of number of participating parties, the utility of individual computation grows (as there are stronger privacy amplification effects), making it suitable for large-scale applications.
The simple paradigm can efficiently solves prevalent tasks that are computationally intractable in MPC approaches (e.g., Ishai’s anonymous model, secure aggregation, or garbled circuits), such as federated deep learning with incentives and large-scale spatial crowdsourcing, necessitating essentially the same overheads as in non-private settings. Moreover, it reduces $40\%$-$90\%$ errors/costs, increases $100\%$-$300\%$ utility, when compared to existing tractable approaches (e.g., local DP).
Discussions and future research. The PIC model guarantees privacy gracefully in cases of malicious clients or collusion, and can adapt to fully decentralized settings without an orchestrating server. Given its alignment with random digital identities in Blockchains and in Ad tracking systems [54], it offers potential to enhance formal privacy in these systems.

### v2

@@
-Beyond Statistical Estimation: Differentially Private Individual Computation in the Shuffle Model
+Beyond Statistical Estimation: Differentially Private Individual Computation via Shuffling
 Abstract
-The shuffle model of differential privacy (DP) has recently emerged as a powerful one for decentralized computation without fully trustable parties.
-Since it anonymizes and permutes messages from clients through a shuffler, the privacy can be amplified and utility can be improved.
-However, the shuffling procedure in turn restricts its applications only to statistical tasks that are permutation-invariant.
-This work explores the feasibility of shuffle privacy amplification for prevalent non-statistical computations: spatial crowdsourcing, combinatorial optimization, location-based social systems, and federated learning with incentives, which suffer either computationally intractability or intolerable utility loss in existing approaches (e.g., secure MPC and local DP).
-We proposes a new paradigm of shuffle model that can provide critical security functionalities like message authorization and result access control pertained to permutation-equivariant tasks, meanwhile maintaining the most of privacy amplification effects.
-It incurs almost the same computation/communication costs on the client side as the non-private setting, and permits the server to run arbitrary algorithms on (noisy) client information in plaintext.
-Our novel technique is introducing statistically random identity into DP and force identical random distribution on all clients, so as to support secure functionalities even after message shuffling and to maintain privacy amplification simultaneously.
-Given that existing DP randomizers fails in the new shuffle model, we also propose a new mechanism and prove its optimality therein.
-Experimental results on spatial crowdsourcing, location-based social system, and federated learning with incentives, show that our paradigm and mechanism is fast as non-private settings, while reducing up to $90\%$ error and increasing utility performance indicates by $100\%$-$300\%$ relatively, and can be practical under reasonable privacy budget.
+In data-driven applications, preserving user privacy while enabling valuable computations remains a critical challenge.
+Technologies like Differential Privacy (DP) have been pivotal in addressing these concerns.
+The shuffle model of DP requires no trusted curators and can achieve high utility by leveraging the privacy amplification effect yielded from shuffling.
+These benefits have led to significant interest in the shuffle model.
+However, the computation tasks in the shuffle model are limited to statistical estimation, making the shuffle model inapplicable to real-world scenarios in which each user requires a personalized output.
+This paper introduces a novel paradigm termed Private Individual Computation (PIC), expanding the shuffle model to support a broader range of permutation-equivariant computations.
+PIC enables personalized outputs while preserving privacy, and enjoys privacy amplification through shuffling.
+We propose a concrete protocol that realizes PIC.
+By using one-time public keys, our protocol enables users to receive their outputs without compromising anonymity, which is essential for privacy amplification.
+Additionally, we present an optimal randomizer, the Minkowski Response, designed for the PIC model to enhance utility.
+We formally prove the security and privacy properties of the PIC protocol.
+Theoretical analysis and empirical evaluations demonstrate PIC’s capability in handling non-statistical computation tasks, and the efficacy of PIC and the Minkowski randomizer in achieving superior utility compared to existing solutions.
 Index Terms:
-differential privacy, shuffle model, digital identities, spatial crowdsourcing, federated learning
+Differential Privacy, Private Individual Computation, Shuffle Model, Privacy-Utility Trade-offs
 I Introduction
 Personal information fuels a wide array of data-driven applications, e.g.
-statistical analytics, machine learning, recommendation systems, spatial crowdsourcing, mobile crowdsensing, combinatorial optimization, and location-based social systems.
-These applications deliver substantial value but depend on data collected from users, which is a prime target for attacks and carries a high risk of leakage or abuse.
+statistical analytics, machine learning, recommendation systems, spatial crowdsourcing, e-health, social networks, and smart cities.
+These applications deliver substantial value but rely on data collected from users, which is a prime target for attacks and carries a high risk of leakage or abuse.
 Data privacy concerns are escalating, especially after several high-profile data breach incidents.
-Despite the introduction of stricter privacy laws such as the EU’s General Data Protection Regulation, the California Consumer Privacy Act, and China’s Personal Information Protection Law, many users still distrust service providers and are hesitant to consent to the use of their data.
-To bridge this trust gap and encourage user participation, significant efforts have been made recently to develop privacy-enhancing technologies that enable private data processing without relying on a trusted third party.
+Despite the introduction of stricter privacy laws such as the EU’s General Data Protection Regulation and the California Consumer Privacy Act, many users still distrust service providers and are hesitant to consent to the use of their data.
+To bridge this trust gap and encourage user participation, significant efforts have been made recently to develop privacy-enhancing technologies that enable private data processing.
 Two prominent technologies addressing this problem are secure multiparty computation (MPC) [91] and differential privacy (DP) [31].
-MPC uses interactive cryptographic protocols to allow mutually untrusted parties to jointly evaluate a function on their private data, with each party receiving an output but nothing else.
-However, the cryptographic nature of MPC imposes heavy overheads, making it challenging to scale for practical use.
-DP, on the other hand, adds random noise to data to ensure privacy, making it computationally more feasible since the computation is carried out on sanitized data without heavy cryptographic machinery.
-Traditional central DP still requires a trusted curator [28], but the local model (LDP [53]) allows each user to sanitize their data locally before sending it out.
-Due to its practicality and low trust assumption, LDP has been adopted by companies like Apple [74], Microsoft [24], and Google [30] in their real-world systems.
-However, LDP’s drawback is that each user must independently inject sufficient random noise to guarantee privacy, leading to significant utility loss.
+MPC employs interactive cryptographic protocols that enable mutually untrusted parties to jointly compute a function on their private data, ensuring each party receives an output while learning nothing else.
+Despite its strong privacy guarantees, the cryptographic nature of MPC results in substantial overhead, posing scalability challenges for practical applications.
+In contrast, DP enhances privacy by adding random noise to the data, allowing computation to be performed on sanitized data without the need for heavy cryptographic machinery.
+Traditional central DP relies on a trusted curator [31], but the local model (LDP [54]) enables users to sanitize their data locally before sharing it.
+LDP’s practicality and minimal trust requirements have led to its adoption by companies such as Apple [77], Microsoft [27], and Google [33] in their real-world systems.
+However, LDP’s primary drawback is that each user must independently add sufficient random noise to ensure privacy, which can significantly impact the utility of the data.
 Recently, within the realm of DP research, the shuffle model [13, 32] has emerged.
-This model introduces a shuffler that randomly permutes messages from users and then sends these anonymized messages to the computing server or analyzer.
+This model introduces a shuffler that randomly permutes messages from users and then sends these anonymized messages to a computing server or analyzer.
 The server can then compute on these messages to derive the result.
@@
 For instance, shuffled messages from $n$ users each adopting local $\epsilon$-DP actually preserve differential privacy at the level $\epsilon_{c}=\tilde{O}(\sqrt{e^{\epsilon}/n})$ [34, 35].
-Consequently, to achieve a predefined global privacy goal, each user consumes less privacy budget when sanitizing their data locally, meaning less noise needs to be added.
+Consequently, to achieve a predefined global privacy goal, less noise must be added when users sanitize their data locally.
 This significantly improves the accuracy of the final result.
@@
 The amount of rewarding incentive must be computed for individual users based on their contribution (e.g., via Shapley values [71]).
-At first glance, private individual computation seems unattainable within the shuffle model because the need for individualized output clashes with the anonymization required for privacy amplification.
-However, this is not necessarily true.
-Our observation is that many individual computing functionalities are equivariant to shuffling.
-That is, the permutation applied to the inputs does not affect the computation.
-This means shuffling does not prevent the server from producing personalized answers for each client—the server does not know and does not need to know which answer is for whom.
-Nevertheless, there is a problem: how to pass the output back to the right user without breaking anonymity.
-A straightforward approach is to let the shuffler maintain a long-term duplex connection channel between each client and the server (e.g., as in an Onion routing network [41]), but this method is costly (due to storing communication states) and may be subject to de-anonymization attacks on anonymous channels [63, 65].
-Additionally, the duplex connection is not supported in current shuffler implementations, such as the one in the seminal work [17].
-This is the first technical challenge we need to address.
-The second challenge we face is how to design optimal randomizers for individual computation in the shuffle model.
-Often the randomizer in a shuffle model protocol needs to be tailored to the tasks.
-For statistical tasks within the shuffle model, several studies have developed near-optimal randomizers, as seen in histogram estimation [36, 31] and one-dimensional summation estimation [9].
-However, the new setting of individual computation is different: the focus is on the accuracy of the output for each user (instead of the statistical accuracy of the population).
-This gap renders existing randomization strategies (e.g., randomizers utilizing dimension sampling, budget splitting, or data sketching, as reviewed in [86]), along with prevalent randomizers (such as adding Laplace [28] or Staircase [35] noises), less effective for the new setting.
-Therefore we need to take a fresh look at the fundamental privacy-utility trade-offs and re-design the underlying randomizers.
-In this work, we introduce a new paradigm of the shuffle model that provides security functionalities (e.g., message authorization, integrity verification, and access control) for permutation-equivariant computations, and term it the private individual computation model.
-The properties of the model are as follows.
-  • I.
-    Security: The server and other clients can verify the owner of a message using a fingerprint submitted alongside the message.
-The server can also securely deliver personalized backward information to each owner (i.e., access control).
-These functionalities enable the secure completion of equivariant tasks.
-  • II.
-    Privacy: There is no degradation of local privacy while providing these additional security functionalities.
-Furthermore, when task results are sparse (e.g., each user is matched to one taxi driver in spatial crowdsourcing), the degradation of shuffle privacy amplification is negligible.
-Additionally, our local and shuffle privacy guarantees are information-theoretic (as long as the shuffling is statistically random).
-This distinguishes our new model from computational DP [62], which also combines cryptographic functions and privacy but guarantees differential privacy only against polynomial-power adversaries.
-  • III.
-    Utility: Since clients still benefit from privacy amplification via shuffling, the utility is significantly improved.
-With the increasing of number of clients, the utility increases and approaches to non-private settings.
-In addition, the server in our model can run arbitrary algorithms on the (noisy) messages in plaintext.
-Our model is compatible with existing achievements and future advancements in server-side algorithms (e.g., noisy-aware matching algorithms for spatial crowdsourcing).
-  • IV.
-    Scalability: As the server runs algorithms on plain-text messages and there is no heavy computation burden on the client side, our model is highly efficient and scales to systems with massive clients.
-This eliminates the need for secure multi-party computation which necessitates impractical costs.
-In our model, the computation/communication costs are essentially the same to non-private settings (see Appendix -I).
-Main techniques.
-Our approach to realizing additional security functionalities is based on two critical observations: (1) If the message payload satisfies $\epsilon$-LDP, appending arbitrary random variables (independent to the private value) to the payload incurs no additional privacy loss.
-Specifically, when random variables are sampled from an identical distribution for all clients, there is no excess loss in privacy amplification via shuffling.
-(2) Post-processing the message payload and appended random variables does not incur excess privacy loss.
-The first observation motivates us to append a random fingerprint (e.g., a public key) to the payload, allowing the client to use the corresponding secret key to negotiate shared secret and establish secure communication with other parties in later phases.
-A side consequence is that the server can use the public key pertained to each message to securely deliver information to the message owner.
-The second observation motivates us to further append a signature to protect the integrity of the fingerprint and payload.
-Challenge (2): Optimal randomizers for non-statistical computations in the shuffle model.
-In the local model of DP, existing research [83, 66, 77, 81, 82] primarily focuses on designing randomizers in a high privacy regime (e.g., local budget $\epsilon=O(1)$).
-Conversely, in the shuffle model, the local budget can extend to $\tilde{O}(\log(n))$.
-For statistical tasks within this model, several studies have developed near-optimal randomizers, as seen in histogram estimation [36, 31] and one-dimensional summation estimation [9].
-However, the new setting of non-statistical computation tasks in the shuffle model—often involving multi-dimensional user data and emphasizing an individual report’s accuracy (instead of statistical accuracy of the population)—has not been explored.
-This gap renders existing randomization strategies (e.g., randomizers utilizing dimension sampling, budget splitting, or data sketching, as reviewed in [86]), along with prevalent randomizers (such as adding Laplace [28] or Staircase [35] noises), less effective for non-statistical computation in the shuffle model.
-The fundamental privacy-utility trade-offs in non-statistical computation for this new model also present significant theoretical interest.
-Our work establishes the error lower bounds for single reports in non-statistical computation within the shuffle model, demonstrating the sub-optimality of existing randomizers.
-To address this, we introduce a novel series of local randomizers—the Minkowski response—aligned with these error lower bounds.
-The core principle involves reporting values within a calibrated Minkowski distance from the true value with high probability.
-The main contributions of this work are summarized as follows:
-  • We propose equivariant shuffle model of DP, which supports various security functionalities and benefit large-scale equivariant computations from privacy amplification.
-The new model exhibits almost the same computation/communication costs on client/server side as non-private settings.
-  • We establish tight minimax error lower bounds of $\tilde{O}((n\epsilon^{2}_{c})^{-\frac{2}{d+2}})$ for the new model of DP, where $d$ is the user data dimension, $\epsilon_{c}$ is global privacy budget, and $n$ is the number of users.
-Since existing data randomizers fail, we then design an attainable randomizer: Minkowski response, for the new model to match the lower bound.
-  • Through experiments on three typical permutation-equivariant tasks: privacy-preserving spatial crowdsourcing, location-based nearest neighbor queries, and federated learning with incentives, we demonstrate that the new model, in conjunction with proposed randomizers, reduces error by $90\%$ meanwhile incurring negligible excess burdens in both the client and server side.
+At first glance, private individual computation appears unattainable within the shuffle model because the need for personalized output conflicts with the anonymization required for privacy amplification.
+However, this is not necessarily the case.
+Our observation is that many individual computing tasks are equivariant to shuffling, meaning that the permutation applied to the inputs does not affect the computation.
+Therefore, shuffling does not prevent the server from producing personalized answers for each client – the server does not need to know which answer belongs to whom.
+Nevertheless, there is a challenge: how to return the output to the correct user without compromising anonymity.
+One straightforward approach is to have the shuffler maintain a long-term duplex connection channel between each client and the server (e.g., as in an Onion routing network [44]).
+However, this method is costly due to the need to store communication states and may be vulnerable to de-anonymization attacks on anonymous channels [64, 66].
+Additionally, current shuffler implementations, e.g.
+the one described in the seminal work [13], do not support such duplex connections.
+Moreover, after receiving the computation results, many individual computation tasks require establishing party-to-party communication (e.g., a user communicates with the matched driver in taxi-hailing services, a user communicates with matched near users in social systems), which is intractable in the duplex shuffle channel.
+Addressing this issue is the first technical challenge we need to overcome.
+The second challenge we face is designing optimal randomizers for individual computation within the shuffle model.
+Often, the randomizer in a shuffle model protocol should be tailored to specific tasks.
+For statistical tasks within the shuffle model, several studies have developed near-optimal randomizers, as demonstrated in histogram estimation [39, 34] and one-dimensional summation estimation [7].
+However, the new setting of individual computation is different: the focus is on the accuracy of the output for each user rather than the statistical accuracy of the population.
+This difference renders existing randomization strategies (e.g., randomizers utilizing dimension sampling, budget splitting, or data sketching, as reviewed in [88]), as well as prevalent randomizers (such as adding Laplace [31] or Staircase [38] noises), less effective for the new setting.
+Therefore, we need to reconsider the fundamental privacy-utility trade-offs and redesign the underlying randomizers to better suit individual computation requirements.
+Our contributions In this work, we introduce a new paradigm extending the shuffle model that allows a wider range of permutation-equivariant tasks to be computed with DP guarantee and can enjoy privacy amplification provided by shuffling.
+We term the new paradigm Private Individual Computation (PIC in short).
+We define PIC formally as an ideal functionality, which captures its functional and security properties.
+We also provide a concrete protocol, with formal proof, that can realize the ideal functionality.
+Similar to the shuffle-DP protocols, each user adds noise locally to their data and encapsulates it (and possibly other auxiliary data) into an encrypted message under the public key of the computation server.
+The shuffler then shuffles the encrypted messages, before sending them to the computation server.
+The server can decrypt these messages and perform the permutation-equivariant computation.
+What differs is that each user also includes a one-time public key within the encrypted message.
+This one-time public key serves two purposes: (1) it allows the server to encrypt the computation result such that it can only be decrypted by the owner of the corresponding private key, and (2) it acts as a pseudonym for the key owner.
+This approach addresses our first challenge with minimal overhead: the server can publish a list where each entry consists of a public key along with the computation result encrypted under this key.
+Users can download the entire list and decrypt the entry associated with their own public key, maintaining their anonymity.
+Additionally, users can establish secure communication channels with other matched parties using their public keys to eventually complete PIC tasks.
+Another main contribution is the development of an asymptotically optimal randomizer specifically designed for the PIC model.
+This randomizer is based on an LDP mechanism we call Minkowski Response.
+The primary goal of this design is to enhance utility, measured by the single-report error, which is the expected squared error between a user’s true data value and its sanitized version.
+To achieve this, a Minkowski distance $r$ is determined based on the privacy budget.
+The randomizer’s output domain is defined as a ball extending the input domain’s radius by $r$.
+The key to achieving high utility lies in the randomizer’s output: it selects a value close to the true value (within $r$) with a relatively high probability and a value from elsewhere in the entire output domain with a relatively low probability.
+We formally prove the security and privacy properties of the PIC protocol.
+Additionally, we provide a theoretical analysis of the utility bounds achievable by protocols in the PIC model and the Minkowski Response mechanism.
+Our analysis demonstrates that asymptotically, the error upper bound of the Minkowski Response matches the error lower bound for all possible randomizers in the PIC model, thereby achieving optimality.
+Alongside theoretical analysis, we conducted extensive experiments using real-world applications and datasets.
+The evaluation confirms that computations conducted in the PIC model exhibit significantly better utility than those in the LDP model.
+Furthermore, the performance of the Minkowski randomizer, measured by single-report error and task-specific utility metrics, surpasses that of existing LDP randomizers commonly used in the shuffle model.
 Organization.
@@
 Section IV formalizes the problem setting.
-Section V presents the PIC model.
-Section VI provides optimal LDP randomizers.
+Section V presents the PIC protocol.
+Section VI provides optimal user-side randomizers.
 Section VII evaluates the utility and efficiency performances of our proposals.
@@
 II Related Works
-This section reviews various approaches to private computation, primarily concentrating on permutation-equivariant tasks.
-II-A Cryptography Methods
-Numerous works aim to securely solve matching and allocation problems involving personal sensitive information from multiple parties, revealing only the output.
-They utilize homomorphic encryption [2, 44], secret sharing [3] or garbled circuits (e.g., in [40, 15]) to perform secure multi-party computations (MPC).
-Specifically, when there is a trusted message shuffler, every client can establish information-theoretically secure communication with other clients and accomplish arbitrary multi-party computation (including DP algorithms) within a few rounds [48, 13].
-In general, MPC-based approaches impose high computation and interaction overheads, especially when the computation algorithm gets sophisticated.
-As comparison, the proposed paradigm in this work permits running arbitrary algorithms on plaintext, and requires only semi-trustness/honest-but-curious assumption on the shuffler (as the shuffler only sees ciphertexts, see Section V).
+This section reviews various approaches to private computation, primarily concentrating on non-statistical tasks.
+II-A Secure Multiparty Computation
+Secure Multiparty Computation (MPC) is a fundamental cryptographic primitive that enables multiple parties to jointly compute a function over their inputs while no party learns anything beyond their own input and the final output of the computation.
+MPC was first conceptualized by Andrew Yao in the 1980s, and it has been proven that any computable function can be realized by MPC [90].
+MPC relies on cryptographic protocols that exchange encrypted messages among parties.
+To allow computing on encrypted data, primitives such as homomorphic encryption [67, 19], secret sharing [14, 26], or garbled circuits [91] can be employed.
+The primary challenge in MPC lies in balancing privacy and efficiency.
+While MPC offers strong privacy guarantees, it often suffers from significant computational and communication overheads, which makes scaling to large datasets or numerous parties difficult.
+Recent research in MPC has been focusing on optimizing protocols and practical implementations [23, 55, 16, 24, 70, 25, 75, 15].
+Despite a significant improvement, MPC still faces efficiency issues that hinder its widespread real-world deployment.
 II-B Curator and Local DP Methods
@@
 This kind of duplex-communication shuffle model can be vulnerable to anonymity attacks [64, 66], and is pertained to the linear computation in private information retrieval.
-It can not be extended to other PIC tasks that involve with non-linear computations, and can not handle party-to-party communication in tasks like spatial crowdsourcing and social systems.
-Overall, existing works in the shuffle model primarily focus on statistical queries that are naturally permutation-invariant.
+It can not be applied to other PIC tasks (such as combinatorial optimization and federated learning with incentive) that involve with non-linear computations, and can not provide secure user-to-user communication needed in tasks like spatial crowdsourcing and social systems.
+Overall, existing works in the shuffle model primarily focus on statistical queries.
 This work, for the first time, explore the shuffle model for non-statistical applications (i.e., combinatorial optimization, location-based social systems, and incentive mechanisms).
-Technically, our work extends the shuffle model to support identity-related security primitives, such as message authentication and access control.
 II-D Combining Cryptography and DP
-While cryptographic tools can protect data secrecy during multi-party computation, they do not necessarily preserve output’s privacy.
+While cryptographic tools can protect data secrecy during multiparty computation, they do not necessarily preserve output’s privacy.
 DP can be employed to enhance the privacy of the outputting result of secure multiparty computation through decentralized noisy addition [45].
 To account for privacy loss due to intermediate encrypted views in MPC, researchers have proposed the relaxed notion of computational DP [63] against polynomial-time adversaries.
-Computational DP protocols inherited the computation/communication complexity of MPC (refer to garbled-circuits-based approaches in Table VIII).
-In contrast to enhancing relaxed privacy in the cryptographic cipher domain, our work initializes the study in the opposite direction, by integrating simple security functionalities into the privacy realm.
-Our model exposes intermediate results with information-theoretical DP guarantee.
-Therefore, it is significantly more efficient (i.e., without computation on ciphertext and requiring few rounds of interaction) and flexible (i.e., supporting arbitrary computations on noisy plaintext) than MPC with computational DP.
-III Background and Problem Settings
+Computational DP protocols often inherited the computation/communication complexity of MPC (refer to approaches in Table IX).
+III Preliminaries
+In this section, we provide a concise introduction to the preliminaries. A list of notations used throughout the paper can be found in Table I.
 TABLE I: List of notations.
@@
 A randomization mechanism $\mathcal{R}$ satisfies local $\epsilon$-DP iff $D_{\epsilon}(\mathcal{R}(x)||\mathcal{R}(x^{\prime}))=0$ for any $x,x^{\prime}\in\mathbb{X}$.
-The data processing inequality is considered a key feature of distance measures used for evaluating privacy.
-It asserts that the privacy guarantee cannot be weakened by further analysis of a private mechanism’s output.
+The data processing inequality is considered a key feature of distance measures (e.g., Hockey-stick divergence) used for evaluating privacy.
+It asserts that the privacy guarantee cannot be weakened by further analysis of a mechanism’s output.
 Definition III.4 (Data processing inequality)
@@
 the tight value of which can also be numerically computed [57, 87].
-III-C Security Primitives
-We employ several well-established security primitives for secure point-to-point communication, including key aggreement, symmetry encryption and digital signature.
-A Diffie-Hellman key agreement protocol with security parameter $\lambda$ comprises two primary steps:
-  • $sk_{i},pk_{i}\leftarrow\textsf{DH-Gen}(\lambda)$: a party $i$ generates a public-private key pair $(sk_{i},pk_{i})$, where $pk_{i}$ is the public key and $sk_{i}$ is the secret/private key;
-  • $K_{ij}\leftarrow\textsf{DH-Compute}(sk_{i},pk_{j})$: the two parties involved in the key exchange use their private keys $sk_{i}$ and $sk_{j}$ to compute a shared key $K_{ij}$ from the public keys $pk_{j}$ and $pk_{i}$ of each party (i.e., $K_{ij}=K_{ji}$.
-The resulting shared secret key, $K_{ij}$, can be used for subsequent data symmetry encryption/decryption. Specifically, a symmetry encryption scheme with secret key $sk$ consists of two steps:
-  • $c\leftarrow\textsf{Enc}(sk,x)$: generates a ciphertext $c$ by encrypting data $x$ with the encryption key $sk$;
-  • $x\leftarrow\textsf{Dec}(sk,c)$: decrypts ciphertext $c$ with the key $sk$, outputting $x$.
-A digital signature scheme permits the receipt to authorize ciphertexts and to ensure message integrity, through the following steps:
-  • $t\leftarrow\textsf{Sign}(sk,c)$: returns a tag $t$ as the output of signing data $c$ with the private key $sk$;
-  • $accept/reject\leftarrow\textsf{Verify}(pub,t,c)$: outputs $accept$ when $t=\textsf{Sign}(sk,c)$, and outputs $reject$ with a probability of $1-negl(k)$ when $t\neq\textsf{Sign}(sk,c)$.
+III-C Public Key Encryption
+[…]
+  • Gen: takes as input the security parameter $\lambda$ and outputs a pair of keys $(pk,sk)$, where $pk$ denotes the public key and $sk$ the private key.
+  • Enc: takes as input a public key $pk$ and a message $m$ from the plaintext space. It outputs a ciphertext $c\rightarrow\textsf{Enc}_{pk}(m)$.
+  • Dec: takes as input a private key $sk$ and a ciphertext $c$, and outputs a message $m$ or a special symbol $\bot$ denoting failure.
+It is required that for every $(pk,sk)$ and plaintext message $m$, it holds that $\textsf{Dec}_{sk}(\textsf{Enc}_{pk}(m))=m$.
+The IND-CPA security ensures the scheme leaks no useful information under a chosen plaintext attack.
 IV Problem Settings
-We aim to extend the capability of the shuffle model to support private individual computation.
-This would have multiple benefits: (1) we can get a formal privacy guarantee (in the DP sense) on a much wider range of data-driven computational tasks; (2) because the computation is done on already sanitized data, we can avoid heavy cryptographic protocols and scale to large data and/or user size; (3) the privacy amplification effect will allow much better utility, compared to sanitizing and computing directly using LDP.
-In this section, we will set the stage by first presenting a few motivating applications, and then formally defining private individual computation as an ideal functionality.
-Figure 1: An illustration of conceptual overview of PIC model for spatial crowdsourcing.
+We now introduce Private Individual Computation, a new paradigm for privacy-preserving computation that offers several benefits: (1) It provides a formal privacy guarantee (in the differential privacy sense) for a wide range of data-driven computational tasks.
+(2) Since the computation is performed on user-sanitized data, it avoids the need for heavy cryptographic protocols, allowing for scalability to handle large datasets and user bases.
+(3) The privacy amplification effect results in significantly better utility compared to direct sanitization and computation using Local Differential Privacy (LDP).
+In this section, we will first present a few motivating applications and then formally define Private Individual Computation as an ideal functionality.
+Figure 1: An illustration of taxi-hailing in the PIC model. Besides (sanitized) location information, each user also encapsulates a one-time random public key into the message to the shuffler.
 IV-A Motivating Applications
-We describe three prevalent exemplar computation tasks: spatial crowdsourcing, location-based services, and federated learning with incentives:
-Spatial crowdsourcing. A spatial crowdsourcing system normally consists of three roles: users, workers, and the orchestrating server (the service platform), and proceed as following major steps:
+We describe three prevalent exemplar computation tasks: spatial crowdsourcing, location-based social systems, and federated learning with incentives:
+Spatial crowdsourcing.
+A spatial crowdsourcing system typically consists of three roles: users, workers, and the orchestrating server (the service platform).
+It proceeds through the following major steps:
   • I.
-    Task submission: each participated user $i\in G_{a}$ submits a task (e.g., taxi calling requests, sensing/photo requests) $x_{i}=(i,l_{i},v_{i})$ containing the location $l_{i}$ and possibly other information $v_{i}$;
+    Task submission: each user $i\in G_{a}$ submits a task (e.g., taxi calling requests, sensing/photo requests) $x_{i}=(i,l_{i},v_{i})$ containing the location $l_{i}$ and possibly other information $v_{i}$;
   • II.
@@
   • IV.
-    Task performing: users and workers retrieve matching results $M$ and collaboratively complete the task.
-Location-based services. In a location-based service, a set of users and a server interact as the following:
+    Task performing: users and workers retrieve matching results and collaboratively complete the task.
+One example of taxi-haling is illustrated in Figure 1.
+Location-based social systems. In a location-based social system, a set of users and a server interact as the following:
   • I.
@@
     Retrieving: Finally, each user retrieves the recommendation results.
-Federated learning with incentives. Federated learning involves a set of clients and a server:
+Federated learning with incentives. Federated learning involves a set of users and a server:
   • I.
-    Submitting gradient: Each client $i\in G_{a}$ in each epoch computes an intermediate gradient information $x_{i}\in[-1,1]^{d}$ with its local model and data, then submits to the server;
+    Submitting gradient: Each user $i\in G_{a}$ in each epoch computes an intermediate gradient information $x_{i}\in[-1,1]^{d}$ with its local model and data, then submits to the server;
   • II.
-    Computing incentive: The server computes the averaged gradient $\overline{x}=\frac{1}{n}\sum_{i\in[n]}x_{i}$.
-To incentive participation, the server may reward clients with monetary tokens (e.g., via cryptocurrency) according to a profit allocation algorithm $V:[-1,1]^{d\times n}\times[-1,1]^{d}\mapsto[0,1]$ (e.g., Shapley value [69]).
+    Computing incentive: The server computes the average gradient $\overline{x}=\frac{1}{n}\sum_{i\in[n]}x_{i}$.
+To incentive participation, the server may reward users with monetary tokens (e.g., via cryptocurrency) according to a profit allocation algorithm $V:[-1,1]^{d\times n}\times[-1,1]^{d}\mapsto[0,1]$ (e.g., Shapley value [71]).
   • II.
-    Receiving incentive: Each client retrieves the token and claims its monetary incentive.
-A fundamental distinction between the above applications and those currently studied in the shuffle model is that now each participant expects an output that differs individually, instead of a single collectively aggregated output.
-Informally, in such applications, we would need to safeguard data privacy so that apart from the party who generates the data, no one should be certain about that party’s data (up to the leakage allowed by DP).
-To be able to amplify privacy, we also need to maintain anonymity so that for any message, an adversary (e.g.
-the server, an observer, an unmatched user) should only know this message comes from a user belonging to a particular group, but nothing more.
-Note that following the convention in the shuffle model, the shuffler is trusted for providing anonymity.
-So the shuffler will know the random permutation used in the shuffling process and will not leak it to other parties, although we prevent the shuffler from observing plaintext messages through encryption.
+    Receiving incentive: Each user retrieves the token and claims its monetary incentive.
+A fundamental distinction between the above applications and those currently studied in the shuffle model is that each participant now expects an output that differs individually, rather than a single collectively aggregated output.
+Another distinction is that a participant might need to securely communicate with other participants after receiving the individualized computation results, such as the matched user and worker in spatial crowdsourcing will communicate with each other to accomplish the task, and the matched users in location-based social systems would like to securely contact each other afterward.
+Informally, in such applications, it is necessary to safeguard data privacy so that, aside from the party who generates the data, no one can be certain about that party’s data (up to the leakage allowed by differential privacy).
+To amplify privacy, we also need to maintain anonymity so that for any given message, an adversary (such as the server, an observer, or an unmatched user) should only know that this message comes from a user within a particular group, but nothing more.
+Following the convention in the shuffle model, the shuffler is trusted to provide anonymity.
+The shuffler knows the random permutation used in the shuffling process and will not leak it to other parties, although the shuffler is prevented from observing plaintext messages through encryption.
 IV-B The Ideal Functionality
@@
 Functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ Parameters: $m\in\mathbb{N}$; $m$ groups of parties $G_{1},G_{2},\cdots,G_{m}$, where $n_{i}=|G_{i}|$ denotes the number of parties in group $G_{i}$; the data randomization mechanisms $\mathcal{R}_{i}$ for group $G_{i}$; a server ${S}$.
-Functionality: Upon receiving $n=\sum_{i\in[m]}n_{i}$ inputs $\{x_{i,j}\}_{i\in[m],j\in[|G_{i}|]}$ from all clients, and the description of a function $f$ to be computed over parties’ inputs from the server, do the following: Compute $x^{\prime}_{i,j}\leftarrow R_{i}(x_{i,j})$ for all $i\in[m]$ and $j\in[|G_{i}|]$.
-Sample $m$ random permutations $\pi_{1},\pi_{2},\cdots,\pi_{m}$, where $\pi_{i}:[n_{i}]\rightarrow[n_{i}]$, perform shuffle over inputs and obtain $L=\left(\{x_{1,\pi_{1}(j)}^{\prime}\}_{j\in[n_{1}]},\cdots,x_{m,\pi_{m}(j)}^{\prime}\}_{j\in[n_{m}]}\right)$, and compute ($\left(\{y_{1,\pi_{1}(j)}\}_{j\in[n_{1}]},\cdots,\{y_{m,\pi_{m}(j)}\}_{j\in[n_{m}]}\right)\leftarrow f(L)$.
+Functionality: Upon receiving $n=\sum_{i\in[m]}n_{i}$ inputs $\{x_{i,j}\}_{i\in[m],j\in[|G_{i}|]}$ from all users, and the description of a function $f$ to be computed over parties’ inputs from the server, do the following: Compute $x^{\prime}_{i,j}\leftarrow R_{i}(x_{i,j})$ for all $i\in[m]$ and $j\in[|G_{i}|]$.
+Sample $m$ random permutations $\pi_{1},\pi_{2},\cdots,\pi_{m}$, where $\pi_{i}:[n_{i}]\rightarrow[n_{i}]$, perform shuffle over inputs and obtain $L=\left(\{x_{1,\pi_{1}(j)}^{\prime}\}_{j\in[n_{1}]},\cdots,\{x_{m,\pi_{m}(j)}^{\prime}\}_{j\in[n_{m}]}\right)$, and compute ($\left(\{y_{1,\pi_{1}(j)}\}_{j\in[n_{1}]},\cdots,\{y_{m,\pi_{m}(j)}\}_{j\in[n_{m}]}\right)\leftarrow f(L)$.
 Send $y_{i,\pi_{i}(j)}$ to party $u_{i,j}$ for all $i\in[m]$ and $j\in n_{i}$.
@@
 Figure 2: The functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$
-Essentially, the ideal functionality represents a fully trusted party that interacts with $m$ groups of clients and one server $S$.
-It sanitizes the input from each client, then shuffles them randomly, and applies a function from the server to compute the output for each client.
-At the end, each client receives its function output, and the server gets the list of shuffled sanitized data from the clients and a list of the function output.
-This captures the functional requirement of individual computation in the real world: a computing task is performed by a server using the joint inputs from a set of clients (which are sanitized and shuffled), and each client gets an individualized output.
-It also captures the privacy requirements: each party should receive precisely the specified output, nothing more.
-Remark A careful reader may notice that the ideal function does not capture differential privacy.
-This is intentional for the sake of security analysis.
-Later in the security analysis, we will decouple different privacy requirements into two sets of proofs: the first set shows that our concrete protocol executed by real-world parties realizes the ideal functionality, meaning that no additional information about parties’ inputs is revealed, except the output given to each party; then we prove that the output conforms with differential privacy.
+Essentially, the ideal functionality represents a fully trusted party that interacts with $m$ groups of users and one server $S$.
+It sanitizes the input from each user, shuffles the inputs randomly, and then applies a function to compute the output for each user.
+In the end, each user receives their individual function output, while the server receives a list of the shuffled sanitized user data and a list of the function outputs.
+It captures the functional requirements of individual computation in the real world: a server performs a computing task using the joint inputs from a set of users (which are sanitized and shuffled), and each user receives an individualized output.
+It also captures the security requirements: each party receives precisely the specified output, and nothing more.
+Remark 1 A careful reader may notice that the ideal functionality does not explicitly capture differential privacy.
+This omission is intentional for the sake of security analysis.
+In the security analysis, we decouple the privacy requirements into two sets of proofs: the first set demonstrates that our concrete protocol, when executed by real-world parties, realizes the ideal functionality.
+This means no additional information about parties’ inputs is revealed, except for the output given to each party.
+The second set of proofs establishes that the output given to each party conforms with differential privacy.
 V A Concrete Protocol
 V-A The Protocol
-We now present a concrete protocol for private individual computation.
+We now present a concrete protocol for PIC.
 Because there are already secure protocols for shuffling, to simplify the description, we present the protocol in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model, in which parties can communicate as usual, and in addition have access to an ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ that does the shuffling.
@@
 Figure 3: The functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$
-The Private Individual Computation protocol is outlined below:
+The PIC protocol is outlined below:
   1. 1.
@@
   2. 2.
-    We denote the $j$-th client in group $G_{i}$ as $u_{i,j}$.
-Each client generates a key pair $(pk_{i,j},sk_{i,j})\leftarrow\textsf{Gen}(\lambda)$.
-Each client then randomizes its private data $x_{i,j}$ with mechanism $\mathcal{R}_{i}$ and obtains $x_{i,j}^{\prime}\leftarrow\mathcal{R}_{i}(x_{i,j})$.
-Then the sanitized input is concatenated with its own public key, and encrypted with the server’s public key $x_{i,j}^{\prime\prime}\leftarrow\textsf{Enc}_{pk_{c}}(pk_{i,j}||x^{\prime}_{i,j})$.
+    We denote the $j$-th user in group $G_{i}$ as $u_{i,j}$.
+Each user generates a key pair $(pk_{i,j},sk_{i,j})\leftarrow\textsf{Gen}(\lambda)$.
+Each user then randomizes their private data $x_{i,j}$ with mechanism $\mathcal{R}_{i}$ and obtains $x_{i,j}^{\prime}\leftarrow\mathcal{R}_{i}(x_{i,j})$.
+Then the sanitized input is concatenated with their own public key, and encrypted with the server’s public key $x_{i,j}^{\prime\prime}\leftarrow\textsf{Enc}_{pk_{c}}(pk_{i,j}||x^{\prime}_{i,j})$.
   3. 3.
-    Each client in group $G_{i}$ invokes $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ with $x_{i,j}^{\prime\prime}$, and $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ outputs the shuffled messages $\{x_{i,\pi_{i}(j)}^{\prime\prime}\}_{j\in[n_{i}]}\leftarrow\mathcal{S}(\{x_{i,j}^{\prime\prime}\}_{j\in[n_{i}]})$ to the server, where $\pi^{-1}_{i}$ is the (secret) random permutation used during $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ for group $G_{i}$.
+    Each user in group $G_{i}$ invokes $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ with $x_{i,j}^{\prime\prime}$, and $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ outputs the shuffled messages $\{x_{i,\pi_{i}(j)}^{\prime\prime}\}_{j\in[n_{i}]}\leftarrow\mathcal{S}(\{x_{i,j}^{\prime\prime}\}_{j\in[n_{i}]})$ to the server, where $\pi^{-1}_{i}$ is the (secret) random permutation used during $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ for group $G_{i}$.
   4. 4.
@@
     $\ \ \ \left.\{pk_{m,\pi_{m}(j)}||x_{m,\pi_{m}(j)}^{\prime}\}_{j\in[n_{m}]}\right)$
-    It then computes the function $f$ over $L$ to produce output for each anonymous client:
-    ($\left(\{y_{1,\pi_{1}(j)}\}_{j\in[n_{1}]},\cdots,\{y_{m,\pi_{m}(j)}\}_{j\in[n_{m}]}\right)\leftarrow f(L)$
+    It then computes the function $f$ over $L$ to produce output for each anonymous user:
+    ($\left(\{y_{1,\pi_{1}(j)}\}_{j\in[n_{1}]},\cdots,\{y_{m,\pi_{m}(j)}\}_{j\in[n_{m}]}\right)\leftarrow f(L).$
   5. 5.
@@
   6. 6.
-    Every client downloads the list, finds in the list the entry with its own public key, and decrypts the payload to get the computation result.
+    Every user downloads the list, finds in the list the entry with their own public key, and decrypts the payload to get the computation result.
+Remark 2 In the final step, we adopt the simplest strategy for users to retrieve their results without the server knowing which entry belongs to whom.
+If bandwidth is a concern, it can be replaced by a more sophisticated (and computationally more expensive) Private Information Retrieval protocol (e.g.
+[46]).
+Remark 3 To eventually accomplish the spatial crowdsourcing (e.g., taxi-hailing services) or location-based social system tasks, the computation result $y_{i,\pi_{i}(j)}$ will contain the matched users of the user $\pi_{i}(j)\in G_{i}$.
+That is, $y_{i,\pi_{i}(j)}$ will encapsulate a list of public keys and noisy location information about the matched users of the user $\pi_{i}(j)\in G_{i}$.
+After receiving the individualized computation results in the final step, each user $\pi_{i}(j)$ can then securely contact the matched users using the public keys in $y_{i,\pi_{i}(j)}$ (possibly with the help of a public communication channel, such as a public bulletin).
 V-B Security Analysis
@@
 This means the protocol leaks no more information than what is allowed by $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ and $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$.
-More precisely, each client gets their output from $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$, the server gets $(L,f(L))$, and in the case of colluding with some clients in $G_{i}$, the partial permutation $\mathcal{L}(\pi_{i})$.
+More precisely, each user gets their output from $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$, the server gets $(L,f(L))$, and in the case of colluding with some users in $G_{i}$, the partial permutation $\mathcal{L}(\pi_{i})$.
 Formally we have the following theorem:
@@
 The Private Individual Computation (PIC) protocol in §V securely computes $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model in the presence of any PPT adversary with static corruption.
-The proof can be found in Appendix -J.
+The proof can be found in Appendix -A.
 The proof is simulation-based.
@@
 $\max(D_{\epsilon_{c}}(\mathcal{A}(X)||\mathcal{A}(X^{\prime})),D_{\epsilon_{c}}(\mathcal{A}(X^{\prime})||\mathcal{A}(X)))\leq\delta.$
-In particular, when all users in $G_{i^{*}}$ use an identical $\epsilon$-LDP mechanism as the data randomizer $\mathcal{R}_{i^{*}}$, and for $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|\geq\tilde{\Omega}(e^{\epsilon})$ we have:
-$\small\epsilon_{c}=\log\Bigg{(}1+\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\Big{(}\sqrt{\frac{32(e^{\epsilon}+1)\log 4/\delta}{n^{\prime}_{i^{*}}}}+\frac{4(e^{\epsilon}+1)}{n^{\prime}_{i^{*}}}\Big{)}\Bigg{)}.\normalsize$ (1)
+In particular, when all users in $G_{i^{*}}$ use an identical $\epsilon$-LDP mechanism as the data randomizer $\mathcal{R}_{i^{*}}$, and for $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|\geq 8(e^{\epsilon}+1)\log(2/\delta)$ we have:
+$\epsilon_{c}=\log\Bigg{(}1+\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\Big{(}\sqrt{\frac{32(e^{\epsilon}+1)\log 4/\delta}{n^{\prime}_{i^{*}}}}+\frac{4(e^{\epsilon}+1)}{n^{\prime}_{i^{*}}}\Big{)}\Bigg{)}.$ (1)
 The analysis can be divided into two cases: in the first case, $S\in\mathcal{C}$, i.e.
 S is corrupted.
-In this case, since the honest client locally randomizes their input, the input enjoys at least $\epsilon$-DP.
+In this case, since the honest user locally randomizes their input, the input enjoys at least $\epsilon$-DP.
 Then the shuffling will amplify the privacy guarantee.
@@
 Continuing along the same line of thought, we conclude that the level of differential privacy assurance is no less than in the first case.
-The full proof can be found in Appendix -A.
+The full proof can be found in Appendix -B.
+Remark 4 After decryption, the server obtains $pk_{i,j}||x^{\prime}_{i,j}$, where $pk_{i,j}$ is a public key not sanitized by the local randomizer.
+Despite the presence of the public key, $pk_{i,j}||x^{\prime}_{i,j}$ and $x^{\prime}_{i,j}$ are equivalent in terms of privacy amplification.
+This is because (1) the public key is random and generated independently of $x_{i,j}$, so prepending it to $x^{\prime}_{i,j}$ does not affect the local DP guarantee—it is the same as $x^{\prime}_{i,j}$ itself; and (2) all public keys follow an identical probability distribution across all users in the group $G_{i}$, ensuring that the privacy amplification effect via shuffling is not degraded.
 V-C Discussion on Post-computation Communication
-In certain scenarios, such as spatial crowdsourcing and location-based social systems, there may be additional client-to-client communications following the execution of the PIC protocol.
-For instance, consider a ridesharing application where passengers are in a group $G_{1}$ and taxi drivers in a group $G_{2}$.
+In certain scenarios, such as spatial crowdsourcing and location-based social systems, there may be additional user-to-user communications following the execution of the PIC protocol.
+For instance, consider a taxi-hailing application where passengers are in a group $G_{1}$ and taxi drivers in a group $G_{2}$.
 After receiving the matching result at the end of the protocol, the passengers must send their locations to the matched drivers, who need to know where to pick them up.
@@
 Inevitably, a party has to sacrifice their privacy to the matched parties.
-The post-computation communication may also have privacy implications for other clients not in the matched pair.
+The post-computation communication may also have privacy implications for other users not in the matched pair.
 Privacy amplification via shuffling against an adversary relies on the number of parties that remain anonymous.
 Recall that in Equation 1, the amplified $\epsilon_{c}$ relies on the number $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|$ of uncorrupted users in a particular group $G_{i^{*}}$.
-From the perspective of the adversary, if the post-computation communication compromises the anonymity of an additional set $U$ of the clients in $G_{i^{*}}$, then $n^{\prime}_{i^{*}}$ becomes $|G_{i^{*}}-C_{i^{*}}-U|$.
+From the perspective of the adversary, if the post-computation communication compromises the anonymity of an additional set $U$ of the users in $G_{i^{*}}$, then $n^{\prime}_{i^{*}}$ becomes $|G_{i^{*}}-C_{i^{*}}-U|$.
 Accordingly, the privacy amplification effect for users in $G_{i^{*}}$ that are still anonymous is weakened.
@@
 VI-A Inadequacy of Existing Randomizers
-In the PIC model, each user first sanitizes their data using an $\epsilon$-LDP randomizer.
+In PIC, each user first sanitizes their data using an $\epsilon$-LDP randomizer.
 The design of the randomizer significantly impacts the utility of the tasks.
@@
 The shuffle model aims to estimate certain statistics from the noisy data collected from users, so it cares about how close the estimation is to the true value of the desired statistic.
-In the literature, utility is often measured by the expected square error (i.e., the variance) of the estimation:
-$\mathbb{E}_{\scriptscriptstyle T\in\mathbb{X}^{n}}[\|\tilde{f}(T)-f(T)\|_{2}^{2}]=\textsf{Var}_{\scriptscriptstyle T\in\mathbb{X}^{n}}[\tilde{f}(T)]$
+In the literature, utility is often measured by the expected square error (i.e., the variance) bound of the estimation:
+$\max_{\scriptscriptstyle T\in\mathbb{X}^{n}}\mathbb{E}[\|\tilde{f}(T)-f(T)\|_{2}^{2}]=\max_{\scriptscriptstyle T\in\mathbb{X}^{n}}\textsf{Var}[\tilde{f}(T)]$
 where $f$ is a statistical function, and $\tilde{f}$ is its estimation output by the shuffle protocol.
@@
 These factors together create issues when existing LDP randomizers are applied directly in the PIC model.
-To see the problem, we first examine a class of LDP randomizers [64, 84, 58, 50]) that work by sampling a few dimensions from $[d]$.
-Each user submits an incomplete report that contains only the sampled dimensions (with added noise) in their local data.
-This is an effective strategy for statistical estimation, but will totally ruin the utility of a single report in PIC as unsampled dimensions are missing.
-Clearly, they are unsuitable for use in the PIC model.
+To understand the problem, we first examine a class of LDP randomizers [65, 86, 59, 51] that operate by sampling a few dimensions from $[d]$.
+Each user submits an incomplete report that contains only the sampled dimensions (with added noise) from their local data.
+On the positive side, this strategy reduces the amount of noise added to the sampled dimensions.
+On the negative side, the unsampled dimensions are missing.
+In statistical estimation tasks, the estimation is made using all reports, each covering some dimensions.
+Therefore, the incompleteness of a single report is less important, and better utility can be achieved.
+However, in the PIC model, where the focus shifts to the error of individual reports, this strategy may lead to worse results.
 There have been LDP randomizers that submit complete reports.
@@
 The Laplace/PlanarLaplace mechanisms introduce an MSE rate of $\frac{d}{\epsilon^{2}}$, while the PrivUnit/PrivUnitG mechanisms incur an MSE rate of $\frac{d}{\min\{\epsilon,\epsilon^{2}\}}$.
-In contrast, later we will show that the MSE rate can be improved to $(e^{\epsilon}-1)^{-2/(d+2)}$ (see Section VI-C and Appendix -B).
+In contrast, later we will show that the MSE rate can be improved to $(e^{\epsilon}-1)^{-2/(d+2)}$ (see Section VI-C and Appendix -D).
 Another strategy for submitting complete reports involves using additional randomization techniques such as random projection, data sketches, public randomness, or quantization.
@@
 VI-B Randomizer Design
-We now introduce an asymptotic optimal randomizer, tailored for the PIC model.
+We now introduce an asymptotically optimal randomizer, tailored for the PIC model.
 The randomizer uses an LDP mechanism, which we termed as Minkowski Response.
-For ease of presentation, here we will focus primarily on the $\ell_{2}$ case (and the $\ell_{+\infty}$ case in Appendix -B), but the mechanism can be generalized to other Minkowski distances.
-Without loss of generality, we assume the user data domain to be a $\ell_{2}$-bounded hyperball $\mathbb{X}=\{x\ |\ x\in\mathbb{R}^{d}\ and\ \|x\|_{2}\leq 1\}$.
+For ease of presentation, here we will focus primarily on the $\ell_{2}$ case (and the $\ell_{+\infty}$ case in Appendix -D), but the mechanism can be generalized to other Minkowski distances.
+Without loss of generality, we assume the user data domain to be an $\ell_{2}$-bounded hyperball $\mathbb{X}=\{x\ |\ x\in\mathbb{R}^{d}\ and\ \|x\|_{2}\leq 1\}$.
 Most, if not all, real-world data domains can be normalized to $\mathbb{X}$ (e.g.
@@
 $\mathbb{Y}_{r}=\{y\ |\ y\in\mathbb{R}^{d}\ and\ \exists x\in\mathbb{X}\ that\ y\in\mathbb{B}_{r}(x)\}.$
-Figure 4: Illustrative probability design of Minkowski response mechanism over two dimensions with cap area radius $r$, given different inputs $x_{i}$ in (a) and (b) respectively.
+Figure 4: The probability design of Minkowski response mechanism with a radius $r$. Illustrated are three inputs $x_{1},x_{2}$ and $x_{3}$, along with their respective cap areas.
 For any input $x\in\mathbb{X}$, Minkowski Response outputs an output $y\in\mathbb{Y}_{r}$ with relatively high probability in the cap area $\mathbb{B}_{r}(x)$ and relatively low probability in remaining domain $\mathbb{Y}_{r}\backslash\mathbb{B}_{r}(x)$ (see Figure 4).
@@
 Also, intuitively it has a better utility because the mechanism is more likely to output a value near the true value (i.e., in $\mathbb{B}_{r}(x)$), than from other parts of the output domain (i.e.
-$\mathbb{Y}_{r}$).
+$\mathbb{Y}_{r}\backslash\mathbb{B}_{r}(x)$).
 When the budget $\epsilon$ gets large, the error rate of Minkowski Response decays faster than in previous LDP mechanisms.
-More specifically, the decay rate of Minkowski Response is $(e^{\epsilon}-1)^{-2/(d+2)}$ (Equation 5 in Appendix -B), while that of the previous mechanisms is $d/(e^{\epsilon/d}-1)^{2/3}$ or $d/\epsilon^{2}$.
+More specifically, the decay rate of Minkowski Response is $(e^{\epsilon}-1)^{-2/(d+2)}$ (Equation 6 in Appendix -D), while that of the previous mechanisms is $d/(e^{\epsilon/d}-1)^{2/3}$ or $d/\epsilon^{2}$.
 Therefore, the utility advantage of Minkowski Response becomes more significant when $\epsilon$ gets larger.
@@
 Theorem VI.2 (Error Lower Bounds)
-Given $d\in\mathbb{N}$, $\epsilon_{c}>0$, $\delta\in(0,0.5]$, $\mathbb{X}=\mathbb{B}_{1}(\{0\}^{d})$, then for any randomizer $\mathcal{R}:\mathbb{X}\mapsto\mathbb{R}^{d^{\prime}}$ such that $\mathcal{S}\circ\mathcal{R}(X)$ and $\mathcal{S}\circ\mathcal{R}(X^{\prime})$ are $(\epsilon_{c},\delta)$-indistinguishable for all possible neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, and for any estimator $f:\mathbb{R}^{d^{\prime}}\mapsto\mathbb{X}$, we have $\max_{x\in\mathbb{X}}\mathbb{E}\big{[}\|f\circ\mathcal{R}(x)-x\|_{2}^{2}\big{]}\geq\tilde{\Omega}\big{(}{1}/{n^{\frac{2}{d+2}}}\big{)}.$
+Given $d\in\mathbb{N}$, $\epsilon_{c}>0$, $\delta\in(0,0.5]$, $\mathbb{X}=\mathbb{B}_{1}(\{0\}^{d})$, then for any randomizer $\mathcal{R}:\mathbb{X}\mapsto\mathbb{R}^{d^{\prime}}$ such that $\mathcal{S}\circ\mathcal{R}(X)$ and $\mathcal{S}\circ\mathcal{R}(X^{\prime})$ are $(\epsilon_{c},\delta)$-indistinguishable for all possible neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, and for any estimator $f:\mathbb{R}^{d^{\prime}}\mapsto\mathbb{R}^{d}$, we have:
+$\max_{x\in\mathbb{X}}\mathbb{E}\big{[}\|f\circ\mathcal{R}(x)-x\|_{2}^{2}\big{]}\geq\tilde{\Omega}\big{(}{1}/{n^{\frac{2}{d+2}}}\big{)}.$
 The above theorem suggests that in the PIC model, for any randomizer, the single report error is at least $\tilde{\Omega}\big{(}{1}/{n^{\frac{2}{d+2}}}\big{)}$.
@@
 For the randomizer using Minkowski Response, as presented in section VI-B, we can demonstrate that its single report error has an upper bound.
-This is summarized in Theorem VI.3, with the proof provided in Appendix -B.
+This is summarized in Theorem VI.3, with the proof provided in Appendix -D.
 Theorem VI.3 (Error Upper Bounds)
@@
 $\max_{x\in\mathbb{X}}\mathbb{E}\big{[}\|\mathcal{R}(x)-x\|_{2}^{2}\big{]}\leq O\Bigg{(}\Big{(}\frac{\log(1/\delta)}{n\epsilon_{c}^{2}}\Big{)}^{\frac{2}{d+2}}\Bigg{)}.$
-Given Theorems VI.2 and VI.3, we observe that as $n\rightarrow\infty$, the error upper bound $\tilde{O}(\frac{1}{n^{2/(d+2)}})$ of the Minkowski Response randomizer is the same as the error lower bound of the PIC model (the $\tilde{O}$ notation suppresses factors about $\epsilon_{c},\delta$ for simplicity).
-This means that the utility of the Minkowski Response is asymptotically optimal.
-In contrast, existing LDP randomizers with privacy amplification in the PIC model will suffer the single report error of $\tilde{O}(\frac{d}{n^{2/(3d)}})$ or $\tilde{O}(\frac{d}{\log^{2}n})$, whose denominator grows much slower than $n^{2/(d+2)}$ when $d>1$ and $n$ is relatively large.
-In practice, $n$ does not need to be very large; in our experiments, Minkowski Response outperforms existing LDP randomizers in PIC model once $n$ reaches the order of $10^{2}$.
-When comparing the PIC model to local DP model, the gap is undoubtedly larger, as numerical methods of privacy amplification [56, 85] will always give a local budget $\epsilon$ that is larger than the global privacy $\epsilon_{c}$ as long as $n\geq 2$, creating an extra gap in between the local and PIC models of these LDP randomizers.
-Actually, any randomizers in the LDP model must endures an error of $\Omega({d}/{\epsilon_{c}^{2}})$ for single report when $\epsilon_{c}\leq O(1)$ [18, 25].
+We observe that in most applications, $(\epsilon_{c},\delta)$ are given as fixed system parameters.
+If we consider $(\epsilon_{c},\delta)$ as constants, then the error upper bound of the Minkowski Response randomizer in Theorem VI.3 is $\tilde{O}\left(\frac{1}{n^{2/(d+2)}}\right)$, which matches the error lower bound of the PIC model in Theorem VI.2.
+This implies that the utility of the Minkowski Response is asymptotically optimal.
+In contrast, using existing LDP randomizers in the PIC model results in a larger single report error of $\tilde{O}\left(\frac{d}{n^{2/(3d)}}\right)$ or $\tilde{O}\left(\frac{d}{\log^{2}n}\right)$ (note that $d>1$ and $n$ is often not small).
+Although asymptotic notations describe behavior as $n\rightarrow\infty$, in practice, the utility advantage of the Minkowski Response becomes noticeable without $n$ being very large: in our experiments, the Minkowski Response randomizer outperforms existing LDP randomizers in the PIC model once $n$ reaches the order of $10^{2}$.
+If we compare Minkowski Response in the PIC model to using LDP directly (without shuffling), the utility advantage is even greater: any randomizers in the LDP model must endure a single report error of $\Omega\left(\frac{d}{\epsilon_{c}^{2}}\right)$ when $\epsilon_{c}\leq O(1)$ [12, 28].
 VII Experimental Evaluation
-We evaluate the utility and efficiency performances of proposed PIC model in three representative individual computation tasks: spatial crowdsourcing, location-based nearest neighbors, and federated learning with incentives, aim to answer the following research questions:
-  • RQ1.
-    How much do our proposals outperform existing approaches in accuracy/utility?
-  • RQ2.
-    What the computational/communication costs of our model for supporting security functionalities?
-TABLE II: Mean $\ell_{2}$-error (and running time) comparison of local $\epsilon$-LDP randomizers on location domain $[-1,1]\times[-1,1]$.
-All results are the expected value of $1000$ repeated experiments.
-randomizer $\epsilon=0.5$ $\epsilon=1.0$ $\epsilon=2.0$ $\epsilon=3.0$ $\epsilon=5.0$ $\epsilon=8.0$ $\epsilon=10.0$
-PrivUnit [16] 8.94 (49us) 4.68 (42us) 2.25 (61us) 1.44 (63us) 0.81 (113us) 0.32 (305us) 0.18 (868us)
-PrivUnitG[8] 8.73 (76ms) 4.63 (75ms) 2.27 (77ms) 1.51 (76ms) 0.96 (74ms) 0.63 (75ms) 0.53 (81ms)
-Laplace [28] 12.97 (0.1us) 6.56 (0.1us) 3.27 (0.1us) 2.13 (0.1us) 1.30 (0.1us) 0.81 (0.1us) 0.64 (0.1us)
-PlanarLaplace [4] 11.17 (12us) 5.63 (11us) 2.84 (11us) 1.88 (11us) 1.14 (11us) 0.71 (12us) 0.56 (12us)
-Staircase[35] 13.19 (49us) 6.40 (48us) 3.13 (46us) 2.01 (48us) 1.05 (44us) 0.46 (51us) 0.28 (50us)
-SquareWave[57] 11.87 (1.9us) 5.72 (2.5us) 2.65 (1.9us) 1.68 (2.2us) 0.92 (1.7us) 0.53 (2.2us) 0.42 (2.1us)
-MinkowskiResponse 10.42 (0.7us) 4.50 (0.6us) 1.78 (0.8us) 0.98 (0.8us) 0.39 (0.6us) 0.14 (0.7us) 0.074 (0.8us)
+We evaluate the efficacy of our PIC protocol and Minkowski randomizer.
+We compare the utility of our proposal against state-of-the-art works in the context of three representative individual computation tasks: spatial crowdsourcing, location-based social systems, and federated learning with incentives.
 VII-A Spatial Crowdsourcing
-Datasets and settings.
-In the experiments for spatial crowdsourcing, we use use two real datasets, GMission dataset [20] and EverySender dataset [79], for scientific simulation and campus-based micro-task completion respectively.
-The number of users/workers, location domain range, and serving radius of workers about these two datasets are summarized in Table III (more information are presented in Appendix -F).
-In our experiments, for compliance with data randomizers (see following parts), we rescale the location domain range to $[-1,1]\times[-1,1]$ and rescale the corresponding serving radius to $0.41.0\cdot\frac{1-(-1)}{5-0}=$ 0.4.
-TABLE III: Summary statistics of spatial crowdsourcing datasets.
+Datasets We use two real-world datasets: GMission dataset [18] for scientific simulation, and EverySender dataset [82] for campus-based micro-task completion.
+Details about the two datasets are summarized in Table II, including the number of users/workers, location domain range, and serving radius of workers about these two datasets (more information can be found in Appendix -F).
+We normalize location data to the domain $[-1,1]\times[-1,1]$ and scale the serving radius to $0.41.0\cdot\frac{1-(-1)}{5-0}=$ 0.4 correspondingly.
+TABLE II: Summary statistics of spatial crowdsourcing datasets.
 Dataset users workers location domain serving radius
@@
 EverySender 4036 817 $[0,5.0]\times[0,5.0]$ $1.0$
-Cryptography implementation.
-Follow the protocol in Section V, we implement the PIC model using Elliptical Curve Diffie-Hellman key agreement (ECDHE [11]) for establishing shared secret between client/server and matched user/worker, meanwhile using AES encryption [68] and HMAC signature [14] for secure point-to-point communication after key agreement.
-Data randomizers and server-side algorithms.
-We use the Minkowski response for location randomization.
-As comparison, in the local model of DP (e.g., in [83, 77, 81]), we compare with existing mechanisms including Laplace [28], Staircase [35], PlanarLaplace with geo-indistinguishability [4], SquareWave [57], PrivUnit [16] and its Gaussian variant PrivUnitG in [8].
-On the server side, we use either minimum weighted full matching [51] or maximum matching [45] over bipartite graph as the user/worker matching algorithm based on (noisy) locations, to obtain a match $M:G_{a}\times G_{b}\mapsto\mathbb{R}$.
-Specifically, the minimum weighted full matching aims to minimize the overall travelling costs between users and works over the (noisy) fully-connected bipartite graph; the maximum matching aims to maximize the number of successfully matched user/worker pairs over the (noisy) partially-connected bipartite graph, where user/worker pairs with distance surpassing $0.4$ (i.e.
-the serving radius) are deemed as not connected.
-Performance metrics.
-We evaluate the performance of spatial crowdsourcing in terms of travel costs, successful matching ratio, and running time.
-The travel costs are the total actual Euclidean distance between all pairing matched user/workers:
-$\text{travel costs}=\sum\nolimits_{(i,j)\in G_{a}\times G_{b}}\llbracket M(i,j)>0\rrbracket\cdot\|l_{i}-l_{j}\|_{2},$
+LDP randomizers We use the Minkowski response for location randomization.
+As a comparison, in the local model of DP (e.g., in [85, 80, 84]), we compare with existing mechanisms including Laplace [31], Staircase [38], PlanarLaplace with geo-indistinguishability [2], SquareWave [58], PrivUnit [12] and its Gaussian variant PrivUnitG in [6].
+Server-side algorithms Two commonly used server-side algorithms are evaluated as concrete tasks: minimum weighted full matching [52] and maximum matching [47].
+The two algorithms have different optimization objectives, thus later we will show the results for task-specific utility for each of them in addition to single report errors.
+The minimum weighted full matching aims to minimize the overall traveling costs between users and workers.
+The maximum matching aims to maximize the number of successfully matched user/worker pairs, where users outside the workers’ serving radius ($0.4$) are deemed unreachable.
+Experimental results We first present the single report errors, quantified by the expected $\ell_{2}$ distance between the reported location and the true location.
+In Fig.
+5 (a) and (b), we compare the single report errors within the LDP model.
+The Minkowski randomizer’s error is on par with other state-of-the-art mechanisms when $\epsilon\leq 1.0$ and significantly lower when $\epsilon\geq 2.0$.
+Fig.
+5 (c) and (d) illustrate the single report errors in the PIC model.
+For both datasets, the Minkowski randomizer performs better, with its utility advantage increasing as the global privacy budget $\epsilon_{c}$ grows.
+Additionally, it is evident that the error is generally higher for all randomizers in the LDP model compared to the PIC model.
+This discrepancy is due to the lack of privacy amplification in the LDP model, highlighting the benefits of employing the PIC model.
+Figure 5: Expected $\ell_{2}$ distances of reported locations to true locations on GMission and EverySender dataset.
+Next, we compare the utility when using the minimum weight matching algorithm, as shown in Fig.
+6.
+Task-specific utility is evaluated by the total travel costs, which is the sum of the actual Euclidean distances between all matched users/workers:
+$\sum\nolimits_{(i,j)\in G_{a}\times G_{b}}\llbracket M(i,j)>0\rrbracket\cdot\|l_{i}-l_{j}\|_{2},$
 […]
-$\text{success ratio}=\frac{\sum_{(i,j)\in G_{a}\times G_{b}}\llbracket M(i,j)>0\rrbracket\cdot\llbracket\|l_{i}-l_{j}\|_{2}\leq\tau\rrbracket}{\min\{|G_{a}|,|G_{b}|\}}.$
-We evaluate both the client-side and server-side running time of our protocol on a laptop computer embedded with Intel i5555-8250U CPU @1.6GHz and $8$GB memory.
-Figure 5: Expected $\ell_{2}$ error of reported locations.
-Performance of location reports.
-Before exploring spatial crowdsourcing tasks, we first evaluate our proposed Minkowski response mechanism against existing mechanisms, as presented in Table II.
-This evaluation focuses on errors and running times associated with reporting locations under local $\epsilon$-DP constraints.
-We quantify the error using the expected $\ell_{2}$ distance between each true and noisy location pair.
-Notably, the Minkowski response showcases an $\ell_{2}$ error comparable to state-of-the-art (SOTA) mechanisms when $\epsilon\leq 1.0$.
-Its superiority is evident when $\epsilon\geq 2.0$, displaying a significantly reduced error.
-This supports the theoretical assertions made in Section VI, underscoring the importance of the Minkowski response for the shuffle model, over other existing randomizers.
-Additionally, the Minkowski response is highly efficient, demanding less than $1$us on the user end.
-This speed is nearly equivalent to merely adding Laplace noises and is roughly $50$x faster than the previously SOTA PrivUnit.
-We further contrast the utility performances of the mechanisms in the local model with those in the shuffle model.
-Privacy amplification in the shuffle model is achieved using the latest analyses from [32].
-As illustrated in Figure 5, the shuffle model presents a significantly reduced location reporting error compared to the local model.
-This disparity becomes more pronounced as the number of users increases (notably, the EverySender dataset comprises more users than GMission).
-The introduction of an apt randomizer—such as the Minkowski response—can further elevate utility.
-This is evident when comparing the Minkowski response and PrivUnit, both of which operate within the shuffle model.
-Performance of minimum weight matching.
-Figure 6 presents the expected travel costs associated with minimum weighted full matching.
-For a comprehensive overview of the experimental results involving more competitors, readers are directed to -F.
-Notably, the Squarewave mechanism outperforms PrivUnit in matchings, prompting its primary representation in the shuffle model discussions.
-The equivalent shuffle model yields a cost reduction of approximately $50\%$, with the savings escalating as the number of users increases.
-The Minkowski response consistently exhibits the lowest costs across all scenarios.
-Together, they result in a substantial reduction of about $75\%$ when compared to existing approaches.
-Figure 6: Travel costs of minimum weighted matching.
-Performance of maximum matching.
-Figure 7 illustrates the success ratios for maximum matching.
-The equivalent shuffle model shows a relative improvement of approximately $200\%$ in the success ratio.
-The Minkowski response consistently yields higher ratios across all scenarios.
-When the user count becomes relatively substantial (for instance, a few thousands as in EverySender), our results indicate that spatial crowdsourcing maximum matching, with a reasonable level of privacy protection (specifically, $\epsilon\leq 3$), can be made practical, achieving a success ratio exceeding $90\%$.
-Figure 7: Success ratios of maximum matching.
-TABLE IV: Running time and communication overheads of spatial crowdsourcing in the PIC model.
-User-side Procedures Time Communication Server-side Procedures Time Communication
-location randomization 0.7us - decrypt one message 2.1ms -
-encrypt information 2.9ms - min-weight match (GMission) 31ms 1.2MB
-send message to shuffler - 1.3KB maximum match (GMission) 15ms 1.2MB
-retrieve matching result - 1.8KB min-weight match (EverySender) 79ms 4.3MB
-key agreement with worker 2.6ms - maximum match (EverySender) 63ms 4.3MB
-Running time and communication overheads.
-We present the computation and communication overheads of the PIC model in Table IV.
-When considering multiple procedures on the user side, the total running time for each user is only a few milliseconds (with the data randomization time being negligible), and the total communication overhead is under $4$KB.
-On the server side, the running time of the matching algorithm on noisy plaintext is consistent with non-private settings, scaling with the number of users/workers and taking dozens of milliseconds.
-The additional decryption/encryption runtime is capped at several seconds, mirroring the costs of HTTPS decryption/encryption.
-Specifically, it utilizes ECDHE for key agreement and AES encryption as with HTTPS on TLS 1.3 [67] for client-server communication.
-In essence, the running times on both the client and server sides of the PIC model closely resemble those in non-private settings that use HTTPS communication (for a fair comparison).
+Figure 6: Travel costs of minimum weighted matching on GMission and EverySender dataset.
+Finally, we present the utility comparison using the maximum matching algorithm, as shown in Fig. 7. In this case, task-specific utility is assessed by the successful matching ratio:
+$\frac{\sum_{(i,j)\in G_{a}\times G_{b}}\llbracket M(i,j)>0\rrbracket\cdot\llbracket\|l_{i}-l_{j}\|_{2}\leq\tau\rrbracket}{\min\{|G_{a}|,|G_{b}|\}}.$
+For both datasets, the matching ratio over clear data is 100%.
+The figure demonstrates that the PIC model enhances the matching ratio for all randomizers due to privacy amplification effects.
+The Minkowski randomizer in the PIC model significantly outperforms the others and the matching ratio approaches an acceptable level for practical use with a reasonable degree of privacy protection.
+Figure 7: Success ratios of maximum matching in spatial crowdsourcing on GMission and EverySender dataset.
 VII-B Location-based Social Systems
-In this part, we evaluate the performance of PIC model for privacy preserving nearest neighbor (NN) queries on users, a prevalent task in location-based social networks [22].
-Like the exemplar spatial crowdsourcing tasks, the nearest location senders’ information (including their corresponding public keys) w.r.t.
-$\ell_{2}$-distance) of a user is securely return to the user.
-These matched neighboring users can then securely communicate with other to accomplish subsequent tasks.
-TABLE V: Statistics of location social network datasets.
-Dataset users check-ins location domain
-Gowalla 138368 6442892 $[37.54,37.79]\times[-122.51,-122.38]$
-Foursquare 1083 227428 $[40.55,40.99]\times[-74.27,-73.68]$
-Datasets and settings.
-We use use two real check-in datasets, Gowalla dataset [22] and Foursquare dataset [87].
-The statistics of two datasets are summarized in Table V.
-The location domains are rescaled to the standard domain $[-1,1]\times[-1,1]$.
-As for nearest neighbor queries, we set the nearest radius $\tau$ for Gowalla to $0.1$ and the nearest radius $\tau$ for Foursquare to $0.2$, so that users have a few tens or hundreds neighbors (vary due to check-ins densities).
-Accordingly, the population size for privacy amplification in PIC model is set to $n-1000$.
-The concrete cryptography implementations and data randomizers for nearest neighbor queries are the same as the one for spatial crowdsourcing, while the server simply runs a radius-based nearest neighbor algorithm and return the neighboring (public key, noisy location) lists to each user.
-Performance metrics.
-We evaluate the performance of nearest neighboring queries in terms of precision and recall, and $F_{1}$ score of neighboring lists.
-Let $N_{i}$ denote the true neighbor lists that have $\ell_{2}$ distance to user $i$ less than the parameter $\tau$, and let $\widehat{N}_{i}$ denote the retrieved neighbor lists that have $\ell_{2}$ distance to user $i$ based on noisy (sanitized) locations less than the parameter $\tau$.
-The precision of nearest neighboring queries is defined as $\text{precision}={(\sum_{i\in[n]}|N_{i}\bigcap\widehat{N}_{i}|)}/{(\sum_{i\in[n]}|N_{i}|)}$, the recall of nearest neighboring queries is $\text{recall}={(\sum_{i\in[n]}|N_{i}\bigcap\widehat{N}_{i}|)}/{(\sum_{i\in[n]}|\widehat{N}_{i}|)}$.
-We use a balanced indicate
-$F_{1}\text{ score}=2\cdot\text{precision}\cdot\text{recall}/(\text{precision}+\text{recall})$
-as the major measurement. All presented results are the average value of $10$ experimental runs.
-Experimental results.
-Figure 8 displays the $F_{1}$ scores for radius nearest neighbor queries.
-To evaluate the effects of varying user numbers, we adjusted the number of users $n$ in the Gowalla dataset from $10000$ to $40000$.
-Notably, as $n$ surpasses $50000$, the PIC model’s scores see an approximate relative increase of $100\%$ compared to the local model.
-The Minkowski response consistently records the highest scores.
-When used in conjunction, they can achieve up to $90\%$ of the $F_{1}$ scores within a reasonable privacy loss (e.g., $\epsilon\leq 3$), especially when $n$ is relatively large.
-Figure 8: F1 scores of radius nearest neighbor queries in location-based social systems.
+TABLE III: Details of location social network datasets.
+Dataset check-ins location domain
+Gowalla (San Francisco) 138368 $[37.54,37.79]$$\times$$[-122.51,-122.38]$
+Foursquare (New York) 227428 $[40.55,40.99]$$\times$$[-74.27,-73.68]$
+Datasets We use two real-world datasets: Gowalla dataset [21] and Foursquare dataset [89].
+Gowalla and Foursquare are location-based social network websites where users share their locations by checking-in.
+Details about the two datasets are summarized in Table III.
+As before, the location data is normalized to $[-1,1]\times[-1,1]$.
+LDP randomizers The LDP randomizers used are the same as those in Section VII-A.
+Server-side algorithm The server performs the radius-based nearest neighbor (NN) search for the users, which is a common task in location-based social networks [21].
+We set the search radius $\tau$ to 0.2, so that each user has several hundreds or thousands of neighbors, varying due to check-in densities.
+Note that in this application scenario, the returned neighbors may be deanonymized in the post-computation phase.
+Hence, the actual privacy guarantee in the PIC model depends on the number of users who remain anonymous (see discussion in Section V-C).
+Considering this, and each user normally has no more than $n\cdot\frac{\pi\tau^{2}}{2^{2}}<n\cdot 3.2\%$ neighbors, we use privacy amplification population $\lfloor n\cdot 90\%\rfloor$ instead of the group size $n$ when calculating the local budget given the global $\epsilon_{c}$.
+Figure 8: Expected $\ell_{2}$ distances of reported locations in location-based social systems with the local model of DP.
+Figure 9: Expected $\ell_{2}$ distances of reported locations in location-based social systems with PIC model.
+Experimental results We first present the single report errors in the LDP model (Fig.
+8) and the PIC model (Fig.
+9).
+In our experiments, each check-in report is treated as if it were submitted by a separate user, with all users in the same group.
+We tested two scenarios: one with a random subset of 10,000 check-ins and the other using all check-ins.
+On both datasets, PIC demonstrates better utility than LDP, and the Minkowski randomizer consistently performs the best across all settings.
+Additionally, it is evident that the number of anonymous users is a crucial parameter, significantly impacting utility.
+Figure 10: F1 scores of nearest neighbor queries (LDP model).
+Figure 11: F1 scores of nearest neighbor queries (PIC model).
+The task-specific utility metric we use for nearest neighbor queries is the $F_{1}$ score.
+Let $N_{i}$ denote the set of true neighbors of user $i$ within an $\ell_{2}$-distance $\tau$, and let $\widehat{N}_{i}$ denote the retrieved neighbor set computed using the sanitized reports.
+The precision of nearest neighbor queries is defined as $\frac{\sum_{i\in[n]}|N_{i}\cap\widehat{N}_{i}|}{\sum_{i\in[n]}|\widehat{N}_{i}|}$, and the recall is defined as $\frac{\sum_{i\in[n]}|N_{i}\cap\widehat{N}_{i}|}{\sum_{i\in[n]}|N_{i}|}$.
+The $F_{1}$ score is then calculated as:
+$F_{1}\text{ score}=2\cdot\frac{\text{precision}\cdot\text{recall}}{\text{precision}+\text{recall}}.$
+In the LDP model (Fig.
+10), the $F_{1}$ score is low, even with a large $\epsilon$.
+While the group size affects the number of neighbors each user has, it does not impact the $F_{1}$ score.
+Conversely, in the PIC model (Fig.
+11), the $F_{1}$ score is significantly higher.
+With a large group, the Minkowski randomizer can achieve an $F_{1}$ score of 0.9 with a stringent global budget $\epsilon_{c}=1$ and nearly 1 if the budget is loosened to $\epsilon_{c}=3$.
 VII-C Federated Learning with Incentives
-Recall that in federated learning with rigorous privacy protection, many clients collaboratively contribute gradient updates of model parameters so as to build a machine learning model.
-Each client $i$ holds a local dataset $D_{i}$.
-In each iteration $t$, a client $i$ downloads the latest model parameter $W_{t-1}$, and computes the local gradient $g_{i}$ based on local dataset $D_{i}$.
-Then, the client sanitizes $g_{i}$ and sends it to the shuffler/server.
-Without loss of generality, we assume the server uses Shapley value [69] to determine the gradient payoffs, where the utility function for computing Shapley value is based on the model parameters’ performance (e.g., accuracy or the negative of loss function) on a validation dataset $D_{val}$.
-The Shapley values themselves may severely leak sensitive information about user data [59].
-Despite the vast non-linear computations involved in evaluating Shapley values with neural networks, no rigorous protection has been offered for federated learning with incentives in existing literature.
-In this study, we demonstrate the feasibility of computing Shapley value without incurring additional privacy loss.
-TABLE VI: Model architecture with 4292 trainable parameters.
+In this application, the users collaboratively train a model by Federated learning, and the server decides each user’s incentive by how much their local gradient contributes to the global model.
+To deliver monetary incentive rewards, we assume the use of an untraceable cryptocurrency (e.g., ZCash [47]).
+Each user can generate an additional public/privacy key pair according to the specification of the cryptocurrency, derive a wallet address from the public key, and append the wallet address to the end of their report.
+Based on the Shapley values, the server determines the monetary incentives and distributes them to the users using appended wallet addresses.
+Datasets We utilize the MNIST handwritten digit dataset to train a simple neural network using federated learning.
+The MNIST dataset comprises 60,000 images, with 50,000 designated as training samples.
+Each user receives one training sample and trains a neural network model, as detailed in Table IV, which contains $d=4292$ trainable parameters.
+In each round, $s=10,000$ users are randomly selected.
+These selected users locally compute the gradient vector using Stochastic Gradient Descent (SGD), then subsample 0.15% of the gradient vector dimensions (with unsampled dimensions set to zero), and clip the gradient values at a threshold of $c=0.00015$.
+Each user submits a sanitized version of the subsampled, clipped gradient vector as their report.
+TABLE IV: Model architecture of the neural network.
 Layer Parameters
@@
 Softmax 10 units
-Datasets and settings.
-Our experiments with federated learning use the MNIST handwritten digit dataset, comprising $60,000$ training, $5,000$ validation, and $5,000$ test images.
-We employ a neural network outlined in Table VI with $d=4292$ trainable parameters.
-It achieves $98.6\%$ accuracy using non-private, untruncated vanilla SGD and $92.7\%$ accuracy with non-private, truncated SGD.
-For our experiments, we simulate $50,000$ clients with each client possessing one sample (i.e., $n=50,000$).
-During each private SGD step, $s=10,000$ clients are randomly selected.
-Each client subsamples $0.5\%$ of the gradient vector dimensions, then clips the $\ell_{1}$-norm of the block gradient with a threshold of $c=0.001$.
-We compare our PIC model against a local model, both implementing Laplace mechanism for gradient randomization.
-Implementation details.
-We use Elliptical curve encryption to perform secure point-to-point information delivery.
-Specifically, to delivering monetary incentives with anonymity protection, clients and server may employ untraceble crypto-currency (e.g., ZCash [46]).
-Each client creates one random temporary wallet address and append it to the message, the server computes incentive payments of the message and then send corresponding amount of coins to the wallet address.
-Performance metrics.
-We assess performances using gradient/model accuracy and Shapley value accuracy.
-Gradient accuracy is gauged using the mean squared error (MSE) between the subsampled, clipped, and privatized gradients $\hat{g}_{i}$ and the corresponding non-private gradients ${g}_{i}$:
-$\text{gradient}\ MSE=\|\sum\nolimits_{i\in[s]}\hat{g}_{i}-\sum\nolimits_{i\in[s]}{g}_{i}\|_{2}^{2}.$
-Model accuracy refers to the image classification accuracy on the test dataset.
-Shapley value accuracy is the MSE between Shapley values of privatized gradients $\hat{g}_{i}$ and non-private gradients ${g}_{i}$:
-$\text{shapley}\ MSE=\sum\nolimits_{i\in[s]}(\widehat{\texttt{Shapley}}_{i}-{\texttt{Shapley}}_{i})^{2},$
-where the Shapley values for privatized gradients use the utility function (see Appendix -H for detail):
-$\widehat{U}(S)=\frac{<\texttt{grad}_{val},\sum_{i\in S}\hat{g}_{i}>}{\|\texttt{grad}_{val}\|_{2}\cdot\|\sum_{i\in S}\hat{g}_{i}\|_{2}}.$
-The presented gradient/shapley MSE averages over $50$ aggregation rounds, and the model accuracy is from the final round.
-Experimental results.
-Since the curator model of DP and other cyrptographic MPC approaches (e.g., secure aggregation [18]) are not applicable to privacy preserving joint gradient aggregation and Shapley value computation (with complicate non-linear computation), we compare equivalent shuffle model with the local model of DP.
-Both models use Laplace noises [28] to privatizing local gradients, and use tight numerical composition for sequential gradient averaging [42].
-We present the experimental results in Table VII, with varying overall privacy budget $\epsilon$.
-Every result is the average value of $5$ simulations.
-TABLE VII: Utility performance of joint gradient aggregation and Shapley-value incentive for federated learning with differential privacy guarantee on MNIST dataset.
+LDP randomizers We use all LDP randomizers as before except for PlanarLaplace, which is unsuitable for multi-dimensional gradient vectors.
+The randomizers sanitize the non-zero values in the subsampled gradient vectors while leaving the zero values unaffected.
+Server-side algorithm The server performs the following steps: first, it takes a sanitized gradient vector and uses randomizer-specific algorithms to estimate the true values of the sampled dimensions.
+These vectors with estimated values are denoted as $\hat{g}_{i}$.
+The server then aggregates the $\hat{g}_{i}$ vectors into a global gradient vector using a simple sum-and-average method.
+This global gradient vector is published so that users can update their local models.
+Additionally, the server computes the Shapley value for each user, which measures the marginal contribution of each $\hat{g}_{i}$ (see Appendix -H for details).
+Experimental results We train each model for 80 rounds using Federated Learning.
+The utility comparison results are summarized in Tables V and VI, each reflecting a different global privacy budget.
+In these tables, we present the single report utility (Gradient $\ell_{2}$ error) and the task-specific utility (Shapley $\ell_{2}$ error).
+The Gradient $\ell_{2}$ error is calculated using the estimated gradients $\hat{g}_{i}$ and the true gradients $g_{i}$ (clipped and with all unsampled dimensions set to 0): $\frac{\sum_{i\in S}\|\hat{g}_{i}-{g}_{i}\|_{2}}{|S|}$, where $S$ is the set of randomly selected users.
+The Shapley $\ell_{2}$ error is computed as follows:$\frac{\sum\nolimits_{i\in S}\|\widehat{\texttt{Shapley}}_{i}-{\texttt{Shapley}}_{i}\|_{2}}{|S|}$, where $\widehat{\texttt{Shapley}}_{i}$ is derived from $\hat{g}_{i}$ and ${\texttt{Shapley}}_{i}$ is obtained from the unsanitized $g_{i}$.
+For reference, we also include the final models’ accuracy in the tables.
+TABLE V: Utility comparison of federated learning with incentives: global privacy budget $(1,0.01/50000)$.
 Setting
-Privacy
-Budget
+Randomization
+Mechanism
 Test
@@
 Gradient
-MSE
+$\ell_{2}$ Error
 Shapley
-MSE
-local model $\epsilon=1$ $19.1\%$ $1.7$e-4 $8.9$e-3
-$\epsilon=3$ $28.9\%$ $1.3$e-4 $6.1$e-3
-$\epsilon=5$ $35.6\%$ $1.1$e-4 $5.2$e-3
-PIC model $\epsilon=1$ $77.6\%$ $2.5$e-5 $4.6$e-3
-$\epsilon=3$ $81.2\%$ $1.4$e-5 $4.2$e-3
-$\epsilon=5$ $85.1\%$ $1.1$e-5 $3.9$e-3
-It is observed that PIC model enjoys significantly better gradient accuracy, increase model test accuracy by about $50\%$ in the absolute sense, reduces about $90\%$ gradient aggregation error, when compared to the local model.
-The PIC model also reduces about $40\%$ error on Shapley values, permit accurate payoff distribution to incentive honest/informative participants.
+$\ell_{2}$ Error
+local model Staircase [38] $18.29\%$ $15.30$ $0.0586$
+Squarewave [58] $13.91\%$ $13.23$ $0.0590$
+PrivHS [29] $28.17\%$ $2.44$ $0.0587$
+Laplace [31] $21.22\%$ $2.53$ $0.0575$
+PrinUnit [12] $23.28\%$ $2.37$ $0.0592$
+Minkowski $18.10\%$ $14.45$ $0.0571$
+PIC model Staircase [38] $32.51\%$ $0.781$ $0.0563$
+Squarewave [58] $33.75\%$ $0.604$ $0.0585$
+PrivHS [29] $65.47\%$ $0.267$ $0.0428$
+Laplace [31] $72.6\%$ $0.154$ $0.0508$
+PrinUnit [12] $74.02\%$ $0.157$ $0.0438$
+Minkowski $\mathbf{78.68\%}$ $\mathbf{0.093}$ $\mathbf{0.0363}$
+TABLE VI: Utility comparison of federated learning with incentives: global privacy budget $(3,0.01/50000)$.
+Setting
+Randomization
+Mechanism
+Test
+Accuracy
+Gradient
+$\ell_{2}$ Error
+Shapley
+$\ell_{2}$ Error
+local model Staircase [38] $19.68\%$ $5.09$ $0.0519$
+Squarewave [58] $18.24\%$ $4.34$ $0.0574$
+PrivHS [29] $41.52\%$ $0.850$ $0.0561$
+Laplace [31] $28.64\%$ $0.845$ $0.0575$
+PrinUnit [12] $39.02\%$ $0.803$ $0.0537$
+Minkowski $20.73\%$ $3.731$ $0.0545$
+PIC model Staircase [38] $44.58\%$ $0.545$ $0.0563$
+Squarewave [58] $53.02\%$ $0.407$ $0.0552$
+PrivHS [29] $74.49\%$ $0.190$ $0.0383$
+Laplace [31] $74.86\%$ $0.104$ $0.0335$
+PrinUnit [12] $77.42\%$ $0.098$ $0.0253$
+Minkowski $\mathbf{83.43\%}$ $\mathbf{0.055}$ $\mathbf{0.0219}$
+We observe in the tables that the utility in the LDP model is significantly worse than in the PIC model, as expected.
+This is consistent across all metrics and the final model accuracy.
+In the local model, the performance of Staircase, Squarewave, and Minkowski mechanisms is poorer compared to the others, but for different reasons.
+Staircase and Squarewave are designed for single-dimensional data, requiring the privacy budget to be split across dimensions when sanitizing vectors, which leads to reduced utility.
+For Minkowski, the issue lies in the 80-round federated learning process, where the privacy budget for each round is relatively small (e.g., between 0.2 and 0.75).
+As previously mentioned, the Minkowski mechanism is intended to achieve better utility with a large local privacy budget.
+In small privacy budget scenarios, it offers no advantage and may even perform worse.
+In the PIC model, the performance of Staircase and Squarewave remains poor.
+However, Minkowski now performs the best among all randomizers.
+This improvement is due to the privacy amplification effect in PIC, which increases the local privacy budget for each round to between 4.0 and 6.0.
 VIII Conclusion
-In this paper, we proposed private individual computation model, for multi-party individual computation with DP preservation, having broad application to spatial crowdsourcing, resource allocation, location-based systems, and federated learning with incentives.
-The paradigm introduces statistically random identities into the traditional shuffle model of DP, so as to support secure functionalities required by individual computations, meanwhile maintaining the most benefits of privacy amplification via shuffling.
-It incurs only $O(d)$ computation and communication costs (matching the non-private setting) on the client side, meanwhile permitting plaintext computation on the server side.
-Additionally, with the increasing of number of participating parties, the utility of individual computation grows (as there are stronger privacy amplification effects), making it suitable for large-scale applications.
-The simple paradigm can efficiently solves prevalent tasks that are computationally intractable in MPC approaches (e.g., Ishai’s anonymous model, secure aggregation, or garbled circuits), such as federated deep learning with incentives and large-scale spatial crowdsourcing, necessitating essentially the same overheads as in non-private settings.
-Moreover, it reduces $40\%$-$90\%$ errors/costs, increases $100\%$-$300\%$ utility, when compared to existing tractable approaches (e.g., local DP).
-Discussions and future research.
-The PIC model guarantees privacy gracefully in cases of malicious clients or collusion, and can adapt to fully decentralized settings without an orchestrating server.
-Given its alignment with random digital identities in Blockchains and in Ad tracking systems [54], it offers potential to enhance formal privacy in these systems.
+Privacy-preserving computation with Differential Privacy (DP) holds great potential for leveraging personal information.
+While the shuffle model offers a rigorous DP guarantee with enhanced utility, its application is confined to statistical tasks.
+In this paper, we introduce a novel paradigm called Private Individual Computation (PIC), which extends the shuffle model to scenarios where each user requires personalized outputs from the computation.
+We demonstrate that PIC can be realized using an efficient protocol that relies on minimal cryptographic operations while maintaining the advantages of privacy amplification through shuffling.
+To further enhance utility, we developed a local randomizer specifically designed for PIC.
+We provide formal proofs of the protocol’s security and privacy, as well as the asymptotic optimality of the randomizer.
+Extensive experiments validate the superiority of the PIC protocol and the randomizer, showcasing their performance across three major application scenarios and various real-world datasets.

### v4

@@
 In data-driven applications, preserving user privacy while enabling valuable computations remains a critical challenge.
-Technologies like Differential Privacy (DP) have been pivotal in addressing these concerns.
+Technologies like differential privacy have been pivotal in addressing these concerns.
 The shuffle model of DP requires no trusted curators and can achieve high utility by leveraging the privacy amplification effect yielded from shuffling.
 These benefits have led to significant interest in the shuffle model.
-However, the computation tasks in the shuffle model are limited to statistical estimation, making the shuffle model inapplicable to real-world scenarios in which each user requires a personalized output.
+However, the computation tasks in the shuffle model are limited to statistical estimation, making it inapplicable to real-world scenarios in which each user requires a personalized output.
 This paper introduces a novel paradigm termed Private Individual Computation (PIC), expanding the shuffle model to support a broader range of permutation-equivariant computations.
@@
 Theoretical analysis and empirical evaluations demonstrate PIC’s capability in handling non-statistical computation tasks, and the efficacy of PIC and the Minkowski randomizer in achieving superior utility compared to existing solutions.
-Index Terms:
-Differential Privacy, Private Individual Computation, Shuffle Model, Privacy-Utility Trade-offs
-I Introduction
+1 Introduction
 Personal information fuels a wide array of data-driven applications, e.g.
@@
 However, LDP’s primary drawback is that each user must independently add sufficient random noise to ensure privacy, which can significantly impact the utility of the data.
-Recently, within the realm of DP research, the shuffle model [13, 32] has emerged.
-This model introduces a shuffler that randomly permutes messages from users and then sends these anonymized messages to a computing server or analyzer.
-The server can then compute on these messages to derive the result.
-Trust in the shuffler is minimized because each user encrypts their locally sanitized data using the server’s public key.
-This way, the shuffler remains oblivious to the messages received from the users.
-A key advantage of the shuffle model over LDP is utility.
-It has been proven that anonymizing/shuffling messages amplifies the privacy guarantee provided by the local randomizer used by the users.
-For instance, shuffled messages from $n$ users each adopting local $\epsilon$-DP actually preserve differential privacy at the level $\epsilon_{c}=\tilde{O}(\sqrt{e^{\epsilon}/n})$ [34, 35].
-Consequently, to achieve a predefined global privacy goal, less noise must be added when users sanitize their data locally.
-This significantly improves the accuracy of the final result.
-Owing to these utility advantages, extensive studies have been conducted within the shuffle model, e.g., [8, 40, 41, 41, 42].
+[…]
 That said, the shuffle model has a noticeable limitation.
@@
 Examples of individual computation tasks include:
-  • I.
   • Combinatorial optimization: Spatial crowdsourcing [87], advertisement allocation [66], and general combinatorial optimization [59], where two or more parties are often matched together based on their private information.
 Each party should get their own list of “best matches” whatever that means.
-  • II.
   • Information retrieval: Mobile search [56], location-based systems [2, 21], where the query results (e.g., nearby restaurants or neighboring users) depend on the private information of the inquirer.
-  • III.
-    Incentive mechanisms: In federated learning [92] or data crowdsensing [74], incentives play a vital role to encourage well-behaved participation.
-The amount of rewarding incentive must be computed for individual users based on their contribution (e.g., via Shapley values [71]).
+  • Incentive mechanisms: In federated learning [97] or crowd sensing [78], incentives play a vital role to encourage well-behaved participation.
+The amount of rewarding incentive must be computed for individuals based on their contribution (e.g., via Shapley values [75]).
 At first glance, private individual computation appears unattainable within the shuffle model because the need for personalized output conflicts with the anonymization required for privacy amplification.
@@
 the one described in the seminal work [13], do not support such duplex connections.
-Moreover, after receiving the computation results, many individual computation tasks require establishing party-to-party communication (e.g., a user communicates with the matched driver in taxi-hailing services, a user communicates with matched near users in social systems), which is intractable in the duplex shuffle channel.
+Moreover, after receiving the computation results, many individual computation tasks require establishing party-to-party communication (e.g., a user communicates with the matched driver in taxi-hailing services, a user communicates with matched near users in social systems), which can be hard to implement in the duplex shuffle channel.
 Addressing this issue is the first technical challenge we need to overcome.
@@
 However, the new setting of individual computation is different: the focus is on the accuracy of the output for each user rather than the statistical accuracy of the population.
-This difference renders existing randomization strategies (e.g., randomizers utilizing dimension sampling, budget splitting, or data sketching, as reviewed in [88]), as well as prevalent randomizers (such as adding Laplace [31] or Staircase [38] noises), less effective for the new setting.
+This difference renders existing randomization strategies (e.g., randomizers utilizing dimension sampling, or budget splitting, as reviewed in [93]), as well as prevalent randomizers (such as adding Laplace noise [31]), less effective for the new setting.
 Therefore, we need to reconsider the fundamental privacy-utility trade-offs and redesign the underlying randomizers to better suit individual computation requirements.
@@
 The randomizer’s output domain is defined as a ball extending the input domain’s radius by $r$.
-The key to achieving high utility lies in the randomizer’s output: it selects a value close to the true value (within $r$) with a relatively high probability and a value from elsewhere in the entire output domain with a relatively low probability.
+The key to achieving high utility lies in the randomizer’s output: it selects a value close to the true value (within $r$) with a relatively high probability and a value from elsewhere with a relatively low probability.
 We formally prove the security and privacy properties of the PIC protocol.
@@
 The remainder of this paper is organized as follows.
-Section II reviews related works.
-Section III provides preliminary knowledge about privacy definitions and security primitives.
-Section IV formalizes the problem setting.
-Section V presents the PIC protocol.
-Section VI provides optimal user-side randomizers.
-Section VII evaluates the utility and efficiency performances of our proposals.
-Finally, Section VIII concludes the paper.
-II Related Works
+Section 2 reviews related works.
+Section 3 provides preliminary knowledge about privacy definitions and security primitives.
+Section 4 formalizes the problem setting.
+Section 5 presents the PIC protocol.
+Section 6 provides optimal randomizers.
+Section 7 evaluates the performances of our proposals.
+Section 8 discusses more merits and future directions of the PIC.
+Finally, Section 9 concludes the paper.
+2 Related Works
 This section reviews various approaches to private computation, primarily concentrating on non-statistical tasks.
-II-A Secure Multiparty Computation
+2.1 Secure Multiparty Computation
 Secure Multiparty Computation (MPC) is a fundamental cryptographic primitive that enables multiple parties to jointly compute a function over their inputs while no party learns anything beyond their own input and the final output of the computation.
@@
 Despite a significant improvement, MPC still faces efficiency issues that hinder its widespread real-world deployment.
-II-B Curator and Local DP Methods
+2.2 Curator and Local DP Methods
 Many works study matching, allocation, or general combinatorial optimization problems within the curator DP model [31] in the presence of a trusted party collecting raw data from clients (e.g., in [65, 22, 83, 82]).
@@
 As each client must injects sufficient noises into data to satisfy local DP, the execution results often maintain low utility.
-II-C Shuffle Model of DP
+2.3 Shuffle Model of DP
 The recently proposed shuffle model [13, 32] combines the advantages of the curator model (e.g., high utility) and the local model (e.g., minimal trust).
@@
 Overall, existing works in the shuffle model primarily focus on statistical queries.
-This work, for the first time, explore the shuffle model for non-statistical applications (i.e., combinatorial optimization, location-based social systems, and incentive mechanisms).
-II-D Combining Cryptography and DP
+This work, for the first time, explore the shuffle model for non-statistical applications (i.e., combinatorial optimization, social systems, and incentive mechanisms).
+2.4 Combining Cryptography and DP
 While cryptographic tools can protect data secrecy during multiparty computation, they do not necessarily preserve output’s privacy.
@@
 To account for privacy loss due to intermediate encrypted views in MPC, researchers have proposed the relaxed notion of computational DP [67] against polynomial-time adversaries.
-Computational DP protocols often inherited the computation/communication complexity of MPC (refer to approaches in Table IX).
-III Preliminaries
-In this section, we provide a concise introduction to the preliminaries. A list of notations used throughout the paper can be found in Table I.
-TABLE I: List of notations.
+Computational DP protocols often inherited the computation/communication complexity of MPC (refer to approaches in Table 9 of the full version [91]).
+3 Preliminaries
+In this section, we provide a concise introduction to the preliminaries. A list of notations can be found in Table 1.
+Table 1: List of notations.
 Notation Description
@@
 $u_{i,j}$ the $j$-th user in group $G_{i}$ where $j\in[n_{i}]$
-$d$ the dimension of user’s data
-$\mathbb{X}$ the domain of input data
-$\mathbb{Y}$ the domain of a sanitized message
 $\epsilon$ the local privacy budget
 $\epsilon_{c}$ the amplified privacy level
-$sk$, $pk$ the secret key and the public key, respectively
+$sk$, $pk$ the secret key and public key, respectively
 $\lambda$ the security parameter of cryptography
-III-A Privacy Definitions
-Definition III.1 (Hockey-stick divergence [72])
+3.1 Privacy Definitions
+Definition 3.1 (Hockey-stick divergence [76]).
 For two probability distributions $P$ and $Q$, the Hockey-stick divergence between them with parameter $e^{\epsilon}$ is as follows:
@@
 Differential privacy ensures that the Hockey-stick divergence between $\mathcal{R}(T)$ and $\mathcal{R}(T^{\prime})$ is bounded by a sufficiently small value (i.e., $\delta=O(1/n)$).
-The formal definition of curator differential privacy is as follows:
-Definition III.2 ($(\epsilon,\delta)$-DP [31])
+Definition 3.2 ($(\epsilon,\delta)$-DP [31]).
 A randomization mechanism $\mathcal{R}$ satisfies $(\epsilon,\delta)$-differential privacy iff $\mathcal{R}(T)$ and $\mathcal{R}(T^{\prime})$ are $(\epsilon,\delta)$-indistinguishable for any neighboring datasets $T,T^{\prime}\in\mathbb{X}^{n}$.
 That is, $\max(D_{\epsilon}(\mathcal{R}(T)\|\mathcal{R}(T^{\prime})),D_{\epsilon}(\mathcal{R}(T^{\prime})\|\mathcal{R}(T)))\leq\delta$.
-In the local model of differential privacy, each user applies a randomization mechanism $\mathcal{R}$ to their own data $x_{i}$, with the objective of ensuring that the Hockey-stick divergence between $\mathcal{R}(x)$ and $\mathcal{R}(x^{\prime})$ is 00 for any $x,x^{\prime}\in\mathbb{X}$ (see Definition III.3).
-Definition III.3 (local $\epsilon$-DP [54])
-A randomization mechanism $\mathcal{R}$ satisfies local $\epsilon$-DP iff $D_{\epsilon}(\mathcal{R}(x)||\mathcal{R}(x^{\prime}))=0$ for any $x,x^{\prime}\in\mathbb{X}$.
-The data processing inequality is considered a key feature of distance measures (e.g., Hockey-stick divergence) used for evaluating privacy.
-It asserts that the privacy guarantee cannot be weakened by further analysis of a mechanism’s output.
-Definition III.4 (Data processing inequality)
-A distance measure $D:\Delta(\mathcal{S})\times\Delta(\mathcal{S})\to[0,\infty]$ on the space of probability distributions satisfies the data processing inequality if, for all distributions $P$ and $Q$ in $\Delta(\mathcal{S})$ and for all (possibly randomized) functions $g:\mathcal{S}\to\mathcal{S^{\prime}}$,
-$D(g(P)\|g(Q))\leq D(P\|Q).$
-III-B The Classical Shuffle Model
+In the local DP model, each user applies a randomization mechanism $\mathcal{R}$ to their own data $x_{i}$, with the objective of ensuring that the Hockey-stick divergence between $\mathcal{R}(x)$ and $\mathcal{R}(x^{\prime})$ is 00 for any $x,x^{\prime}\in\mathbb{X}$ (see Definition 3.3).
+Definition 3.3 (local $\epsilon$-DP [57]).
+A mechanism $\mathcal{R}$ satisfies local $\epsilon$-DP iff $D_{\epsilon}(\mathcal{R}(x)||\mathcal{R}(x^{\prime}))=0$ for any $x,x^{\prime}\in\mathbb{X}$.
+3.2 The Classical Shuffle Model
 Following the conventions of the randomize-then-shuffle model [20, 7], we define a single-message shuffle protocol $\mathcal{P}$ as a list of algorithms $\mathcal{P}=(\mathcal{R},\mathcal{A})$, where $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ is local randomizer on client side, and $\mathcal{A}:\mathbb{Y}^{n}\to\mathbb{Z}$ is the analyzer on the server side.
@@
 In summary, the output of $\mathcal{P}(x_{1},\ldots,x_{n})$ is represented by $\mathcal{A}\circ\mathcal{S}\circ\mathcal{R}(X)=\mathcal{A}(\mathcal{S}(\mathcal{R}(x_{1}),\ldots,\mathcal{R}(x_{n})))$.
-In particular, when all users adopt an identical $\epsilon$-LDP mechanism $\mathcal{R}$, recent works [34, 35] have derived that $n$ shuffled $\epsilon_{l}$-LDP messages satisfy $(O((1-e^{-\epsilon})\sqrt{e^{\epsilon}\log(1/\delta)/n}),\delta)$-DP.
+In particular, when all users adopt an identical $\epsilon$-LDP mechanism $\mathcal{R}$, recent works [34, 35] have derived that $n$ shuffled $\epsilon$-LDP messages satisfy $(O((1-e^{-\epsilon})\sqrt{e^{\epsilon}\log(1/\delta)/n}),\delta)$-DP.
 We denote the amplified privacy level as:
 $\epsilon_{c}=\texttt{Amplify}(\epsilon,\delta,n),$
-the tight value of which can also be numerically computed [57, 87].
-III-C Public Key Encryption
+the tight value of which can also be numerically computed [60, 92]. Our experiments will use numerical bounds in [60].
+Figure 1: An illustration of taxi-hailing in the PIC model. Besides (sanitized) location information, each user also encapsulates a one-time random public key into the message to the shuffler.
+3.3 Public Key Encryption
 […]
@@
 The IND-CPA security ensures the scheme leaks no useful information under a chosen plaintext attack.
-IV Problem Settings
-We now introduce Private Individual Computation, a new paradigm for privacy-preserving computation that offers several benefits: (1) It provides a formal privacy guarantee (in the differential privacy sense) for a wide range of data-driven computational tasks.
-(2) Since the computation is performed on user-sanitized data, it avoids the need for heavy cryptographic protocols, allowing for scalability to handle large datasets and user bases.
-(3) The privacy amplification effect results in significantly better utility compared to direct sanitization and computation using Local Differential Privacy (LDP).
+4 Problem Settings
+We now introduce Private Individual Computation, a new paradigm for privacy-preserving computation that offers several benefits: (1) It provides a formal privacy guarantee (in the DP sense) for a wide range of computation tasks.
+(2) Since the computation is performed on user-sanitized data, it avoids the need for heavy cryptographic protocols, allowing for scalability to handle large user bases.
+(3) The privacy amplification effect results in significantly better utility compared to local DP.
 In this section, we will first present a few motivating applications and then formally define Private Individual Computation as an ideal functionality.
-Figure 1: An illustration of taxi-hailing in the PIC model. Besides (sanitized) location information, each user also encapsulates a one-time random public key into the message to the shuffler.
-IV-A Motivating Applications
+4.1 Motivating Applications
 We describe three prevalent exemplar computation tasks: spatial crowdsourcing, location-based social systems, and federated learning with incentives:
-Spatial crowdsourcing.
-A spatial crowdsourcing system typically consists of three roles: users, workers, and the orchestrating server (the service platform).
-It proceeds through the following major steps:
+Spatial crowdsourcing. A spatial crowdsourcing system typically consists of three roles: users, workers, and the orchestrating server. It proceeds through the following major steps:
   • I.
-    Task submission: each user $i\in G_{a}$ submits a task (e.g., taxi calling requests, sensing/photo requests) $x_{i}=(i,l_{i},v_{i})$ containing the location $l_{i}$ and possibly other information $v_{i}$;
+    Task submission: each user $i\in G_{a}$ submits a task (e.g., taxi calling requests, sensing requests) $x_{i}=(i,l_{i},v_{i})$ containing the location $l_{i}$ and possibly other information $v_{i}$;
   • II.
@@
   • I.
-    Querying: Each participating user $i\in G_{a}$ submits a request $x_{i}=(i,l_{i},v_{i})$, which includes information like location $l_{i}$ and preferences $v_{i}$.
+    Querying: each user $i\in G_{a}$ submits a request $x_{i}=(i,l_{i},v_{i})$, which might include location $l_{i}$ and preferences $v_{i}$.
   • II.
-    Generating recommendation: The server receives $\{x_{i}\}_{i\in G_{a}}$ and generates a list of recommendations for each user based on specific criteria, such as proximity and user preferences.
+    Generating recommendation: the server receives $\{x_{i}\}_{i\in G_{a}}$ and generates a list of recommendations for each user based on specific criteria (e.g., proximity and preferences).
   • III.
-    Retrieving: Finally, each user retrieves the recommendation results.
+    Retrieving: each user retrieves the recommendations.
 Federated learning with incentives. Federated learning involves a set of users and a server:
@@
     Computing incentive: The server computes the average gradient $\overline{x}=\frac{1}{n}\sum_{i\in[n]}x_{i}$.
-To incentive participation, the server may reward users with monetary tokens (e.g., via cryptocurrency) according to a profit allocation algorithm $V:[-1,1]^{d\times n}\times[-1,1]^{d}\mapsto[0,1]$ (e.g., Shapley value [71]).
+To incentive participation, the server may reward users with monetary tokens (e.g., via cryptocurrency) according to a profit allocation algorithm $V:[-1,1]^{d\times n}\times[-1,1]^{d}\mapsto[0,1]$.
   • II.
@@
 The shuffler knows the random permutation used in the shuffling process and will not leak it to other parties, although the shuffler is prevented from observing plaintext messages through encryption.
-IV-B The Ideal Functionality
+4.2 The Ideal Functionality
 Following the ideal-real world paradigm, we capture private individual computation formally as an ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$, which is shown in Figure 2.
+[…]
+Remark 1 A careful reader may notice that the ideal functionality does not explicitly capture differential privacy.
+This omission is intentional for the sake of security analysis.
+In the security analysis, we decouple the privacy requirements into two sets of proofs: the first set demonstrates that our concrete protocol, when executed by real-world parties, realizes the ideal functionality.
+This means no additional information about parties’ inputs is revealed, except for the output given to each party.
+The second set of proofs establishes that the output given to each party conforms with differential privacy.
 Functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ Parameters: $m\in\mathbb{N}$; $m$ groups of parties $G_{1},G_{2},\cdots,G_{m}$, where $n_{i}=|G_{i}|$ denotes the number of parties in group $G_{i}$; the data randomization mechanisms $\mathcal{R}_{i}$ for group $G_{i}$; a server ${S}$.
@@
 Figure 2: The functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$
-Essentially, the ideal functionality represents a fully trusted party that interacts with $m$ groups of users and one server $S$.
-It sanitizes the input from each user, shuffles the inputs randomly, and then applies a function to compute the output for each user.
-In the end, each user receives their individual function output, while the server receives a list of the shuffled sanitized user data and a list of the function outputs.
-It captures the functional requirements of individual computation in the real world: a server performs a computing task using the joint inputs from a set of users (which are sanitized and shuffled), and each user receives an individualized output.
-It also captures the security requirements: each party receives precisely the specified output, and nothing more.
-Remark 1 A careful reader may notice that the ideal functionality does not explicitly capture differential privacy.
-This omission is intentional for the sake of security analysis.
-In the security analysis, we decouple the privacy requirements into two sets of proofs: the first set demonstrates that our concrete protocol, when executed by real-world parties, realizes the ideal functionality.
-This means no additional information about parties’ inputs is revealed, except for the output given to each party.
-The second set of proofs establishes that the output given to each party conforms with differential privacy.
-V A Concrete Protocol
-V-A The Protocol
-We now present a concrete protocol for PIC.
+5 A Concrete Protocol
+5.1 The Protocol
+[…]
+We now present a concrete protocol for PIC in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model.
+The $\mathcal{F}$-hybrid model in MPC is a conceptual framework that simplifies the design and analysis of secure protocols.
+In many MPC protocols, parties rely on existing sub-protocols that have already been proven secure for specific tasks.
+To help protocol designers focus on other aspects of the protocol, these sub-protocols can be abstracted as ideal functionalities, denoted by $\mathcal{F}$, which are assumed to be computed securely by a trusted third party.
+This results in a hybrid protocol, combining concrete cryptographic operations from the real world with calls to ideal functionalities, thus avoiding the complexity of specifying and analyzing the sub-protocols in detail.
+In our case, we define the PIC protocol in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model, where $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ replaces the secure shuffle sub-protocol.We now present a concrete protocol for PIC.
 Because there are already secure protocols for shuffling, to simplify the description, we present the protocol in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model, in which parties can communicate as usual, and in addition have access to an ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ that does the shuffling.
@@
 [47]).
-Remark 3 To eventually accomplish the spatial crowdsourcing (e.g., taxi-hailing services) or location-based social system tasks, the computation result $y_{i,\pi_{i}(j)}$ will contain the matched users of the user $\pi_{i}(j)\in G_{i}$.
-That is, $y_{i,\pi_{i}(j)}$ will encapsulate a list of public keys and noisy location information about the matched users of the user $\pi_{i}(j)\in G_{i}$.
-After receiving the individualized computation results in the final step, each user $\pi_{i}(j)$ can then securely contact the matched users using the public keys in $y_{i,\pi_{i}(j)}$ (possibly with the help of a public communication channel, such as a public bulletin).
-V-B Security Analysis
+[…]
+5.2 Security Analysis
 This section presents the formal security analysis of the protocol in the previous section.
@@
 Formally we have the following theorem:
-Theorem V.1 (Security)
-The Private Individual Computation (PIC) protocol in §V securely computes $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model in the presence of any PPT adversary with static corruption.
-The proof can be found in Appendix -A.
+Theorem 5.1 (Security).
+The Private Individual Computation (PIC) protocol in §5 securely computes $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model in the presence of any PPT adversary with static corruption.
 The proof is simulation-based.
@@
 Security means that the simulated view is indistinguishable from the view of real-world execution.
+The detailed proof can be found in Appendix G of the full version [91].
 The above theorem states that the adversary learns strictly no more than the allowed output and leakage by engaging in the protocol execution.
@@
 Formally, we have the following theorem:
-Theorem V.2 (Differential Privacy)
+Theorem 5.2 (Differential Privacy).
 The Private Individual Computation protocol satisfies $(\epsilon_{c},\delta)$-DP, i.e.
 $\max(D_{\epsilon_{c}}(\mathcal{A}(X)||\mathcal{A}(X^{\prime})),D_{\epsilon_{c}}(\mathcal{A}(X^{\prime})||\mathcal{A}(X)))\leq\delta.$
-In particular, when all users in $G_{i^{*}}$ use an identical $\epsilon$-LDP mechanism as the data randomizer $\mathcal{R}_{i^{*}}$, and for $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|\geq 8(e^{\epsilon}+1)\log(2/\delta)$ we have:
-$\epsilon_{c}=\log\Bigg{(}1+\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\Big{(}\sqrt{\frac{32(e^{\epsilon}+1)\log 4/\delta}{n^{\prime}_{i^{*}}}}+\frac{4(e^{\epsilon}+1)}{n^{\prime}_{i^{*}}}\Big{)}\Bigg{)}.$ (1)
-The analysis can be divided into two cases: in the first case, $S\in\mathcal{C}$, i.e.
-S is corrupted.
-In this case, since the honest user locally randomizes their input, the input enjoys at least $\epsilon$-DP.
-Then the shuffling will amplify the privacy guarantee.
-Since the corrupted server receives the partial permutation as the leakage, the amplification depends on the number of uncorrupted users in the same group ($n^{\prime}_{i^{*}}$) as the honest user.
-The amplified $\epsilon_{c}$ can then be derived following [34, 35].
-In the second case where $S\not\in\mathcal{C}$, the knowledge of the adversary is $y_{i,j}$ for each corrupted user $u_{i,j}\in\mathcal{C}$.
-The tricky part is that how much information $y_{i,j}$ leaks depends on the function $f$ being evaluated by the server.
-Hence we consider the worst case where $f$ output $y_{i,j}=L$.
-Continuing along the same line of thought, we conclude that the level of differential privacy assurance is no less than in the first case.
-The full proof can be found in Appendix -B.
-Remark 4 After decryption, the server obtains $pk_{i,j}||x^{\prime}_{i,j}$, where $pk_{i,j}$ is a public key not sanitized by the local randomizer.
-Despite the presence of the public key, $pk_{i,j}||x^{\prime}_{i,j}$ and $x^{\prime}_{i,j}$ are equivalent in terms of privacy amplification.
-This is because (1) the public key is random and generated independently of $x_{i,j}$, so prepending it to $x^{\prime}_{i,j}$ does not affect the local DP guarantee—it is the same as $x^{\prime}_{i,j}$ itself; and (2) all public keys follow an identical probability distribution across all users in the group $G_{i}$, ensuring that the privacy amplification effect via shuffling is not degraded.
-V-C Discussion on Post-computation Communication
+In particular, when all users in $G_{i^{*}}$ use an identical $\epsilon$-LDP mechanism as the data randomizer $\mathcal{R}_{i^{*}}$, and for $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|\geq 16e^{\epsilon}\log(2/\delta)$ we have:
+$\epsilon_{c}=\log\Bigg{(}1+\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\Big{(}\sqrt{\frac{64e^{\epsilon}\log 4/\delta}{n^{\prime}_{i^{*}}}}+\frac{8e^{\epsilon}}{n^{\prime}_{i^{*}}}\Big{)}\Bigg{)}.$ (1)
+[…]
+[…]
+5.3 Discussion on Post-computation Communication
 In certain scenarios, such as spatial crowdsourcing and location-based social systems, there may be additional user-to-user communications following the execution of the PIC protocol.
@@
 Inevitably, a party has to sacrifice their privacy to the matched parties.
-The post-computation communication may also have privacy implications for other users not in the matched pair.
-Privacy amplification via shuffling against an adversary relies on the number of parties that remain anonymous.
-Recall that in Equation 1, the amplified $\epsilon_{c}$ relies on the number $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|$ of uncorrupted users in a particular group $G_{i^{*}}$.
-From the perspective of the adversary, if the post-computation communication compromises the anonymity of an additional set $U$ of the users in $G_{i^{*}}$, then $n^{\prime}_{i^{*}}$ becomes $|G_{i^{*}}-C_{i^{*}}-U|$.
-Accordingly, the privacy amplification effect for users in $G_{i^{*}}$ that are still anonymous is weakened.
+[…]
 We emphasize that preventing the loss or weakening of privacy via technical means is not feasible, because the information is necessary for the proper functioning of the application.
 However, managerial countermeasures, such as ensuring sufficiently large user groups and limiting post-computation exposure according to the need-to-know principle, can mitigate potential privacy risks arising from post-computation communication.
-VI Optimal Randomizers
-VI-A Inadequacy of Existing Randomizers
+6 Optimal Randomizers
+6.1 Inadequacy of Existing Randomizers
 In PIC, each user first sanitizes their data using an $\epsilon$-LDP randomizer.
@@
 $\max_{x_{j}\in\mathbb{X}}\mathbb{E}[\|\mathcal{R}(x_{j})-x_{j}\|_{2}^{2}],$
-where $\mathcal{R}$ is the LDP-randomizer employed by the users.
-Additionally, in the PIC model, user data is typically multi-dimensional (e.g., location, gradient), making sanitization significantly more difficult compared to scalar data.
-Another notable characteristic of the PIC model is that the local differential privacy budget $\epsilon$ is relatively large, often scaling linearly with $\log(n^{\prime}_{i^{*}})$, as implied by Theorem V.2.
-These factors together create issues when existing LDP randomizers are applied directly in the PIC model.
+[…]
 To understand the problem, we first examine a class of LDP randomizers [69, 90, 63, 53] that operate by sampling a few dimensions from $[d]$.
@@
 The Laplace/PlanarLaplace mechanisms introduce an MSE rate of $\frac{d}{\epsilon^{2}}$, while the PrivUnit/PrivUnitG mechanisms incur an MSE rate of $\frac{d}{\min\{\epsilon,\epsilon^{2}\}}$.
-In contrast, later we will show that the MSE rate can be improved to $(e^{\epsilon}-1)^{-2/(d+2)}$ (see Section VI-C and Appendix -D).
+In contrast, later we will show that the MSE rate can be improved to $(e^{\epsilon}-1)^{-2/(d+2)}$ (see Section 6.3 and Appendix A).
 Another strategy for submitting complete reports involves using additional randomization techniques such as random projection, data sketches, public randomness, or quantization.
@@
 The resulting additional error will never diminish even when $\epsilon\to+\infty$.
-VI-B Randomizer Design
+6.2 Randomizer Design
 We now introduce an asymptotically optimal randomizer, tailored for the PIC model.
 The randomizer uses an LDP mechanism, which we termed as Minkowski Response.
-For ease of presentation, here we will focus primarily on the $\ell_{2}$ case (and the $\ell_{+\infty}$ case in Appendix -D), but the mechanism can be generalized to other Minkowski distances.
+For ease of presentation, here we will focus primarily on the $\ell_{2}$ case (and the $\ell_{+\infty}$ case in Appendix A), but the mechanism can be generalized to other Minkowski distances.
 Without loss of generality, we assume the user data domain to be an $\ell_{2}$-bounded hyperball $\mathbb{X}=\{x\ |\ x\in\mathbb{R}^{d}\ and\ \|x\|_{2}\leq 1\}$.
@@
 Figure 4: The probability design of Minkowski response mechanism with a radius $r$. Illustrated are three inputs $x_{1},x_{2}$ and $x_{3}$, along with their respective cap areas.
-For any input $x\in\mathbb{X}$, Minkowski Response outputs an output $y\in\mathbb{Y}_{r}$ with relatively high probability in the cap area $\mathbb{B}_{r}(x)$ and relatively low probability in remaining domain $\mathbb{Y}_{r}\backslash\mathbb{B}_{r}(x)$ (see Figure 4).
-Formally:
+[…]
 $y=\left\{\begin{array}[]{@{}ll@{}}\text{uniform}(\mathbb{B}_{r}(x)),&\text{with prob.
@@
 When the budget $\epsilon$ gets large, the error rate of Minkowski Response decays faster than in previous LDP mechanisms.
-More specifically, the decay rate of Minkowski Response is $(e^{\epsilon}-1)^{-2/(d+2)}$ (Equation 6 in Appendix -D), while that of the previous mechanisms is $d/(e^{\epsilon/d}-1)^{2/3}$ or $d/\epsilon^{2}$.
+More specifically, the decay rate of Minkowski Response is $(e^{\epsilon}-1)^{-2/(d+2)}$ (Equation 4 in Appendix A), while that of the previous mechanisms is $d/(e^{\epsilon/d}-1)^{2/3}$ or $d/\epsilon^{2}$.
 Therefore, the utility advantage of Minkowski Response becomes more significant when $\epsilon$ gets larger.
@@
 no additional error).
-VI-C Analysis of Minkowski Response
-The local privacy guarantee of the randomizer is presented in Theorem VI.1.
-Theorem VI.1 (Local Privacy Guarantee)
+6.3 Analysis of Minkowski Response
+The local privacy guarantee of the randomizer is presented in Theorem 6.1.
+Theorem 6.1 (Local Privacy Guarantee).
 Given input domain $\mathbb{X}=\mathbb{B}_{1}$, the Minkowski response mechanism defined in Equation 2 satisfies $\epsilon$-LDP.
-Proof:
+Proof.
 It is observed that the output probability distribution in Equation 2 is valid for any input $x\in\mathbb{X}$, the probability density in the cap area $\mathbb{B}_{r}(x)$ is $\frac{e^{\epsilon}}{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}$, and the density in the non-cap area $\mathbb{Y}_{r}\backslash\mathbb{B}_{r}(x)$ is $\frac{1}{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}$.
@@
 ∎
-Next, we analyze the utility of the Minkowski Response in the PIC model.
-As previously mentioned, in the PIC model, the single report error is a more appropriate measure of utility compared to the statistical errors used in the conventional shuffle model.
-In Theorem VI.2, we examine the single report error in the PIC model and establish its lower bound.
-The proof is provided in Appendix -C.
-Theorem VI.2 (Error Lower Bounds)
-Given $d\in\mathbb{N}$, $\epsilon_{c}>0$, $\delta\in(0,0.5]$, $\mathbb{X}=\mathbb{B}_{1}(\{0\}^{d})$, then for any randomizer $\mathcal{R}:\mathbb{X}\mapsto\mathbb{R}^{d^{\prime}}$ such that $\mathcal{S}\circ\mathcal{R}(X)$ and $\mathcal{S}\circ\mathcal{R}(X^{\prime})$ are $(\epsilon_{c},\delta)$-indistinguishable for all possible neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, and for any estimator $f:\mathbb{R}^{d^{\prime}}\mapsto\mathbb{R}^{d}$, we have:
+[…]
+Theorem 6.2 (Error Lower Bounds).
+[…]
 $\max_{x\in\mathbb{X}}\mathbb{E}\big{[}\|f\circ\mathcal{R}(x)-x\|_{2}^{2}\big{]}\geq\tilde{\Omega}\big{(}{1}/{n^{\frac{2}{d+2}}}\big{)}.$
-The above theorem suggests that in the PIC model, for any randomizer, the single report error is at least $\tilde{\Omega}\big{(}{1}/{n^{\frac{2}{d+2}}}\big{)}$.
-Clearly, a randomizer offers better utility if its error is closer to this bound.
-For the randomizer using Minkowski Response, as presented in section VI-B, we can demonstrate that its single report error has an upper bound.
-This is summarized in Theorem VI.3, with the proof provided in Appendix -D.
-Theorem VI.3 (Error Upper Bounds)
+[…]
+Theorem 6.3 (Error Upper Bounds).
 For any $d\in\mathbb{N}$, $\epsilon_{c}>0$, $\delta\in(0,0.5]$, $\mathbb{X}=\mathbb{B}_{1}(\{0\}^{d})$, if $\epsilon_{c}\leq O(1)$ and $n>\max\{16\log(1/\delta),\frac{2^{d+7}\log(1/\delta)}{(e^{\epsilon_{c}}-1)^{2}}\}$, then there exist a randomizer $\mathcal{R}:\mathbb{X}\mapsto\mathbb{R}^{d}$ such that $\mathcal{S}\circ\mathcal{R}(X)$ and $\mathcal{S}\circ\mathcal{R}(X^{\prime})$ are $(\epsilon_{c},\delta)$-indistinguishable for all possible neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, and:
@@
 We observe that in most applications, $(\epsilon_{c},\delta)$ are given as fixed system parameters.
-If we consider $(\epsilon_{c},\delta)$ as constants, then the error upper bound of the Minkowski Response randomizer in Theorem VI.3 is $\tilde{O}\left(\frac{1}{n^{2/(d+2)}}\right)$, which matches the error lower bound of the PIC model in Theorem VI.2.
+If we consider $(\epsilon_{c},\delta)$ as constants, then the error upper bound of the Minkowski Response randomizer in Theorem 6.3 is $\tilde{O}\left(\frac{1}{n^{2/(d+2)}}\right)$, which matches the error lower bound of the PIC model in Theorem 6.2.
 This implies that the utility of the Minkowski Response is asymptotically optimal.
@@
 If we compare Minkowski Response in the PIC model to using LDP directly (without shuffling), the utility advantage is even greater: any randomizers in the LDP model must endure a single report error of $\Omega\left(\frac{d}{\epsilon_{c}^{2}}\right)$ when $\epsilon_{c}\leq O(1)$ [12, 28].
-VII Experimental Evaluation
+7 Experimental Evaluation
 We evaluate the efficacy of our PIC protocol and Minkowski randomizer.
 We compare the utility of our proposal against state-of-the-art works in the context of three representative individual computation tasks: spatial crowdsourcing, location-based social systems, and federated learning with incentives.
-VII-A Spatial Crowdsourcing
+7.1 Spatial Crowdsourcing
 Datasets We use two real-world datasets: GMission dataset [18] for scientific simulation, and EverySender dataset [86] for campus-based micro-task completion.
-Details about the two datasets are summarized in Table II, including the number of users/workers, location domain range, and serving radius of workers about these two datasets (more information can be found in Appendix -F).
+Details about the two datasets are summarized in Table 2, including the number of users/workers, location domain range, and serving radius of workers about these two datasets (more information can be found in Appendix C of the full version [91]).
 We normalize location data to the domain $[-1,1]\times[-1,1]$ and scale the serving radius to $0.41.0\cdot\frac{1-(-1)}{5-0}=$ 0.4 correspondingly.
-TABLE II: Summary statistics of spatial crowdsourcing datasets.
+Table 2: Summary statistics of spatial crowdsourcing datasets.
 Dataset users workers location domain serving radius
@@
 EverySender 4036 817 $[0,5.0]\times[0,5.0]$ $1.0$
-LDP randomizers We use the Minkowski response for location randomization.
-As a comparison, in the local model of DP (e.g., in [85, 80, 84]), we compare with existing mechanisms including Laplace [31], Staircase [38], PlanarLaplace with geo-indistinguishability [2], SquareWave [58], PrivUnit [12] and its Gaussian variant PrivUnitG in [6].
+[…]
 Server-side algorithms Two commonly used server-side algorithms are evaluated as concrete tasks: minimum weighted full matching [54] and maximum matching [48].
@@
 Figure 6: Travel costs of minimum weighted matching on GMission and EverySender dataset.
-Finally, we present the utility comparison using the maximum matching algorithm, as shown in Fig. 7. In this case, task-specific utility is assessed by the successful matching ratio:
+Finally, we present the utility comparison using the maximum matching algorithm, as shown in Fig. 7. In this case, utility is assessed by the successful matching ratio:
 $\frac{\sum_{(i,j)\in G_{a}\times G_{b}}\llbracket M(i,j)>0\rrbracket\cdot\llbracket\|l_{i}-l_{j}\|_{2}\leq\tau\rrbracket}{\min\{|G_{a}|,|G_{b}|\}}.$
@@
 Figure 7: Success ratios of maximum matching in spatial crowdsourcing on GMission and EverySender dataset.
-VII-B Location-based Social Systems
-TABLE III: Details of location social network datasets.
+7.2 Location-based Social Systems
+Table 3: Details of location social network datasets.
 Dataset check-ins location domain
@@
 Gowalla and Foursquare are location-based social network websites where users share their locations by checking-in.
-Details about the two datasets are summarized in Table III.
+Details about the two datasets are summarized in Table 3.
 As before, the location data is normalized to $[-1,1]\times[-1,1]$.
-LDP randomizers The LDP randomizers used are the same as those in Section VII-A.
+LDP randomizers The LDP randomizers used are the same as those in Section 7.1.
 Server-side algorithm The server performs the radius-based nearest neighbor (NN) search for the users, which is a common task in location-based social networks [21].
@@
 Note that in this application scenario, the returned neighbors may be deanonymized in the post-computation phase.
-Hence, the actual privacy guarantee in the PIC model depends on the number of users who remain anonymous (see discussion in Section V-C).
-Considering this, and each user normally has no more than $n\cdot\frac{\pi\tau^{2}}{2^{2}}<n\cdot 3.2\%$ neighbors, we use privacy amplification population $\lfloor n\cdot 90\%\rfloor$ instead of the group size $n$ when calculating the local budget given the global $\epsilon_{c}$.
+Hence, the actual privacy guarantee in the PIC model depends on the number of users who remain anonymous (see discussion in Section 5.3).
+Considering this, and each user normally has about $n\cdot\frac{\pi\cdot\tau^{2}}{2^{2}}\approx n\cdot 3.2\%$ neighbors, we use privacy amplification population $\lfloor n\cdot 90\%\rfloor$ instead of the group size $n$ when calculating the local budget given the global $\epsilon_{c}$.
 Figure 8: Expected $\ell_{2}$ distances of reported locations in location-based social systems with the local model of DP.
@@
 Let $N_{i}$ denote the set of true neighbors of user $i$ within an $\ell_{2}$-distance $\tau$, and let $\widehat{N}_{i}$ denote the retrieved neighbor set computed using the sanitized reports.
-The precision of nearest neighbor queries is defined as $\frac{\sum_{i\in[n]}|N_{i}\cap\widehat{N}_{i}|}{\sum_{i\in[n]}|\widehat{N}_{i}|}$, and the recall is defined as $\frac{\sum_{i\in[n]}|N_{i}\cap\widehat{N}_{i}|}{\sum_{i\in[n]}|N_{i}|}$.
-The $F_{1}$ score is then calculated as:
+The precision of nearest neighbor queries is defined as ${\sum_{i\in[n]}|N_{i}\cap\widehat{N}_{i}|}/{\sum_{i\in[n]}|\widehat{N}_{i}|}$, and the recall is defined as ${\sum_{i\in[n]}|N_{i}\cap\widehat{N}_{i}|}/{\sum_{i\in[n]}|N_{i}|}$.
+The $F_{1}$ score is as:
 $F_{1}\text{ score}=2\cdot\frac{\text{precision}\cdot\text{recall}}{\text{precision}+\text{recall}}.$
@@
 With a large group, the Minkowski randomizer can achieve an $F_{1}$ score of 0.9 with a stringent global budget $\epsilon_{c}=1$ and nearly 1 if the budget is loosened to $\epsilon_{c}=3$.
-VII-C Federated Learning with Incentives
+7.3 Federated Learning with Incentives
 In this application, the users collaboratively train a model by Federated learning, and the server decides each user’s incentive by how much their local gradient contributes to the global model.
@@
 The MNIST dataset comprises 60,000 images, with 50,000 designated as training samples.
-Each user receives one training sample and trains a neural network model, as detailed in Table IV, which contains $d=4292$ trainable parameters.
+Each user receives one training sample and trains a neural network model, as detailed in Table 4, which contains $d=4292$ trainable parameters.
 In each round, $s=10,000$ users are randomly selected.
@@
 Each user submits a sanitized version of the subsampled, clipped gradient vector as their report.
-TABLE IV: Model architecture of the neural network.
+Table 4: Model architecture of the neural network.
 Layer Parameters
@@
 This global gradient vector is published so that users can update their local models.
-Additionally, the server computes the Shapley value for each user, which measures the marginal contribution of each $\hat{g}_{i}$ (see Appendix -H for details).
+Additionally, the server computes the Shapley value for each user, which measures the marginal contribution of each $\hat{g}_{i}$ (see Appendix E for details in the full version [91]).
 Experimental results We train each model for 80 rounds using Federated Learning.
-The utility comparison results are summarized in Tables V and VI, each reflecting a different global privacy budget.
+The utility comparison results are summarized in Tables 5 and 6, each reflecting a different global privacy budget.
 In these tables, we present the single report utility (Gradient $\ell_{2}$ error) and the task-specific utility (Shapley $\ell_{2}$ error).
@@
 For reference, we also include the final models’ accuracy in the tables.
-TABLE V: Utility comparison of federated learning with incentives: global privacy budget $(1,0.01/50000)$.
+Table 5: Utility comparison of federated learning with incentives: global privacy budget $(1,0.01/50000)$.
 Setting
@@
 Minkowski $\mathbf{78.68\%}$ $\mathbf{0.093}$ $\mathbf{0.0363}$
-TABLE VI: Utility comparison of federated learning with incentives: global privacy budget $(3,0.01/50000)$.
+Table 6: Utility comparison of federated learning with incentives: global privacy budget $(3,0.01/50000)$.
 Setting
@@
 For Minkowski, the issue lies in the 80-round federated learning process, where the privacy budget for each round is relatively small (e.g., between 0.2 and 0.75).
-As previously mentioned, the Minkowski mechanism is intended to achieve better utility with a large local privacy budget.
-In small privacy budget scenarios, it offers no advantage and may even perform worse.
+As previously mentioned, the Minkowski is intended to achieve better utility with a large local privacy budget.
+In small budget scenarios, it offers no advantage and may even perform worse.
 In the PIC model, the performance of Staircase and Squarewave remains poor.
 However, Minkowski now performs the best among all randomizers.
-This improvement is due to the privacy amplification effect in PIC, which increases the local privacy budget for each round to between 4.0 and 6.0.
-VIII Conclusion
-Privacy-preserving computation with Differential Privacy (DP) holds great potential for leveraging personal information.
+This improvement is due to the privacy amplification in PIC, which increases the local budget for each round to between 4.0 and 6.0.
+8 Discussions
+[…]
+Decentralized download-then-compute settings.
+While our protocol initially allows a server to perform the computation $f$, it is also possible for users to first download $L$ and then compute $f$ locally.
+This download-then-compute paradigm can be seen as a special decentralized case of our PIC model (i.e.
+the server simply computes $f(x)=x$).
+It would also rely on the one-time random key to perform post-computation communication.
+The PIC protocol offers several advantages over the download-then-compute paradigm: (1) PIC accommodates more tasks, e.g.
+those dependent on private server information (validation dataset in federated learning with incentives).
+(2) Download-then-compute requires public $f$, while PIC allows proprietary server-side algorithms.
+(3) Better privacy.
+E.g.
+in taxi-hailing, download-then-compute allows each user to compute all matching pairs and location distribution, while PIC publishes encrypted matching results (without leaking unnecessary information).
+[…]
+Necessity of grouping users.
+While many applications involve a single type of participant (e.g., social systems and federated learning), other tasks require interaction between multiple participant types, where the server must use the type information of participants to execute the algorithm $f$ (e.g., in taxi-hailing services with types of drivers and users).
+From a privacy-preservation standpoint, it is also possible to apply varying levels of privacy protection across different groups.
+[…]
+Worst-case leakage and privacy under active attacks.
+We assume that an attacker can corrupt and control a subset of parties, but the data of the uncorrupted users remains $(\epsilon_{c},\delta)$-private against the attacker.
+The specific value of $\epsilon_{c}$ depends on the number of uncorrupted users within the same group (see Theorem 5.2, Eq.
+1).
+In the worst-case scenario, where all but one user are corrupted, the privacy guarantee is reduced to $\epsilon$-LDP, as provided by the local randomizer, i.e., there is no privacy amplification through shuffling.
+Multiple executions and reuse keys.
+When multiple/sequential PIC executions are needed (e.g., in federated learning), a fresh one-time user key for each execution is required, ensuring the adversary cannot correlate user activities.
+According to Eq.
+1 and sequential composition of DP [31], the overall privacy consumption for $k$ executions is $\tilde{O}(\sqrt{k\cdot e^{\epsilon}/{n_{i}^{*}}})$.
+Otherwise, reusing keys in $k$ executions will consume more privacy (to a level of shuffling $n_{i}^{*}$ messages each is $k\cdot\epsilon$-LDP, i.e., $\tilde{O}(\sqrt{e^{k\cdot\epsilon}/{n_{i}^{*}}})$).
+[…]
+Multi-message settings.
+The PIC model allows each client to send one single message, to align with most non-private applications.
+However, restricting each client to sending one message in the shuffle model has intrinsic privacy amplification and utility limitations (as seen in statistical analyses within the shuffle model [40]).
+Extending PIC model to multi-message settings is an interesting future direction, though it is quite challenging (e.g., in taxi-hailing services, the messages from a driver must appear independent, then each message may match a different customer, leading to an oversale).
+9 Conclusion
+Privacy-preserving computation with differential privacy holds great potential for leveraging personal information.
 While the shuffle model offers a rigorous DP guarantee with enhanced utility, its application is confined to statistical tasks.
@@
 Extensive experiments validate the superiority of the PIC protocol and the randomizer, showcasing their performance across three major application scenarios and various real-world datasets.
+[…]
+Future directions.
+The relaxed notion of local metric DP, or geo-indistinguishability [2], also shows shuffle privacy amplification effects [92], making it a promising area for future research to further enhance utility in the PIC model.
+Acknowledgments
+We thank the anonymous reviewers for their insightful suggestions and comments.
+This work is supported by National Key Research and Development (R&D) Program (Young Scientist Scheme No.2022YFB3102400), National Natural Science Foundation of China (No.62372120, No.62302118, No.62372125, No.62261160651, No.62102108), the Guangdong Natural Science Funds for Distinguished Young Scholar under Grant 2023B1515020041, Natural Science Foundation of Guangdong Province of China (No.2022A1515010061), and Guangzhou Basic and Applied Basic Research Foundation (No.2025A03J3182).
+Di Wang is supported in part by the funding BAS/1/1689-01-01, URF/1/4663-01-01, REI/1/5232-01-01, REI/1/5332-01-01, and URF/1/5508-01-01 from KAUST, and funding from KAUST - Center of Excellence for Generative AI, under award number 5940.
+Ethics considerations
+This work does not present any ethical issues.
+In fact, it directly mitigate privacy concerns associated with computational tasks involving sensitive user data, such as location data, preference data, and gradient information.
+All experiments are conducted using public datasets, ensuring no exposure of personal information.
+Moreover, we strictly adhere to ethical guidelines throughout the research process, ensuring that no other ethical concerns arise.
+Open science
+In alignment with the principles of open science, our code and experimental data (for the three exemplar applications of the PIC model in Section 7) are publicly accessible at https://zenodo.org/records/14710367 (please refer to README.md for more details).
+Appendix A Error Upper Bounds of Minkowski Response Mechanism
+We now study the error bound of the Minkowski response in the PIC model.
+We start with analyze the mean squared error formula of the mechanism given fixed local budget $\epsilon$ and cap area radius parameter $r$.
+Then, we apply the global privacy budget $(\epsilon_{c},\delta)$ and the privacy amplification bound in Theorem 5.2, to deduce a feasible local budget $\epsilon$ and optimized radius parameter afterward.
+To deal with both $\ell_{2}$-norm and $\ell_{+\infty}$-norm bounded domain, we introduce a more general notation $\mathbb{B}_{p,r}(x)$ to represent the $\ell_{2}$-norm hyperball with radius $r$ centered at any $x\in\mathbb{R}^{d}$:
+$\mathbb{B}_{p,r}(x)=\{x^{\prime}\ |\ x^{\prime}\in\mathbb{R}^{d}\ and\ \|x^{\prime}-x\|_{p}\leq r\},$
+and a general notation $\mathbb{Y}_{p,q,r}$ to present the following $\ell_{q}$ expanded domain:
+$\mathbb{Y}_{p,q,r}=\{x\ |\ x\in\mathbb{R}^{d}\ and\ \exists x^{\prime}\in\mathbb{B}_{p,1}\ that\ x\in\mathbb{B}_{q,r}(x^{\prime})\}.$
+For hyper-ball domain $\mathbb{B}_{2,1}$.
+When the domain is $\ell_{2}$-norm bounded hyperball $\mathbb{B}_{2,1}$, we use $q=2$ for the cap area as well.
+In this context, we let $\beta=\frac{r^{d}(e^{\epsilon}-1)}{(1+r)^{d}+r^{d}(e^{\epsilon}-1)}$ and obtain the MSE bound given fixed local budget $\epsilon$ and radius $r$ as follows:
+$\max_{x\in\mathbb{B}_{p,1}}\mathbb{E}[\|\tilde{x}-x\|_{2}^{2}]=\max_{x\in\mathbb{B}_{p,1}}\frac{1}{\beta^{2}}\cdot\textsf{Var}[y|x]$
+$=$ $\max_{x\in\mathbb{B}_{p,1}}\frac{1}{\beta^{2}}(\beta\cdot\mathbb{E}[\|\mathbb{B}_{q,r}(x)\|_{2}^{2}]+(1-\beta)\cdot\mathbb{E}[\|\mathbb{Y}_{p,q,r}(x)\|_{2}^{2}]-\beta^{2}\|x\|_{2}^{2})$
+$=$ $\max_{x\in\mathbb{B}_{p,1}}\frac{1}{\beta^{2}}(\beta(\|x\|_{2}^{2}+r^{2})+(1-\beta)(1+r)^{2}-\beta^{2}\|x\|_{2}^{2})$
+$\leq$ $\frac{1}{\beta^{2}}(\beta(1+r^{2})+(1-\beta)(1+r)^{2}-\beta^{2})$
+$\leq$ $\frac{1}{\beta^{2}}(\beta r^{2}+(1-\beta)(1+(1+r)^{2}))$
+where $\mathbb{E}[\|\mathbb{B}\|_{2}^{2}]$ denote the expected squared distance between a (uniform-distributed) space $\mathbb{B}\subseteq\mathbb{R}^{d}$ and the origin point $\{0\}^{d}$.
+If the local privacy budget $\epsilon$ is relatively large (e.g., $\epsilon\geq\log((c+1)^{\frac{d+2}{2}})$ for some constant $c\geq 1$), and we specify $r=(e^{\epsilon}-1)^{2/(d+2)}-1)^{-1}$, we then have $r\leq 1/c$, $\beta\in[1/2,1]$ and:
+$\max_{x\in\mathbb{B}_{p,1}}\mathbb{E}[\|\tilde{x}-x\|_{2}^{2}]$
+$\leq$ $4(r^{2}+5\frac{(1+1/r)^{d}}{(e^{\epsilon}-1)+(1+1/r)^{d}})$
+$\leq$ $4(r^{2}+5\frac{(1+1/r)^{d}}{e^{\epsilon}-1})$
+$\leq$ $4(e^{\epsilon}-1)^{-2/(d+2)}+5\frac{(e^{\epsilon}-1)^{-2/(d+2)}}{e^{\epsilon}-1})$
+$\leq$ $24(e^{\epsilon}-1)^{-2/(d+2)}.$ (4)
+Consider the case $n\geq\max\big{\{}16\log(1/\delta),\frac{256(1+c)^{d+2}\log(1/\delta)}{(e^{\epsilon_{c}}-1)^{2}}\big{\}}$ holds for some constant $c\geq 1$, we specify local budget $\epsilon$ such that $e^{\epsilon}=\frac{n(e^{\epsilon_{c}}-1)^{2}}{256\log(1/\delta)}$ holds according to Theorem 5.2, and specify the radius $r$ to:
+$\Big{(}\big{(}\frac{n(e^{\epsilon_{c}}-1)^{2}}{256\log(1/\delta)}\big{)}^{1/(d+2)}-1\Big{)}^{-1}.$
+Observe that in this setting, we have $\beta\in[1/2,1]$ and $r\leq 1/c$. Then, the MSE is upper bounded as:
+$\max_{x\in\mathbb{B}_{p,1}}\mathbb{E}[\|\tilde{x}-x\|_{2}^{2}]$
+$\leq$ $\frac{1}{\beta^{2}}(\beta r^{2}+(1-\beta)(1+(1+r)^{2}))$
+$\leq$ $4(r^{2}+5\frac{(1+1/r)^{d}}{(e^{\epsilon}-1)+(1+1/r)^{d}})$
+$\leq$ $4(r^{2}+5\frac{(1+1/r)^{d}}{e^{\epsilon}})$
+$\leq$ $4\Big{(}\big{(}\frac{256\log(1/\delta)}{n(e^{\epsilon_{c}}-1)^{2}}\big{)}^{\frac{2}{d+2}}\cdot\frac{(c+1)^{2}}{c^{2}}+\frac{5((e^{\epsilon_{c}}-1)^{2}n/(256\log(1/\delta)))^{\frac{d}{d+2}}}{(e^{\epsilon_{c}}-1)^{2}n/(256\log(1/\delta))}\Big{)}$
+$\leq$ $36\Big{(}\frac{256\log(1/\delta)}{n(e^{\epsilon_{c}}-1)^{2}}\Big{)}^{\frac{2}{d+2}}.$
+Therefore, we establish Theorem 6.3. With sufficiently large size $n$ of the amplification population, the derived error bound matches the lower bound in Theorem 6.2.
+For hyper-cube domain $\mathbb{B}_{\infty,1}$.
+Another data domain that is commonly encountered in practical settings is the $\ell_{+\infty}$-norm bounded hypercube.
+We use $q=+\infty$ as well for the cap area, then we have volumes $V(\mathbb{B}_{2,r})=(2r)^{d}$, $V(\mathbb{Y}_{\infty,\infty,r})=(2+2r)^{d}$, and let $\beta=\frac{r^{d}(e^{\epsilon}-1)}{(1+r)^{d}+r^{d}(e^{\epsilon}-1)}$.
+For fixed local budget $\epsilon$ and radius $r$, the mean squared error bound is:
+$\max_{x\in\mathbb{B}_{p,1}}\mathbb{E}[\|\tilde{x}-x\|_{2}^{2}]=\frac{1}{\beta^{2}}\cdot\text{Var}[y]$
+$\leq$ $\frac{d}{\beta^{2}}(\beta(1+r^{2}/3)+(1-\beta)(1+r)^{2}/3)-d$
+$\leq$ $\frac{d}{3\beta^{2}}(\beta r^{2}+(1-\beta)((1+r)^{2}+3(1-\beta)))$
+Consider the case $n\geq\max\{16\log(1/\delta),\frac{32(1+c)^{d+2}\log(1/\delta)}{(e^{\epsilon_{c}}-1)^{2}}\}$ holds for some constant $c\geq 1$, we specify local budget $\epsilon$ such that $e^{\epsilon}=\frac{n(e^{\epsilon_{c}}-1)^{2}}{256\log(1/\delta)}$ holds according to Equation 1, and specify the radius $r$ as:
+$\Big{(}\big{(}\frac{n(e^{\epsilon_{c}}-1)^{2}}{256\log(1/\delta)}\big{)}^{1/(d+2)}-1\Big{)}^{-1}.$
+In this setting, we have $\beta\in[1/2,1]$ and $r\leq 1/c$, we thus obtain:
+$\max_{x\in\mathbb{B}_{p,1}}\mathbb{E}[\|\tilde{x}-x\|_{2}^{2}]$
+$\leq$ $\frac{d}{3\beta^{2}}(r^{2}+7(1-\beta))$
+$\leq$ $\frac{d}{3\beta^{2}}(r^{2}+7\frac{(1+1/r)^{d}}{(e^{\epsilon}-1)+(1+1/r)^{d}})$
+$\leq$ $\frac{d}{3\beta^{2}}(r^{2}+7\frac{(1+1/r)^{d}}{e^{\epsilon}})$
+$\leq$ $\frac{4d}{3}\Big{(}\big{(}\frac{256\log(1/\delta)}{n(e^{\epsilon_{c}}-1)^{2}}\big{)}^{\frac{2}{d+2}}\cdot\frac{(c+1)^{2}}{c^{2}}+7\big{(}\frac{256\log(1/\delta)}{n(e^{\epsilon_{c}}-1)^{2}}\big{)}^{\frac{2}{d+2}}\Big{)}$
+$\leq$ ${15d}\cdot\Big{(}\frac{256\log(1/\delta)}{n(e^{\epsilon_{c}}-1)^{2}}\Big{)}^{\frac{2}{d+2}}.$
+Alternatively, one may firstly transform the $\ell_{+\infty}$-norm vector into a $\ell_{2}$-norm bounded one, and utilizing the mechanism for hyper-ball.
+Similar utility can be guaranteed for both ways.
+Appendix B Details on Experimental Implementation
+In the experiments related to both spatial crowdsourcing and location-based social systems, user location data is confined within a two-dimensional cube domain $[-1,1]\times[-1,1]$. As a result:
+  • For the Laplace mechanism, the privacy sensitivity parameter related to replacement is defined as $\Delta=4$.
+  • In the PlanarLaplace mechanism, given that the maximum $\ell_{2}$-distance is $2\sqrt{2}$, we set the geo-indistinguishability parameter to $\epsilon/(2\sqrt{2})$ to ensure a fair comparison.
+  • In mechanisms like Staircase and Squarewave, which originally operate in one-dimensional domain, the local budget is evenly distributed across two dimensions.
+This is crucial for generating meaningful location reports pertinent to these tasks.
+  • For the PrivUnit mechanism, which uses an $\ell_{2}$-bounded unit vector as input, we convert the two-dimensional cube domain into a three-dimensional hyper-ball domain.
+After randomization, it’s reverted back to its original two-dimensional form.
+To enhance performance, we further engage in a numerical search for the optimal hyper-parameter, following the approach in [36].
+  • In the Minkowski response mechanism, we set $q=+\infty$ for the cap area to align with the input domain, and engage in a numerical search for the best-suited cap area radius $r$.
+  • We additionally introduce a classical mechanism by Duchi et al. [29], denoted as PrivHS, for comparison.
+In the experiments of federated learning with incentives, the (randomly) subsampled gradient vector has $6$ dimensions, thus the two-dimensional PlanarLaplace mechanism is unapplicable.
+For mechanisms like Staircase and Squarewave, the local budget is evenly distributed across $6$ dimensions.
+When privacy amplification via shuffling is applied in the PIC model, the parameter $\delta$ is fixed to $0.01/n_{i}$, where $n_{i}$ is the number of users in the same group $i$.
+In the spatial crowdsourcing applications, since one user is associated with at most one worker, the effective number $n^{\prime}_{i}$ of amplification population is set to $n_{i}-1$; in the location-based social system applications, the effective number $n^{\prime}$ of amplification population is set to $0.90\cdot n$ when neighboring radius $\tau=0.2$ and is set to $0.98\cdot n$ when neighboring radius $\tau=0.1$.
+We evaluate both the client-side and server-side running time of our protocol on a laptop computer embedded with Intel i5555-8250U CPU @1.6GHz and $8$GB memory.
+Appendix C Spatial Crowdsourcing
+Table 7: Mean $\ell_{2}$-error (and running time) comparison of local $\epsilon$-LDP randomizers on location domain $[-1,1]\times[-1,1]$.
+All results are the expected value of $1000$ repeated experiments.
+randomizer $\epsilon=0.5$ $\epsilon=1.0$ $\epsilon=2.0$ $\epsilon=3.0$ $\epsilon=5.0$ $\epsilon=8.0$ $\epsilon=10.0$
+PrivUnit [12] 8.94 (49us) 4.68 (42us) 2.25 (61us) 1.44 (63us) 0.81 (113us) 0.32 (305us) 0.18 (868us)
+PrivUnitG[6] 8.73 (76ms) 4.63 (75ms) 2.27 (77ms) 1.51 (76ms) 0.96 (74ms) 0.63 (75ms) 0.53 (81ms)
+Laplace [31] 12.97 (0.1us) 6.56 (0.1us) 3.27 (0.1us) 2.13 (0.1us) 1.30 (0.1us) 0.81 (0.1us) 0.64 (0.1us)
+PlanarLaplace [2] 11.17 (12us) 5.63 (11us) 2.84 (11us) 1.88 (11us) 1.14 (11us) 0.71 (12us) 0.56 (12us)
+Staircase[39] 13.19 (49us) 6.40 (48us) 3.13 (46us) 2.01 (48us) 1.05 (44us) 0.46 (51us) 0.28 (50us)
+SquareWave[62] 11.87 (1.9us) 5.72 (2.5us) 2.65 (1.9us) 1.68 (2.2us) 0.92 (1.7us) 0.53 (2.2us) 0.42 (2.1us)
+MinkowskiResponse 10.42 (0.7us) 4.50 (0.6us) 1.78 (0.8us) 0.98 (0.8us) 0.39 (0.6us) 0.14 (0.7us) 0.074 (0.8us)
+Table 8: Running time and communication overheads of spatial crowdsourcing in the PIC model.
+User-side Procedures Time Communication Server-side Procedures Time Communication
+location randomization 0.7us - decrypt one message 2.1ms -
+encrypt information 2.9ms - min-weight match (GMission) 31ms 1.2MB
+send message to shuffler - 1.3KB maximum match (GMission) 15ms 1.2MB
+retrieve matching result - 1.8KB min-weight match (EverySender) 79ms 4.3MB
+key agreement with worker 2.6ms - maximum match (EverySender) 63ms 4.3MB
+Dataset descriptions.
+The GMission dataset originates from a spatial crowdsourcing platform for scientific simulations.
+It contains information about every task, including its description, location, time of assignment, and deadline (in minutes).
+Furthermore, it provides data about each worker, comprising their location, maximum activity range (in kilometers), etc.
+EverySender, on the other hand, represents a campus-based spatial crowdsourcing platform, facilitating everyone to post micro tasks like package collection or to act as a worker.
+Similar to GMission, EverySender dataset also carries detailed information for every task and worker.
+We assume each worker’s capacity as one and, for simplicity, we consider only the location information of the taskers/workers for matching purposes.
+Before exploring spatial crowdsourcing tasks, we first evaluate the utility and efficiency of our proposed Minkowski response mechanism against existing mechanisms in the LDP model, as presented in Table 7.
+This evaluation focuses on errors and running times associated with reporting locations under local $\epsilon$-DP constraints.
+We quantify the error using the expected $\ell_{2}$ distance between each true and noisy location pair.
+Notably, the Minkowski response showcases an $\ell_{2}$ error comparable to state-of-the-art (SOTA) mechanisms when $\epsilon\leq 1.0$.
+Its superiority is evident when $\epsilon\geq 2.0$, displaying a significantly reduced error.
+This supports the theoretical assertions made in Section 6, underscoring the importance of the Minkowski response for the shuffle model, over other existing randomizers.
+Additionally, the Minkowski response is highly efficient, demanding less than $1$us on the user end.
+This speed is nearly equivalent to merely adding Laplace noises and is roughly $50$x faster than the previously SOTA PrivUnit.
+We present the computation and communication overheads of the PIC model for spatial crowdsourcing in Table 8.
+The total running time for each user is only a few milliseconds (with the data randomization time being negligible), and the total communication overhead is under $4$KB.
+On the server side, the running time of the matching algorithm on noisy plaintext is consistent with non-private settings, scaling with the number of users/workers and taking dozens of milliseconds.
+The additional decryption/encryption runtime is capped at several seconds, mirroring the costs of HTTPS decryption/encryption.
+Specifically, it utilizes ECDHE for key agreement and AES encryption as with HTTPS on TLS 1.3 [73] for client-server communication.
+In essence, the running times on both the client and server sides of the PIC model closely resemble those in non-private settings that use HTTPS communication (for a fair comparison).
+Appendix D Location-based Social Systems
+Dataset descriptions.
+Gowalla is a location-based social networking website where users share their locations by checking-in.
+The Gowalla consists a total of 6,442,890 check-ins of $138368$ users over the period of Feb.
+2009 - Oct.
+2010, in the San Francisco, CA.
+The Foursquare dataset contains check-ins in New York city collected for about 10 month (from 12 April 2012 to 16 February 2013).
+It contains 227,428 check-ins in New York city.
+Each check-in is associated with its time stamp, its GPS coordinates and its semantic meaning (represented by fine-grained venue-categories).
+To further illustrate the effects of neighboring radius $\tau$ on performances, we consider $(\tau=0.1)$-radius nearest neighbor queries, where each user has several tens or hundreds of neighbors.
+The experimental results on the LDP model are presented in Figure 12, and the results on the PIC shuffle model are presented in Figures 13 and 14.
+Considering post-computation user-to-user communications, and each user has about $n\cdot\frac{\pi\cdot 0.1^{2}}{2^{2}}\approx n\cdot 0.8\%$ neighbors, we use $\lfloor n\cdot 98\%\rfloor$ when deriving the local budget given the global budget $\epsilon_{c}$.
+The combination of the PIC model and the Minkowski response mechanism outperforms competitors in most scenarios.
+The Minkowski randomizer can achieve an $F_{1}$ score of 0.75 with a stringent global budget $\epsilon_{c}=1$ and over 0.80 with $\epsilon_{c}=3$.
+Compared to the $\tau=0.2$ scenarios, the $\tau=0.1$ cases retrieve nearer (and fewer) neighbors, but can be less stable under DP noises (i.e., has lower F1 scores).
+Figure 12: F1 scores of nearest neighbor queries (LDP model) with radius $\tau=0.1$.
+Figure 13: Expected $\ell_{2}$ distances of reported locations in location-based social systems with PIC model with radius $\tau=0.1$.
+Figure 14: F1 scores of nearest neighbor queries (PIC model) with radius $\tau=0.1$.
+Appendix E Federated Learning with Incentives
+Among 60000 images in the MNIST dataset, 50000 of them are designated as training samples, 5000 of them works as the validation dataset and the remaining 5000 images works as the test dataset.
+During the 80 rounds of (noisy) gradient aggregation and model updating, we use Adam optimizer with learning rate 0.05 and decay rate 0.99.
+Table 9: Comparison of various approaches to multi-party PIC computation, with $C$ gates and $h$ depth of computation circuits, and $(\epsilon,\delta)$-DP where $\epsilon=O(1)$, where $\lambda$ is the security parameter, the $n$ is number of clients/inputs, $d$ is the dimension of inputs.
+The Privatization Error is the expected mean squared error between an input and its reported values due to differential privacy; the Algo.
+on Plaintext indicates whether the computation circuit is running on un-encrypted plaintext.
+Paradigm
+Client
+Comm. Rounds
+Client
+Comm. Bits
+Orchestrator
+Comput. Costs
+Algo. on
+Plaintext?
+Privatization
+Error
+plaintext $1$ $O(d)$ $O(nd)$
+plaintext with secure communication $1$ $O(\max\{d,\lambda\})$ $O(n\cdot\max\{d,\lambda\})$
+GMW/BGW protocols [44, 11] $h$ $O(Poly(n)\cdot C\cdot\lambda)$ $O(Poly(n)\cdot C\cdot\lambda)$
+BMR garbled circuits [96, 9] $3$ $O(Poly(n)\cdot C\cdot\lambda)$ $O(Poly(n)\cdot C\cdot\lambda)$
+Ishai’s anonymous model [51]+ABT[3] $3$ $O(Poly(n)\cdot C\cdot\lambda)$ $O(Poly(n)\cdot C\cdot\lambda)$
+Beimel’s anonymous model [10] $2$ $O(Poly(n)\cdot C\cdot\lambda)$ $O(Poly(n)\cdot C\cdot\lambda)$
+local DP model [57] with secure communication $1$ $O(\max\{d,\lambda\})$ $O(n\cdot\max\{d,\lambda\})$ $O(d/\epsilon^{2})$
+PIC model $1$ $O(\max\{d,\lambda\})$ $O(n\cdot\max\{d,\lambda\})$ $\tilde{O}(1/(n\epsilon^{2})^{\frac{2}{d+2}})$
+We use Shapley value to measure the contribution of each gradient report.
+We define the utility function of Shapley payoff as cosine similarity between aggregated private gradient and the true gradient $\texttt{grad}_{val}$ (which is the average gradient over the validation dataset):
+$U(S)=\frac{<\texttt{grad}_{val},\sum_{i\in S}g_{i}>}{\|\texttt{grad}_{val}\|_{2}\cdot\|\sum_{i\in S}g_{i}\|_{2}},$
+so as to efficiently approximate the negative loss function over the validation dataset.
+We note that one can use other Shapley-value utility functions within our PIC model, we choose this one for efficiency.
+Then, the Shapley value of one single gradient update $g_{i}$ is computed as follows:
+$\texttt{Shapley}_{i}=\frac{1}{n}\sum_{k=1}^{n}{n\choose k-1}\sum_{\begin{subarray}{c}S\subseteq[n]\backslash\{i\},\\ |S|=k-1\end{subarray}}U(S\cup\{i\})-U(S).$
+In cases the server receives sanitized gradient information, the corresponding sanitized version of Shapley values is computed in the same way, except using sanitized gradient in the utility function.
+The Shapley values themselves may severely leak sensitive information about user data [64].
+Despite the vast non-linear computations involved in evaluating Shapley values with neural networks, no rigorous protection has been offered for federated learning with incentives in existing literature.
+In this study, we have demonstrated the feasibility of computing Shapley value without incurring additional privacy loss.
+Appendix F Comparison of Private Permutation-equivariant Multi-party Computation
+We perform comparisons across a broader range of permutation-equivariant computation tasks.
+These tasks include bipartite matching and federated learning with incentives.
+We detail these comparisons in Table 9, emphasizing general permutation-equivariant computation tasks characterized by $C$ arithmetic circuit gates and $h$ circuit depth.
+Specifically:
+  • The nearest neighboring task aims to find $k$ nearest neighbors for each party, the $C$ is of the order $n^{3}$, and the $h$ is of the order $\log n$.
+  • In minimum weight bipartite matching using LAPJVsp algorithm [54], both $C$ and $h$ are of the order $n^{3}$ where $n$ is the number of nodes in the bipartite graph.
+  • In maximum bipartite matching using Hopcroft-Karp algorithm [48], $C$ is of the order $E\sqrt{n}$, $h$ is of the order $\sqrt{n}$, where $E$ is the number of edges in the bipartite graph.
+  • For federated learning with Shapley incentives, which uses accuracy on the validation dataset as the utility function, $C=n_{val}\cdot N\cdot M$ can become exceptionally large, where $N$ isthe number of neurons in neural networks, $n_{val}$ is the number of samples in the validation dataset, and $M$ is the number of Monte Carlo evaluations.
+The $h$ is the depth of the neural network.
+We note that the Ishai’s MPC model [51] and Beimel’s MPC model [10] also use a role of message shuffler, but their shufflers must be fully trustable.
+In general, MPC-based approaches impose high computation and interaction overheads, especially when the computation algorithm gets sophisticated.
+As comparison, the proposed paradigm in this work permits running arbitrary algorithms on plaintext, and requires only semi-trustness/honest-but-curious assumption on the shuffler (as the shuffler only sees ciphertexts, see Section 5).
+Our PIC model does not necessitate algorithmic computation on the user side.
+Additionally, it permits the orchestrating server to operate on plaintext, making it considerably more efficient in computation and communication compared to MPC-based methods.
+In fact, the cost of the model grounded on hybrid encryption (covered in Section 5) virtually mirrors non-private settings that maintain secure communications, such as those employing HTTPS.
+Besides, our model is compatible with existing achievements and future advancements in server-side algorithms (e.g., noise-aware matching algorithms for spatial crowdsourcing and combinatorial optimization).
+Appendix G Security Proof
+This part provides security proof for the PIC protocol in §5.
+G.1 Protocol Setting and Security Goals
+Protocol setting.
+Our permutation-invariant computation (PIC) protocol involves $m$ groups of clients and one single computing server $S$.
+These $m$ groups of parties want to jointly compute some pre-defined computing task, formalized as a multi-input function $f$, with the help of $S$.
+At the end of the protocol, each client may receive a function output.
+Security definitions and goals.
+We follow the simulation-based security model with semi-honest adversaries, who will faithfully follow the protocol specifications but try to learn more information than allowed through protocol interaction.
+Security goals are formally captured as an ideal functionality $\mathcal{F}$.
+$\mathcal{F}$ receives inputs from the parties, performs computation, and sends the computation result back to the parties.
+Roughly, the security goals include privacy and correctness.
+Privacy requires that the adversary can only learn information as allowed but nothing more, while correctness requires that the computation is done correctly.
+We note that correctness is easy to achieve for semi-honest protocols.
+Remark on privacy guarantee.
+It’s known that security proof for a secure computation protocol demonstrates that no additional information about parties’ inputs is revealed, except the computation output and any allowed/inherent leakage, which is well-captured in the definition of an ideal functionality.
+However, we note the computation output (combined with captured leakage) may contain a significant amount of private information about parties’ inputs.
+Our PIC protocol achieves differential privacy, which provides a trade-off between privacy and utility.
+For this part, we refer to §5.2 for a formal analysis of the DP guarantees offered by our protocols.
+G.2 Ideal Functionality
+The ideal shuffle functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$.
+The shuffle functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ receives $n$ inputs from $n$ input providers and outputs $n$ randomly permuted outputs of the original inputs.
+Possible instantiations of $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ include trusted hardware and securely evaluating a permutation network using MPC.
+The ideal Permutation-Invariant-Computation functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$.
+The ideal PIC functionality captures the core features of our permutation invariant computation.
+It receives inputs from $m$ groups of parties, adds noise to each input, randomly shuffles inputs of each group, and performs computation over the noisy inputs.
+At the end of the protocol, $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ sends each party a computation result, and $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ additionally sends all randomized inputs and all function outputs to the server.
+G.3 Proof
+We prove the security of our protocol in the static corruption setting, where the adversary specifies the corruption parties before running the protocol.
+Let $\mathcal{C}$ be the collection of corrupted parties and $\mathcal{C}\subset G_{1}\cup G_{2}\cdots G_{n}\cup\{S\}$, and $\mathcal{H}=G_{1}\cup G_{2}\cdots G_{n}\cup\{S\}\setminus\mathcal{C}$ be the remaining honest parties.
+The proof is simulation-based.
+It shows that for any PPT adversary $\mathcal{A}$ who corrupts a set of parties, there exists a PPT simulator $\mathcal{S}$ that can generate a simulated view indistinguishable from the view of real-world execution.
+Note that the simulator learns the corrupted parties’ input and output (i.e., computation result and allowed leakage) for the view simulation in the semi-honest security model.
+In this proof, we assume some existing ideal functionalities (e.g., $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$) in the hybrid model, and we directly use existing simulators for these functionalities when needed.
+Depending on whether $S\in\mathcal{C}$, we separately prove the security of our PIC protocol in two cases as follows:
+Case 1: $S\notin\mathcal{C}$. In this case, all corrupted parties are clients. We construct a simulator $\mathcal{S}$ for view simulation as follows:
+  1. 1.
+    The simulator $\mathcal{S}$ sets up global parameters of the system as the real protocol execution, including the specification of a public key encryption scheme $\Pi=(\textsf{Gen},\textsf{Enc},\textsf{Dec})$, security parameter $\lambda$, the server’s public key $pk_{c}$ from invoking $\textsf{Gen}(\lambda)$, each client groups $G_{i}$ $(i\in[m])$, and the data randomization mechanisms $\mathcal{R}_{i}$ for each group.
+$\mathcal{S}$ generates public keys for all parties.
+  2. 2.
+    For the $j$-th client $u_{i,j}$ in group $G_{i}$, if $u_{i,j}$ is corrupted by $\mathcal{A}$, $\mathcal{S}$ runs as the real protocol execution: It randomizes $u_{i,j}$’s input $x_{i,j}$ with mechanism $\mathcal{R}_{i}$ and obtains $x_{i,j}^{\prime}=\mathcal{R}_{i}(x_{i,j})$.
+Then the sanitized input is concatenated with $u_{i,j}$’s public key $pk_{i,j}$, and encrypted with the server’s public key $x_{i,j}^{\prime\prime}=\textsf{Enc}_{pk_{c}}(pk_{i,j}||x^{\prime}_{i,j})$.
+If $u_{i,j}$ is not corrupted, $\mathcal{S}$ generates a ciphertext $x_{i,j}^{\prime\prime}=\textsf{Enc}_{pk_{c}}(\mathbf{0})$ from the ciphertext domain (of the public encryption scheme $\Pi$) as the simulated ciphertext from $u_{i,j}$, where $\mathbf{0}$ denotes a vector of 0s of equal length as $pk_{i,j}||x^{\prime}_{i,j}$.
+  3. 3.
+    $\mathcal{S}$ invokes $\mathcal{S}_{\scriptscriptstyle\textsf{Shuffle}}$ to simulate the view involved in the shuffle protocol.
+In particular, $\mathcal{S}$ has ciphertexts $\{x_{i,j}^{\prime\prime}\}_{i\in[m],j\in[n_{i}]}$ as the input to $\mathcal{S}_{\scriptscriptstyle\textsf{Shuffle}}$.
+Additionally, $\mathcal{S}$ generates ciphertexts $\{\widetilde{x}_{i,j}^{\prime\prime}\}_{i\in[m],j\in[n_{i}]}$, where each $\widetilde{x}_{i,j}^{\prime\prime}$ is generated as follows: For each corrupted party $u_{i,j}\in\mathcal{C}$, $\mathcal{S}$ learns $\pi_{i}(j)$ from the leakage profile $\mathcal{L}$.
+Then $\mathcal{S}$ generates $\widetilde{x}^{\prime\prime}_{i,\pi_{i}(j)}$ by randomizing ciphertext $\widetilde{x}^{\prime\prime}_{i,j}$.
+For $\widetilde{x}_{i,j}^{\prime\prime}$ corresponding to an honest party $u_{i,j}$, $\mathcal{S}$ simply generates $\widetilde{x}_{i,j}^{\prime\prime}=\textsf{Enc}_{pk_{c}}(\mathbf{0})$.
+$\mathcal{S}$ then invokes $\mathcal{S}_{\scriptscriptstyle\textsf{Shuffle}}(\{x_{i,j}^{\prime\prime}\}_{i\in[m],j\in[n_{i}]},$$\{\widetilde{x}_{i,j}^{\prime\prime}\}_{i\in[m],j\in[n_{i}]}$, $\mathcal{L})$ to generate the view for the shuffle phase, and appends it as a part of the simulated view for $\mathcal{S}$.
+  4. 4.
+    The last piece of simulation is to ensure consistency between the prior view and the computation result of $f$.
+To this end, $\mathcal{S}$ works as follows: For each corrupted party $u_{i,j}\in\mathcal{C}$, $\mathcal{S}$ obtains $y_{i,\pi_{i}(j)}$ from $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ and $\pi_{i}(j)$ (from the leakage profile $\mathcal{L}$).
+It then generates $(pk_{i,\pi_{i}(j)},\textsf{Enc}_{pk_{i,\pi_{i}(j)}}(y_{i,\pi_{i}(j)})$, which is then arranged as the $j$-th encrypted output within group ${G}_{i}$.
+If party $u_{i,j}$ is not corrupted, $\mathcal{S}$ generates $(pk_{i,j},c_{i,j})$, where $c_{i,j}=\textsf{Enc}(\mathbf{0})$ is ciphertext of $\mathbf{0}$ with the same length as $y_{i,j}$.
+Below, we show the simulated view is indistinguishable from real protocol execution via the following hybrid games.
+$\mathbf{G_{1}}$: This is the real protocol execution.
+$\mathbf{G_{2}}$: $\mathbf{G_{2}}$ is same as $\mathbf{G_{1}}$, except the following difference: For all ciphertexts corresponding to the honest clients, $\mathbf{G_{2}}$ generates the ciphertexts as the encryption of $\mathbf{0}$ with proper length.
+Due to the IND-CPA security of the public encryption scheme $\Pi$, $\mathbf{G_{2}}$ is computationally indistinguishable from $\mathbf{G_{1}}$.
+$\mathbf{G_{3}}$: $\mathbf{G_{3}}$ is same as $\mathbf{G_{2}}$, except the following difference: $\mathbf{G_{3}}$ uses the leakage profile $\mathcal{L}$ to arrange the ciphertexts that should be outputted from the shuffling phase.
+In particular, for each ciphertext corresponding to a corrupted party $j$ of group $i$ after the shuffle phase, $\mathbf{G}_{3}$ puts the input ciphertext to the coordinate $\pi_{i}(j)$ and randomizes the ciphertext.
+The view distribution is identical to $\mathbf{G}_{2}$.
+$\mathbf{G_{4}}$: $\mathbf{G_{4}}$ is same as $\mathbf{G_{3}}$, except the following difference: $\mathbf{G_{4}}$ invokes the simulator $\mathcal{S}_{\scriptscriptstyle\textsf{Shuffle}}$ to simulate the view corresponding to the view of shuffling.
+Since our protocol works in the hybrid model, $\mathbf{G_{4}}$ is identical to the view from $G_{3}$.
+Also note that $\mathbf{G}_{4}$ works the same as the simulator $\mathcal{S}$.
+Overall, the view generated by $\mathcal{S}$ is computationally indistinguishable from the view of real protocol execution in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model.
+Case 2: $S\in\mathcal{C}$. The server $S$ is also corrupted in this case. We construct a simulator $\mathcal{S}$ for view simulation as follows:
+  1. 1.
+    The simulator $\mathcal{S}$ sets up global parameters of the system as the server does in the real protocol execution, including the specification of a public key encryption scheme $\Pi=(\textsf{Gen},\textsf{Enc},\textsf{Dec})$, security parameter $\lambda$, the server’s public key $pk_{c}$ from invoking $\textsf{Gen}(\lambda)$, each client groups $G_{i}$ $(i\in[m])$, and the data randomization mechanisms $\mathcal{R}_{i}$ for each group.
+$\mathcal{S}$ generates public keys for all parties like the real protocol execution.
+  2. 2.
+    To simulate the view corresponding to the encrypted computation output, $\mathcal{S}$ receives $(L,f(L))$ from the output of $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$.
+$\mathcal{S}$ pareses $f(L)$ as ($(\{z_{1,j}\}_{j},\cdots,\{z_{m,j}\}_{j})$; note that ($(\{z_{1,j}\}_{j},\cdots,\{z_{m,j}\}_{j})$ corresponds to $\{y_{1,\pi_{1}(j)}\}_{j\in[n_{1}]},\cdots,\{y_{m,\pi_{m}(j)}\}_{j\in[n_{m}]}$ in the real protocol execution; here $\mathcal{S}$ doesn’t know the involved permutations $\{\pi_{i}\}_{i\in[m]}$, except the information from the leakage profile $\mathcal{L}$.
+$\mathcal{S}$ performs simulation as follows.
+    • For a corrupted party $u_{i,j}$, $\mathcal{S}$ generates a ciphertext $(pk_{i,j},$ $\textsf{Enc}_{pk_{i,j}}(z_{i,\pi^{-1}_{i}(j)}))$ that will be placed to the $\pi^{-1}_{i}(j)$-th encrypted output in group $G_{i}$.
+    • For all other non-corrupted parties, $\mathcal{S}$ randomly selects permutations $\{\widetilde{\pi}_{i}\}_{i\in[m]}$ under the constraint of the leakage profile $\mathcal{L}$, which means that $\widetilde{\pi}_{i}(j)=\pi_{i}(j)$ for each corrupted party $u_{i,j}$, where $\{\pi_{i}\}_{i\in[m]}$ is the collection of secret permutations used for shuffling in the real protocol execution.
+Using $\{\widetilde{\pi}_{i}\}_{i\in[m]}$ and honest parties’ public keys, $\mathcal{S}$ can encrypt the computation results corresponding to the honest parties accordingly.
+Specifically, $\mathcal{S}$ generates $(pk_{i,j},\textsf{Enc}_{pk_{i,j}}(z_{i,\widetilde{\pi}^{-1}_{i}(j)}))$ that will be placed as the $\widetilde{\pi}^{-1}_{i}(j)$-th encrypted output in group $G_{i}$.
+  3. 3.
+    To simulate the view corresponding to the shuffling phase, $\mathcal{S}$ first obtains (shuffled) input lists $L$; these shuffled lists are revealed to the server in real protocol execution, and here $\mathcal{S}$ obtains the information from the ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$.
+Then, for the $j$-th client $u_{i,j}$ in group $G_{i}$, to simulate the input ciphertexts before the shuffling, the simulator $\mathcal{S}$ does the following:
+    • If $u_{i,j}$ is corrupted, $\mathcal{S}$ gets the sanitized, shuffled input $x_{i,\pi_{i}^{-1}(j)}^{\prime}$ from $L$.
+Then $x_{i,\pi_{i}^{-1}(j)}^{\prime}$ is concatenated with $u_{i,j}$’s public key $pk_{i,j}$, and encrypted with the server’s public key $x_{i,j}^{\prime\prime}=\textsf{Enc}_{pk_{c}}(pk_{i,j}||x^{\prime}_{i,\pi_{i}^{-1}(j)})$.
+    • If $u_{i,j}$ is non-corrupted, $\mathcal{S}$ gets the sanitized, shuffled input $x_{i,\widetilde{\pi}_{i}^{-1}(j)}^{\prime}$ from $L$.
+Then $x_{i,\widetilde{\pi}_{i}^{-1}(j)}^{\prime}$ is concatenated with $u_{i,j}$’s public key $pk_{i,j}$, and encrypted with the server’s public key $x_{i,j}^{\prime\prime}=\textsf{Enc}_{pk_{c}}(pk_{i,j}||x^{\prime}_{i,\widetilde{\pi}_{i}^{-1}(j)})$.
+    Clearly, $\mathcal{S}$ uses $\{\widetilde{\pi}_{i}\}_{i\in[m]}$ to arrange these ciphertexts in the above simulation.
+The generated ciphertexts above serve as the inputting ciphertexts for the shuffling phase.
+To simulate the output ciphertexts after the shuffling, $\mathcal{S}$ simply permutes all inputs ciphertexts using $\{\widetilde{\pi}_{i}\}_{i\in[m]}$, and re-randomizes each ciphertext.
+$\mathcal{S}$ then invokes the simulator $\mathcal{S}_{\scriptscriptstyle\textsf{Shuffle}}$ over the inputting ciphertexts, the outputting ciphertexts, and the leakage profile $\mathcal{L}$.
+$\mathcal{S}$ appends the generated view from $\mathcal{S}_{\scriptscriptstyle\textsf{Shuffle}}$ as a part of its own simulation.
+Below, we show the simulated view is indistinguishable from real protocol execution via the following hybrid games.
+$\mathbf{G_{1}}$: This is the real protocol execution.
+$\mathbf{G_{2}}$: $\mathbf{G_{2}}$ is same as $\mathbf{G_{1}}$, except the following difference: $\mathbf{G_{2}}$ uses randomly sampled permutations $\{\widetilde{\pi}_{i}\}_{i\in[m]}$ with the constraint that $\widetilde{\pi}_{i}(j)=\pi_{i}(j)$ for a corrupted party $u_{i,j}$ for $i\in[m]$.
+The security of secure shuffling ensures that the difference from $\widetilde{\pi}_{i}(j)\neq\pi_{i}(j)$ for any non-corrupted party $u_{i,j}$ in $\mathbf{G}_{2}$ is indistinguishable from $\mathbf{G}_{1}$.
+$\mathbf{G_{3}}$: $\mathbf{G_{3}}$ uses the function output $f(L)$ and $\{\widetilde{\pi}_{i}\}_{i\in[m]}$ to generated the final encrypted output for each party.
+For all other parts, $\mathbf{G_{3}}$ remains the same as $\mathbf{G_{2}}$.
+In particular, $\mathbf{G_{3}}$ encrypts the desired outputs for all parties using $f(L)$ and $\{\widetilde{\pi}_{i}\}_{i\in[m]}\}$, as done in Step (2) of $\mathcal{S}$.
+$\mathbf{G_{3}}$ is perfectly indistinguishable from $\mathbf{G_{2}}$.
+$\mathbf{G_{4}}$: $\mathbf{G_{4}}$ uses the shuffled output $L$ and $\{\widetilde{\pi}_{i}\}_{i\in[m]}$ to generated ciphertexts before and after the shuffling.
+For all other parts, $\mathbf{G_{4}}$ remains the same as $\mathbf{G_{3}}$.
+In particular, $\mathbf{G_{4}}$ encrypts the desired outputs for all parties using $L$ and $\{\widetilde{\pi}_{i}\}_{i\in[m]}\}$, as done in Step (3) of $\mathcal{S}$.
+$\mathbf{G_{4}}$ is perfectly indistinguishable from $\mathbf{G_{3}}$.
+$\mathbf{G_{5}}$: The only difference between $\mathbf{G_{5}}$ and $\mathbf{G_{4}}$ is that $\mathbf{G_{5}}$ invoke $\mathcal{S}_{\scriptscriptstyle\textsf{Shuffle}}$ to simulate the view for the shuffling phase.
+Note that $\mathbf{G_{5}}$ works exactly as the simulator $\mathcal{S}$.
+$\mathbf{G_{5}}$ is perfectly indistinguishable from $\mathbf{G_{4}}$.
+Overall, the view generated by $\mathcal{S}$ is perfectly indistinguishable from the view of real protocol execution in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model.
+Intuitively, the only secret information when $S$ is corrupted is the hidden part of the permutations $\{\pi_{i}\}_{i\in[m]}$ corresponding to each honest party $u_{i,j}$ for all $i\in[m]$ and $j\in[|\mathbf{G}_{j}|]$, which can be perfectly simulated in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model.
+Appendix H Privacy Proof for Theorem 5.2
+Our proof frequently uses the data processing inequality of DP.
+The data processing inequality is considered a key feature of distance measures (e.g., Hockey-stick divergence) used for evaluating privacy.
+It asserts that the privacy guarantee cannot be weakened by further analysis of a mechanism’s output.
+Definition H.1 (Data processing inequality).
+A distance measure $D:\Delta(\mathcal{S})\times\Delta(\mathcal{S})\to[0,\infty]$ on the space of probability distributions satisfies the data processing inequality if, for all distributions $P$ and $Q$ in $\Delta(\mathcal{S})$ and for all (possibly randomized) functions $g:\mathcal{S}\to\mathcal{S^{\prime}}$,
+$D(g(P)\|g(Q))\leq D(P\|Q).$
+The situations of victim user $u_{i^{*},j^{*}}$’s privacy can be categorized into three case: (1) $u_{i^{*},j^{*}}$ himself/herself is corrupted; (2) $u_{i^{*},j^{*}}$ is not corrupted and the server $S$ is a corrupted party (i.e., $S\in C$); (3) neither $u_{i^{*},j^{*}}$ nor the server $S$ is corrupted.
+In the first case, protecting the privacy of $u_{i^{*},j^{*}}$ against himself/herself is trivial.
+In the second case, a key observation is that the received computation results of all corrupted parties (including $S$) in $C$ are rooted from the server’s received messages $\{x_{i,\pi_{(}j)}^{\prime\prime}\}_{i\in[m],j\in[n_{i}]}$ in Step (4).
+We let variable $Y^{\prime}$ denote the server’s received messages $\{x_{i,\pi_{(}j)}^{\prime\prime}\}_{i\in[m],j\in[n_{i}]}$ when the input datasets are $X^{\prime}$, and let variable $Y$ denote the received messages when the input datasets are $X$.
+Then, for any distance measure $D$ that satisfies the data processing inequality, we readily have the divergence level of the adversaries’ view (i.e.
+$\mathcal{A}(X)$ when the input datasets are $X$, and $\mathcal{A}(X^{\prime})$ when the input datasets are $X^{\prime}$) is upper bounded by the one about the server’s received messages:
+$D(\mathcal{A}(X)\|\mathcal{A}(X^{\prime}))\leq D(Y\|Y^{\prime}).$ (5)
+We now focus on the DP guarantee of the server’s received messages in Step 4 (i.e.
+the divergence $D(Y\|Y^{\prime})$).
+For neighboring datasets $X$ and $X^{\prime}$ that differ at (and only at) the user $u_{i^{*},j^{*}}$ from group $i^{*}$, we have the $D(Y\|Y^{\prime})$ upper bounded by the divergence of shuffled messages from uncorrupted users in group $X_{i^{*}}$.
+To be precise, we let $\hat{Y}_{i^{*}}$ denote the shuffled messages $\mathcal{S}(\{x_{i^{*},j}^{\prime\prime}\}_{u_{i^{*},j}\in G_{i*}\backslash C_{i^{*}}})$ of group $i^{*}$ when the input dataset is $X$, and let $\hat{Y}^{\prime}_{i^{*}}$ denote the shuffled messages $\mathcal{S}(\{x_{i^{*},j}^{\prime\prime}\}_{u_{i^{*},j}\in G_{i*}\backslash C_{i^{*}}})$ when the input dataset is $X^{\prime}$.
+Then for any distance measure that satisfies the data processing inequality, we have $D(Y\|Y^{\prime})\leq D(\hat{Y}_{i^{*}}\|\hat{Y}^{\prime}_{i^{*}})$ (see Lemma H.2 for proof).
+Lemma H.2.
+For neighboring datasets $X$ and $X^{\prime}$ that differ at (and only at) the user $u_{i^{*},j^{*}}$ from group $i^{*}$, we have:
+$D(Y\|Y^{\prime})\leq D(\hat{Y}_{i^{*}}\|\hat{Y}^{\prime}_{i^{*}}).$
+Proof.
+Without loss of generality (for both corrupted/malicious and uncorrupted users), we let $\mathcal{R}_{i,j}$ denote the (possibly random) local message generation function of user $u_{i,j}$ in Step (2), which takes as input the local information (i.e.
+$x_{i,j}$, $pk_{i,j}$, $sk_{i,j}$) and global parameters.
+To prove the result, we define the following procedures (denoted as $post$) that takes as input the shuffled messages $\mathcal{S}(\{x_{i^{*},j}^{\prime\prime}\}_{u_{i^{*},j}\in G_{i^{*}}\backslash C_{i^{*}}})$ and outputs $\mathcal{S}(\{x_{i,j}^{\prime\prime}\}_{i\in[m],u_{i,j}\in G_{i}})$.
+  • For $u_{i^{*},j}\in C_{i^{*}}$, running $\mathcal{R}_{i^{*},j}$ to obtain $x_{i^{*},j}^{\prime\prime}$;
+  • Uniformly sample a permutation $\pi^{\prime}_{i^{*}}:C_{i^{*}}\mapsto[n_{i^{*}}]$, then construct a $n_{i^{*}}$-length list $Y_{i^{*}}$ such that every $x_{i^{*},j}^{\prime\prime}$ (for corrupted users $u_{i^{*},j}\in C_{i^{*}}$) asides at the $\pi^{\prime}_{i^{*}}(j)$-th position, and then inputted messages $\mathcal{S}(\{x_{i^{*},j}^{\prime\prime}\}_{u_{i^{*},j}\in G_{i^{*}}\backslash C_{i^{*}}})$ sequentially fill up the list;
+  • For all $i\in[m]\backslash\{i^{*}\}$ and all possible $j\in G_{i}$, running $\mathcal{R}_{i,j}$ to obtain $x_{i,j}^{\prime\prime}$;
+  • For all $i\in[m]\backslash\{i^{*}\}$, uniformly sample a permutation $\pi_{i}:G_{i}\mapsto[n_{i}]$ and get $\{x_{i,\pi_{i}(j)}^{\prime\prime}\}_{j\in[n_{i}]}$;
+  • Initialize $Y_{temp}$ as $Y_{i^{*}}$.
+For $i\in[i^{*}-1]$, prepend $\{x_{i,\pi_{i}(j)}^{\prime\prime}\}_{j\in[n_{i}]}$ to $Y_{temp}$; for $i\in[i^{*}+1,m]$, append $\{x_{i,\pi_{i}(j)}^{\prime\prime}\}_{j\in[n_{i}]}$ to $Y_{temp}$;
+  • Return $Y_{temp}$.
+It is oblivious that: (1) when the input dataset is $X$, the $post(\hat{Y}_{i})$ distributionally equals to $Y$; (2) when the input dataset $X^{\prime}$, the $post(\hat{Y}_{i}^{\prime})$ distributionally equals to $Y^{\prime}$.
+Therefore, according to the data processing inequality, we have the conclusion.
+∎
+We proceed to analyze the DP guarantee of shuffled messages from uncorrupted users in group $i^{*}$ (i.e.
+the divergence $D(\hat{Y}_{i^{*}})\|\hat{Y}^{\prime}_{i^{*}})$).
+To this end, we firstly summarize the local message generating procedures aside on each uncorrupted client’s side as an abstracted mechanism, then show the extended & encrypted message outputted by the abstracted mechanism has the same local privacy guarantee as the data randomizer $\mathcal{R}$, and show all clients in the same group follow an identical abstracted mechanism.
+Finally, we show the privacy amplification via shuffling can be applied to this identical local mechanism, as if purely shuffling data randomized by $\mathcal{R}$.
+An abstraction for local client-side procedures.
+Recall that the only message leaving a uncorrupted client $j$ (i.e.
+$u_{i^{*},j}\in G_{i^{*}}\backslash C_{i^{*}}$) in Phases (0)-(4) is (fingerprint/public key $f_{i^{*},j}$, sanitized data $x^{\prime}_{i^{*},j}$, extra data $t_{i^{*},j}$), where the fingerprint $f_{i^{*},j}=F(sk_{i^{*},j})$ is post-processed from the random private key $sk_{i^{*},j}$ using some function $F$, the sanitized data $x^{\prime}_{i^{*},j}$ is the privatized version of secret data $x_{i^{*},j}=((i^{*},j),l_{i^{*},j},v_{i^{*},j})$ via a privatization mechanism $\mathcal{R}$, and the extra data $t_{i^{*},j}$ is post-processed as $T(sk_{i^{*},j},f_{i^{*},j},x^{\prime}_{i^{*},j})$ using some function $T$.
+We summarize and abstract these procedures in Algorithm 1.
+Params : A domain $\mathbb{D}_{k}$, a distribution $P_{k}:\mathbb{D}_{k}\mapsto[0,1]$, any function $F:\mathbb{D}_{k}\mapsto\mathbb{D}_{f}$, any local mechanism $\mathcal{R}:\mathbb{X}\mapsto\mathbb{Y}$, and any function $T:\mathbb{D}_{k}\times\mathbb{D}_{f}\times\mathbb{Y}\mapsto\mathbb{D}_{z}$.
+Input: An input $x_{i^{*},j}\in\mathbb{X}$.
+Output: An extended message in Phase (1).
+1
+[…]
+3sample $sk_{i^{*},j}\sim P_{k}$
+[…]
+[…]
+return $(f_{i^{*},j},\ x^{\prime}_{i^{*},j},\ t_{i^{*},j})$
+Algorithm 1 An abstracted local mechanism $\overline{\mathcal{R}}$
+Identicalness of local mechanisms.
+Recall that privacy amplification via shuffling requires, not only that each local data is randomized, but also that clients follow an identical randomization mechanism (otherwise anonymity might break due to the output domain/distribution difference in different randomizers).
+In Lemma H.3, we show that the messages from every honest client (in the same group) are indeed randomized by an identical mechanism (i.e., the Algorithm 1).
+The proof of the lemma is trivial since every honest client in the same group adopted the same global parameters and the same local randomizer since Phase (0).
+Lemma H.3.
+For each honest/uncorrupted client in the same group, Phase (1) in the conceptual protocol from Section 5 is equivalent to Algorithm 1 with global parameters given in Phase (0).
+Privacy amplification guarantees.
+Since every message from clients in the same group is sanitized by an identical mechanism $\overline{\mathcal{R}}$ (see the former part), and the server (i.e., the potential privacy adversary) only observe the shuffled messages from each group, one can directly apply the amplification bounds in the literature [7, 34, 35, 92].
+Generally, for any distance measure $D$ that satisfies the data processing inequality, we have the corresponding distance between $\hat{Y}_{i^{*}})$ and $\hat{Y}^{\prime}_{i^{*}})$, is upper bounded by the distance between $\mathcal{S}\circ\mathcal{R}(\{x_{i^{*},j}\}_{u_{i,j}\in G_{i^{*}}\backslash C_{i^{*}}})$ and $\mathcal{S}\circ\mathcal{R}(\{x^{\prime}_{i^{*},j}\}_{u_{i,j}\in G_{i^{*}}\backslash C_{i^{*}}})$ (see Lemma H.4).
+Lemma H.4 (Privacy Guarantee of Shuffled Algorithm 1).
+Given two neighboring dataset $\hat{X}_{i^{*}}=\{x_{i^{*},j}\}_{u_{i,j}\in G_{i^{*}}\backslash C_{i^{*}}}$ and $\hat{X}^{\prime}_{i^{*}}=\{x^{\prime}_{i^{*},j}\}_{u_{i,j}\in G_{i^{*}}\backslash C_{i^{*}}}$ that differ only at one element, then for any distance measure $D$ that satisfies the data processing inequality,
+$D(\mathcal{S}\circ\overline{\mathcal{R}}(\hat{X}_{i^{*}})\|\mathcal{S}\circ\overline{\mathcal{R}}(\hat{X}^{\prime}_{i^{*}}))\leq D(\mathcal{S}\circ\mathcal{R}(\hat{X}_{i^{*}})\|\mathcal{S}\circ\mathcal{R}(\hat{X}^{\prime}_{i^{*}})).$
+Proof.
+Given input either $\mathcal{S}\circ\mathcal{R}(\hat{X}_{i^{*}})$ or $\mathcal{S}\circ\mathcal{R}(\hat{X}^{\prime}_{i^{*}})$, we define a following function (denoted as $g_{sr}$):
+  • For each message $x^{\prime}_{i^{*},j}$ in the input, sample $sk_{i^{*},j}\sim P_{k}$, compute $f_{i^{*},j}\leftarrow F(sk_{i^{*},j})$, then compute $t_{i^{*},j}\leftarrow T(sk_{i^{*},j},f_{i^{*},j},x^{\prime}_{i^{*},j})$, and finally get $(f_{i^{*},j},x^{\prime}_{i^{*},j},t_{i^{*},j})$.
+  • Return a sequential of $(f_{i^{*},j},x^{\prime}_{i^{*},j},t_{i^{*},j})$ each is generated by the previous step over each input message.
+It is obvious that the $g_{sr}(\mathcal{S}\circ\mathcal{R}(\hat{X}_{i*}))$ distributionally equals to $\mathcal{S}\circ\overline{\mathcal{R}}(\hat{X}_{i*})$, and $g_{sr}(\mathcal{S}\circ\mathcal{R}(\hat{X}^{\prime}_{i*}))$ distributionally equals to $\mathcal{S}\circ\overline{\mathcal{R}}(\hat{X}^{\prime}_{i*})$.
+Therefore, according to the data processing inequality of the distance measure of $D$, we have the conclusion.
+∎
+In particular, when the mechanism $\mathcal{R}$ satisfies $\epsilon$-LDP, we have $\overline{\mathcal{R}}$ satisfies $\epsilon$-LDP.
+Then according to the latest work [35], the shuffled messages $\mathcal{S}(\{\overline{\mathcal{R}}(x_{i^{*},j})\}_{u_{i,j}\in G_{i^{*}}\backslash C_{i^{*}}})$ from group $G_{i^{*}}$ satisfy $(\epsilon_{c},\delta)$-DP where
+$\epsilon_{c}=\texttt{Amplify}(\epsilon,\delta,|G_{i^{*}}\backslash C_{i^{*}}|)=\tilde{O}(\sqrt{e^{\epsilon}/n^{\prime}_{i^{*}}}).$
+Specifically, when $n\geq 8(e^{\epsilon}+1)\log(2/\delta)$, we have $\epsilon_{c}\leq\log\Big{(}1+\frac{e^{\epsilon}-1}{e^{\epsilon}+1}(\sqrt{\frac{64e^{\epsilon}\log 4/\delta}{n^{\prime}_{i^{*}}}}+\frac{8e^{\epsilon}}{n^{\prime}_{i^{*}}})\Big{)}$ [34, Theorem 3.1].
+This implies the conclusion of the second case.
+Finally, let us consider the third case.
+Simply add the server $S$ to corrupted parties $C$ (i.e.
+more views are leaked to adversaries), and apply the result for the second case, we arrive at the conclusion.
+Appendix I Proof of Error Lower Bounds in PIC Model
+We firstly shift our focus from the original $\ell_{2}$-norm spherical domain $\mathbb{S}_{2,1}({0}^{d})$ to the simpler $\ell_{+\infty}$-norm domain $[0,1]^{d}$.
+To establish the lower bound of the mean square error (MSE) of a single report derived from the single-message shuffle model with $x_{i}\in[0,1]^{d}$ as input, we follow a four-step approach.
+First, we confine the space of local randomizers to the subspace where the domain of the randomizer’s output matches the input.
+Second, we construct a set of discretized and gridded inputs that are well separated.
+Third, we compute their expected MSE using a formula that hinges on the probability transition matrix among these discrete inputs.
+Finally, we leverage the constraints of differential privacy inherent to the shuffle model to ascertain the properties of the probability transition matrix, thereby deducing lower bounds.
+Step (1): We begin by demonstrating that the MSE bound of any local randomizer $R:[0,1]^{d}\mapsto\mathbb{R}^{d^{\prime}}$ coupled with any post-processing function $f:\mathbb{R}^{d^{\prime}}\mapsto\mathbb{R}^{d}$ is lower bounded by a certain local randomizer $R^{\prime}:[0,1]^{d}\mapsto[0,1]^{d}$.
+This local randomizer, $R^{\prime}$, possesses an output domain identical to the input.
+We establish it using a constructive method.
+Defining $R^{\prime}(x_{i})=\max(0,\min(1,f(R(x_{i})))$, where $\min$ and $\max$ execute coordinate-wise min/max truncation, we can say for any $x_{i}\in[0,1]^{d}$ that:
+$\mathbb{E}[\|f(R(x_{i}))-x_{i}\|_{2}^{2}]$
+$=$ $\int_{x\in\mathbb{R}^{d}}\mathbb{P}[f(R(x_{i}))=x]\cdot\|x-x_{i}\|_{2}^{2}\mathrm{d}x$
+$\geq$ $\int_{x\in\mathbb{R}^{d}}\mathbb{P}[f(R(x_{i}))=x]\cdot\|\max(0,\min(1,x))-x_{i}\|_{2}^{2}\mathrm{d}x$
+$\geq$ $\mathbb{E}[\|R^{\prime}(x_{i})-x_{i}\|_{2}^{2}].$
+As a result, we can now narrow our focus to local randomizers with an output domain of $[0,1]^{d}$.
+Step (2): We proceed by partitioning the domain $[0,1]^{d}$ into multi-dimensional grids, such that the center points of these grids are well-separated (i.e., have non-zero distances from each other).
+Specifically, each dimension is segmented into $L$ uniform intervals, where the $l$-th interval is defined as $[\frac{l-1}{L},\frac{l}{L})$ for $l\in[L-1]$, and the $L$-th interval is $[\frac{L-1}{L},1]$.
+Intervals across all dimensions divide the domain into grids or subdomains, yielding a total of $L^{d}$ grids.
+Each grid is indexed by the indices of its intervals in each dimension.
+For instance, the $(1,1,\ldots,1)$-th grid corresponds to the subdomain $[0,\frac{1}{L})\times[0,\frac{1}{L})\times\cdots\times[0,\frac{1}{L})\in[0,1]^{d}$.
+Each grid possesses a center point, with the center point of the $(l_{1},l_{2},\ldots,l_{d})$-th grid being $\big{[}\frac{l_{1}-1/2}{L},\frac{l_{2}-1/2}{L},\ldots,\frac{l_{d}-1/2}{L}\big{]}$ for $l_{1},\dots,l_{d}\in[L]^{d}$.
+We denote all center points as:
+$C=\Big{\{}\big{[}\frac{l_{1}-1/2}{L},\frac{l_{2}-1/2}{L},\ldots,\frac{l_{d}-1/2}{L}\big{]}\ |\ l_{1},\dots,l_{d}\in[L]^{d}\Big{\}}.$
+We use $G(l_{1},l_{2},\ldots,l_{d})$ to denote the subdomain of the $(l_{1},l_{2},\ldots,l_{d})$-th grid.
+Step (3): Following the outcome of Step (1), we examine any local randomizer $R:[0,1]^{d}\mapsto[0,1]^{d}$.
+Given any input $x_{i}\in C$, the randomizer $R$ defines a probabilistic transition from $x_{i}$ to each of the $L^{d}$ grids.
+We represent the transition probability from $x_{i}$ to the $\mathbf{l^{\prime}}$-th grid as $P_{\mathbf{l},\mathbf{l^{\prime}}}$, where $\mathbf{l}$ is the index of the grid that contains $x_{i}$ and $\mathbf{l},\mathbf{l^{\prime}}\in[L]^{d}$.
+By iterating over all possible $x_{i}\in C$, we obtain a transition probability matrix.
+Now, considering that $x$ is randomly sampled from $C$ in a uniform manner, our objective is to analyze the expected MSE of $R$ given $x$ as input.
+Given that a center point maintains a minimum Manhattan distance of $\frac{1}{2L}$ from other grids, we can lower bound the expected MSE as follows:
+$\mathbb{E}_{x\sim uniform(C)}[\|R(x)-x\|_{2}^{2}]$
+$\geq$ $\frac{1}{L^{d}}\sum_{\mathbf{l}\in[L]^{d}}\sum_{\mathbf{l^{\prime}}\in[L]^{d}}P_{\mathbf{l},\mathbf{l^{\prime}}}\cdot\llbracket\mathbf{l}\neq\mathbf{l^{\prime}}\rrbracket\cdot\frac{1}{(2L)^{2}}$
+$\geq$ $\frac{1}{L^{d}}\sum_{\mathbf{l}\in[L]^{d}}\frac{1-P_{\mathbf{l},\mathbf{l}}}{4L^{2}},$
+where $P_{\mathbf{l},\mathbf{l}}$ represents the probability that the output resides within the same interval as the input central point of the $\mathbf{l}$-th grid.
+Additionally, since the squared distance between two points from two different grids (e.g., from the $\mathbf{l}$-th and $\mathbf{l^{\prime}}$-th grid respectively) is at least $\frac{1}{L^{2}}\sum_{j\in[d]}\llbracket\mathbf{l}_{j}\neq\mathbf{l^{\prime}}_{j}\rrbracket(|\mathbf{l}_{j}-\mathbf{l^{\prime}}_{j}|-1/2)^{2}$, we lower bound the expected MSE as follows:
+$\mathbb{E}_{x\sim uniform(C)}[\|R(x)-x\|_{2}^{2}]$
+$\geq$ $\frac{1}{L^{d}}\sum_{\mathbf{l}^{\prime}\in[L]^{d}}\sum_{\mathbf{l}\in[L]^{d}}P_{\mathbf{l^{\prime}},\mathbf{l}}\cdot\frac{\sum_{j\in[d]}\llbracket\mathbf{l}_{j}\neq\mathbf{l^{\prime}}_{j}\rrbracket(|\mathbf{l^{\prime}}_{j}-\mathbf{l}_{j}|-1/2)^{2}}{L^{2}}$
+$\geq$ $\frac{1}{L^{d}}\sum_{\mathbf{l}^{\prime}\in[L]^{d}}\big{(}\min_{\mathbf{l^{\prime\prime}}\in[L]^{d}}P_{\mathbf{l^{\prime\prime}},\mathbf{l}}\big{)}\sum_{\mathbf{l}\in[L]^{d}}\frac{\sum_{j\in[d]}\llbracket\mathbf{l}_{j}\neq\mathbf{l^{\prime}}_{j}\rrbracket(|\mathbf{l^{\prime}}_{j}-\mathbf{l}_{j}|-1/2)^{2}}{L^{2}}$
+$\geq$ $\frac{1}{L^{d}}\sum_{\mathbf{l}^{\prime}\in[L]^{d}}\big{(}\min_{\mathbf{l^{\prime\prime}}\in[L]^{d}}P_{\mathbf{l^{\prime\prime}},\mathbf{l}}\big{)}\sum_{j\in[d]}\sum_{\mathbf{l}\in[L]^{d}}\frac{\llbracket\mathbf{l}_{j}\neq\mathbf{l^{\prime}}_{j}\rrbracket(|\mathbf{l^{\prime}}_{j}-\mathbf{l}_{j}|-1/2)^{2}}{L^{2}}$
+$\geq$ $\frac{1}{L^{d}}\sum_{\mathbf{l}^{\prime}\in[L]^{d}}\big{(}\min_{\mathbf{l^{\prime\prime}}\in[L]^{d}}P_{\mathbf{l^{\prime\prime}},\mathbf{l}}\big{)}\sum_{j\in[d]}\frac{L^{d-1}(L-1/L)}{48},$
+where the last step uses the fact that $\sum_{\mathbf{l}\in[L]^{d}}\frac{\llbracket\mathbf{l}_{j}\neq\mathbf{l^{\prime}}_{j}\rrbracket(|\mathbf{l^{\prime}}_{j}-\mathbf{l}_{j}|-1/2)^{2}}{L^{2}}\geq\frac{L^{d-1}(L-1/L)}{48}$ holds for all possible $\mathbf{l^{\prime}}$.
+The $\min_{\mathbf{l^{\prime\prime}}\in[L]^{d}}P_{\mathbf{l^{\prime\prime}},\mathbf{l}}$ denotes the minimum possible probability that the output falls within the same interval as the central point of the $\mathbf{l}$-th grid, given all possible input $\mathbf{l^{\prime\prime}}\in[L]^{d}$.
+Step (4): We now leverage the DP constraint in the shuffle model to establish a relationship between the following two probabilities:
+$p_{1}=P_{\mathbf{l},\mathbf{l}},$
+$p_{0}=P_{\mathbf{l^{\prime\prime}},\mathbf{l}}$
+when $\mathbf{l},\mathbf{l^{\prime\prime}}$.
+Let’s consider two central points $x,x^{\prime\prime}\in C$ such that $x$ and $x^{\prime\prime}$ belong to the $\mathbf{l}$-th and $\mathbf{l^{\prime\prime}}$-th grid respectively.
+We can then construct two neighboring datasets $T=(x^{\prime\prime},x^{\prime\prime},\ldots,x^{\prime\prime})$ and $T^{\prime\prime}=(x,x^{\prime\prime},\ldots,x^{\prime\prime})$ that both contain $n$ elements.
+Then, in two independent runs of $S\circ R(T)$ and $S\circ R(T^{\prime\prime})$, the corresponding probability that $S\circ R(T)$ (or $S\circ R(T^{\prime\prime})$) contains no elements within the $\mathbf{l}_{j}$-th interval, is constrained by $(\epsilon,\delta)$-differential privacy as follows (assuming $\delta<0.5$):
+$\mathbb{P}[S\circ R(T)\cap G(\mathbf{l})=\emptyset]\leq e^{\epsilon}\mathbb{P}[S\circ R(T^{\prime\prime})\cap G(\mathbf{l})=\emptyset]+\delta,$ (6)
+where $G(\mathbf{l})\subseteq[0,1]^{d}$ denotes the domain of the $\mathbf{l}$-th grid.
+Note that $\mathbb{P}[S\circ R(T)\cap G(\mathbf{l})=\emptyset]=(1-p_{0})^{n}$ and $\mathbb{P}[S\circ R(T)\cap G(\mathbf{l})=\emptyset]=(1-p_{1})(1-p_{0})^{n-1}$.
+Consequently, if $(1-p_{1})\leq 0.5e^{-\epsilon}$, then we have $p_{0}\geq(1/2-\delta)/n$ where $\delta<1/2$, since $(1-p_{0})^{n}>1/2+\delta>e^{\epsilon}(1-p_{1})+\delta\geq 1/2+\delta$ results in a contradiction [7, Lemma 4.5].
+This implies that either $(1-p_{1})>0.5e^{-\epsilon}$ or $p_{0}\geq(1/2-\delta)/n$ holds for all possible $\mathbf{l^{\prime\prime}},\mathbf{l}\in[L]^{d}$ (under the assumption that $\delta<1/2$).
+Therefore, we arrive at the expected MSE as follows:
+$\mathbb{E}_{x\sim uniform(C)}[\|R(x)-x\|_{2}^{2}]$
+$\geq$ $\frac{1}{L^{d}}\sum_{\mathbf{l}\in[L]^{d}}\min\Big{\{}\frac{1-p_{1}}{4L^{2}},\ \frac{dL^{d-1}(L-1/L)p_{0}}{48}\Big{\}}$
+$\geq$ $\frac{1}{L^{d}}\sum_{\mathbf{l}\in[L]^{d}}\min\Big{\{}\frac{e^{-\epsilon}}{8L^{2}},\ \frac{1/2-\delta}{n}\cdot\frac{dL^{d-1}(L-1/L)}{96}\Big{\}}$
+Choosing $L$ at $\lceil{(n/d)}^{1/(d+2)}\rceil$ yields the expected MSE as $\tilde{\Omega}(\frac{d^{2/(d+2)}}{n^{2/(d+2)}})$.
+As we are concerned with the asymptotic MSE with respect to $n$, the parameters $(\epsilon,\delta)$ are suppressed in the bound.
+Because the expected MSE bounds will never exceed the worst-case MSE bounds, we can establish that for an input data domain $\mathbb{X}=[0,1]^{d}$, the MSE lower bound $\max_{x\in\mathbb{X}}R(x)\geq\tilde{\Omega}(\frac{d^{2/(d+2)}}{n^{2/(d+2)}})$.
+Then, for the re-scaled domain $\mathbb{X}^{\prime}=[0,{1}/{\sqrt{d}}]^{d}$, the MSE lower bound is $\max_{x^{\prime}\in\mathbb{X}^{\prime}}R(x^{\prime})\geq\tilde{\Omega}(\frac{d^{-d/(d+2)}}{n^{2/(d+2)}})$.
+Finally, using the fact that $[0,{1}/{\sqrt{d}}]^{d}\subseteq\mathbb{B}_{2,1}(\{0\}^{d})$, we can conclude that the MSE lower bound over the domain $\mathbb{B}_{2,1}(\{0\}^{d})$ is $\tilde{\Omega}(\frac{1}{n^{2/(d+2)}})$ when $d$ is fixed and $n$ is sufficiently large.
+For the $d=1$ case, the [7] has established the $\tilde{\Omega}(\frac{1}{n^{2/3}})$ bound; the concurrent work in [5] also extends to multi-dimension case for the hyperspherical unit domain and obtains similar results.